### v1

# A modern perspective on rational homotopy theory

Abstract.
In Quillen’s paper on rational homotopy theory, the category of $1$-reduced simplicial sets is endowed with a family of model structures, the most prominent of which is the one whose weak equivalences are the rational homotopy equivalences. In this paper, we give a modern approach to this family of model structures.
1. Introduction
Rational homotopy theory studies $1$-connected spaces by simplifying the computation of their homotopy groups. The non-vanishing homotopy groups of a $1$-connected space are abelian, so their torsion can be discarded by tensoring with the rational numbers. Discarding torsion makes homotopy group computations more tractable: in the case of spheres, their rational homotopy groups were computed by Serre in [Ser51]. The success of rational homotopy theory has led to its extension to nilpotent spaces by Bousfield and Kan in [BK72]; Sullivan in [Sul74] and [Sul77]; Hilton, Mislin, and Roitberg in [HMR75]; Bousfield and Gugenheim in [BG76]; and Neisendorfer in [Nei78]. Ultimately, rational homotopy theory has been extended to all spaces by Casacuberta and Peschke in [CP93]; Gómez-Tato, Halperin, and Tanré in [GTHT00]; and Bastardas and Casacuberta in [BC01].
Model categories, as introduced by Quillen in [Qui67], provide a formal framework in which many tools and constructions from homotopy theory are defined. For instance, every model category admits a homotopy category, which is the localization of the model category at its class of weak equivalences. Thus, a model structure on a category endows the category with a homotopy theory and gives us tools to study that homotopy theory.
In [Qui69], Quillen applies his recently introduced notion of a model category to study rational homotopy theory, and especially how rational homotopy theory translates difficult problems in geometry and topology to more feasible problems in algebra. More specifically, Quillen produces model categories both on the topological side and on the algebraic side, and then proves that specific functorial constructions between them induce equivalences of these model categories. The notion of equivalence that we are referring to is that of a Quillen equivalence; in particular, a Quillen equivalence induces an equivalence of homotopy categories. On the algebraic side, Quillen produces model structures on rational differential graded (DG) Lie algebras and rational DG coalgebras, which are important algebraic invariants for computations in rational homotopy theory.
In this paper, we focus on the topological side, where Quillen produces a family of model structures on the category of $1$-reduced simplicial sets: simplicial sets whose $1$-skeleton is a point. As the category of $1$-connected spaces is ill-behaved with respect to important types of limits and colimits, such as pullbacks, the passage to $1$-reduced simplicial sets is necessary for Quillen’s model-categorical approach to rational homotopy theory.
This family of model structures interpolates between ordinary homotopy theory, where the weak equivalences are weak homotopy equivalences, and rational homotopy theory, where the weak equivalences are rational homotopy equivalences. Namely, for a multiplicative subset $M$ of $\mathbb{Z}$, Quillen endows the category of $1$-reduced simplicial sets with a model structure in which a map is a weak equivalence if and only if it induces an isomorphism of homotopy groups after tensoring with $M^{-1}\mathbb{Z}$; we call such a map an $M$-local homotopy equivalence.
Tensoring with $M^{-1}\mathbb{Z}$ for different choices of a multiplicative subset $M$ of $\mathbb{Z}$ allows us to preserve specific types of torsion in homotopy groups. For instance, for a prime $p$, taking $M=\mathbb{Z}\setminus p\mathbb{Z}$ and tensoring with the $p$-local integers $\mathbb{Z}_{(p)}$ specifically preserves $p$-torsion. Thus, we can study the torsion one prime at a time.
The following table depicts the interpolation given by varying our choice of a multiplicative subset $M$ of $\mathbb{Z}$.
$M$$\{1\}$ $\mathbb{Z}\setminus 2\mathbb{Z}$$\mathbb{Z}\setminus\{0\}$
[…]
torsion discarded by $-\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}$ none all but $2$-torsion all
$M$-local homotopy equivalence weak $2$-local rational
To establish this family of model structures, Quillen verifies the model category axioms one-by-one. In later years, though, further advancements in the study of model categories have provided us with powerful tools that enable us to induce new model structures from old without needing to perform such a painstaking verification.
The motivating idea behind our work in this paper is that we can use these modern tools to provide a new approach to Quillen’s family of model structures on $1$-reduced simplicial sets. The first tool we employ is that of left transfer, as developed by Bayeh, Hess, Karpova, Kędziorek, Riehl, and Shipley in [BHK+15]. This tool allows us to transfer the model structure on pointed simplicial sets to a model structure on $1$-reduced simplicial sets whose weak equivalences are the weak homotopy equivalences. Then, we left Bousfield localize the latter model structure to obtain the rest of Quillen’s family. Although the tool of left Bousfield localization dates back to the work of Bousfield in [Bou75] and [Bou77] and Farjoun in [Far92] and [Far95b], we rely on the more recent existence results for left Bousfield localization developed by Hirschhorn in [Hir03] and by Barwick in [Bar10].
For the left transfer, we observe that a $1$-reduced simplicial set has a unique 00-simplex, so there is a forgetful functor from the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets to the category ${\mathcal{SS}\mathrm{ets}}_{\ast}$ of pointed simplicial sets. This functor admits a right adjoint given by the $1$st Eilenberg subcomplex of a pointed simplicial set as defined by May in [May67]; in fact, Quillen uses this construction to establish his family of model structures. As the $1$st Eilenberg subcomplex depends on the basepoint, the passage to pointed simplicial sets, rather than unpointed ones, is necessary. We use this adjunction to left-transfer the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ to the following model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$.
Theorem 1.1.
There is a model structure on the category ${\mathcal{SS}\mathrm{ets}}_{1}$ in which a map $f$ is
  1. (1)
    a weak equivalence if and only if $f$ is a weak homotopy equivalence; and
  2. (2)
    a cofibration if and only if $f$ is a monomorphism.
Then, for the left Bousfield localization, we characterize, for a multiplicative subset $M$ of $\mathbb{Z}$, the $M$-local spaces and the $M$-local homotopy equivalences in the language of function complexes. Our characterizations can be viewed as the enriched-in-simplicial-sets version of the work of Casacuberta, Peschke, and Pfenniger in [CPP92], which uses hom-sets rather than function complexes. Using these characterizations, we left Bousfield localize the model category from Theorem 1.1 to recover the entire family of model structures on ${\mathcal{SS}\mathrm{ets}}_{1}$ from [Qui69]. We say that a simplicial set $X$ is $M$-local if $X$ is a Kan complex whose higher homotopy groups are $M^{-1}\mathbb{Z}$-modules.
Theorem 1.2.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ on the category of $1$-reduced simplicial sets in which a map $f$ is
  1. (1)
    a weak equivalence if and only if $f$ is an $M$-local homotopy equivalence; and
  2. (2)
    a cofibration if and only if $f$ is a monomorphism.
A $1$-reduced simplicial set $X$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $X$ is $M$-local.
We also give an alternative description of this family of model structures using Bousfield localization with respect to homology, as introduced in [Bou75], in conjunction with the result on left-induced model structures from [BHK+15] that we use for Theorem 1.1. The idea is that, rather than first left-transfer the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ and then localize, we can try to localize the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ with respect to homology with $M^{-1}\mathbb{Z}$-coefficients and then left-transfer this localized model structure, which we denote by $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$, to ${\mathcal{SS}\mathrm{ets}}_{1}$. We show that this approach also yields the family of model structures from Theorem 1.2. Schematically, the diagram of model-categorical operations
[…]
commutes, thus giving us two modern descriptions of the family of model structures on ${\mathcal{SS}\mathrm{ets}}_{1}$.
Another motivating goal of ours is to show that these model structures have properties that one generally desires and expects in the context of model categories. The tools that we invoke guarantee us most of these properties. An exception is the property of being a simplicial model category, as defined by Quillen in [Qui67]. We establish this property explicitly by adapting the techniques of Goerss and Jardine in [GJ99] to show that the model category from Theorem 1.1 is simplicial. As left Bousfield localization preserves the simplicial structure, this structure is bestowed upon the entire family of model structures on $1$-reduced simplicial sets from Theorem 1.2. We remark that, in [Bur21], Burke proves the analog of our simplicial structure for reduced simplicial sets: simplicial sets with a unique 00-simplex. However, Burke’s line of argument is different from that in this paper.
Also, we partially extend Theorem 1.2 to the category $\mathcal{SS}\mathrm{ets}$ of simplicial sets. Namely, we produce a family of model structures on simplicial sets that describes a generalization of rational homotopy theory to all spaces. In [Iva22], Ivanov gives an overview of the various approaches to rational homotopy theory beyond the framework of $1$-connectedness or nilpotency. We adopt the approach of Gómez-Tato, Halperin, and Tanré from [GTHT00]. In this approach, one rationalizes a connected space $X$ by applying fiberwise rationalization to the fibration
$\widetilde{X}\rightarrow X\rightarrow K(\pi_{1}(X),1)$
given by the $1$st Postnikov approximation of $X$; its homotopy fiber $\widetilde{X}$ is the universal cover of $X$, which is a $1$-connected space that can be rationalized in the classical sense. The tool of fiberwise localization is invoked by Sullivan in [Sul74] and investigated by Llerena in [Lle82] and [Lle85] and Farjoun in [Far95a]. This rationalization theory for non-nilpotent spaces is also studied in recent work of Rivera, Wierstra, and Zeinalian in [RWZ21].
We partially extend Theorem 1.2 to all simplicial sets as follows.
Theorem 1.3.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ on the category of simplicial sets in which a simplicial set $X$ is fibrant if and only if $X$ is $M$-local.
We also prove the counterpart of this theorem for the category ${\mathcal{T}\mathrm{op}}$ of topological spaces. The reason why Theorem 1.3 only partially extends Theorem 1.2 to all simplicial sets is that we lack a description of the weak equivalences other than the standard one for a left Bousfield localization. This deficiency is due to the fact that fiberwise rationalization satisfies a more intricate universal property than that of the rationalization of $1$-connected spaces, which causes our characterization of the weak equivalences in Theorem 1.2 to break down.
Lastly, we verify that our localization $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ of spaces with respect to rational homotopy from Theorem 1.3 does not coincide with Bousfield’s localization of spaces with respect to rational homology from [Bou75]. This result captures the fact that rational homotopy and rational homology do not agree beyond the framework of $1$-connected spaces; in fact, our proof of the result involves a specific instance of this general phenomenon.
Organization of the paper
Section 2 comprises background on $1$-reduced simplicial sets. In Section 3, we perform the left transfer to obtain Theorem 1.1. In Section 4, we prove some results about the fibrations in this left-induced model structure. In Section 5, we establish that the model category from Theorem 1.1 is simplicial. Section 6 consists of auxiliary characterization results from the rational homotopy theory of $1$-connected spaces. We use these results in Section 7 to perform the left Bousfield localization that yields Theorem 1.2. In Section 8, we give an alternative approach to Theorem 1.2 via Bousfield localization with respect to homology. In Section 9, we give an overview of rational homotopy theory beyond the setting of $1$-connectedness. In Section 10, we perform the left Bousfield localization that yields Theorem 1.3. In Section 11, we establish the analog of Theorem 1.3 for topological spaces. Lastly, in Section 12, we show that our localization of spaces with respect to rational homotopy is not equal to the Bousfield localization of spaces with respect to rational homology.
Acknowledgments
I am grateful to my PhD advisor at the University of Virginia, Julie Bergner, for her guidance, input, feedback, and support that made the work in this paper possible. I would also like to thank Nicholas Kuhn for suggesting the approach of using Bousfield localization with respect to homology, as well as Scott Balchin and Kimball Strong for enlightening conversations about this work.
2. $1$-reduced simplicial sets
We begin with some background on $1$-reduced simplicial sets. We define the $1$-skeleton and the $1$-coskeleton of a simplicial set. Using these notions, we define $1$-reduced simplicial sets. Then, we define the $1$-reduction of a simplicial set, as well as the $1$st Eilenberg subcomplex of a pointed simplicial set. These constructions provide us with adjunctions between $1$-reduced simplicial sets and (pointed) simplicial sets that we rely on in this paper.
2.1. Simplicial sets
We collect background and notation on simplicial sets; a reference is [GJ99, Chapter I].
The combinatorics involved in the definition of a simplicial set can be packaged in the simplex category $\Delta$.
Definition 2.1.
The simplex category $\Delta$ is the category whose
  1. (1)
    objects are the sets $[n]=\{0,1,\dots,n\}$ for a non-negative integer $n$; and
  2. (2)
    morphisms are the order-preserving maps $\theta\colon[n]\rightarrow[m]$.
Example 2.2.
For $0\leq i\leq n$, two special types of maps in $\Delta$ are
  1. (1)
    the $i$-th coface map $d^{i}\colon[n-1]\rightarrow[n]$; and
  2. (2)
    the $i$-th codegeneracy map $s^{i}\colon[n+1]\rightarrow[n]$.
These maps generate the simplex category $\Delta$ subject to a list of identities that determine their composites.
The simplex category $\Delta$ allows for a concise definition of a simplicial set. We write ${\mathcal{S}\mathrm{ets}}$ for the category of sets.
Definition 2.3.
A simplicial set is a functor $X\colon\Delta^{\operatorname{op}}\rightarrow{\mathcal{S}\mathrm{ets}}$. In detail, a simplicial set comprises
  1. (1)
    a set of $n$-simplices $X_{n}$ for every non-negative integer $n$;
  2. (2)
    face maps $d_{i}\colon X_{n}\rightarrow X_{n-1}$ for $0\leq i\leq n$; and
  3. (3)
    degeneracy maps $s_{i}\colon X_{n}\rightarrow X_{n+1}$ for $0\leq i\leq n$;
subject to a list of simplicial identities that govern how the faces and degeneracies are composed.
We write ${\mathcal{SS}\mathrm{ets}}$ for the category of simplicial sets and simplicial maps (natural transformations) between them.
Example 2.4.
The standard $n$-simplex is the representable simplicial set
$\Delta[n]=\operatorname{Hom}_{\Delta}\left(-,[n]\right).$
The identity $\operatorname{id}_{[n]}$ gives an $n$-simplex $\sigma_{n}$ of $\Delta[n]$. For a simplicial set $X$, the Yoneda lemma yields the bijection
$X_{n}\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}}\left(\Delta[n],X\right)$
that is natural in $[n]$ and in $X$. Therefore, we occasionally view an $n$-simplex of $X$ as a map $\Delta[n]\rightarrow X$.
Example 2.5.
The boundary $\partial\Delta[n]$ is the smallest subcomplex of $\Delta[n]$ containing the faces $d_{i}(\sigma_{n})$ for $0\leq i\leq n$. We write
$i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]$
for the boundary inclusion.
Example 2.6.
For $n\geq 1$ and $0\leq k\leq n$, the horn $\Lambda[n,k]$ is the smallest subcomplex of $\Delta[n]$ containing $d_{i}(\sigma_{n})$ for $i\neq k$. We write
$j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]$
for the horn inclusion. Observe that horns are undefined for $n=0$.
We can geometrically realize a simplicial set $X$ to obtain a CW complex $|X|$. Writing ${\mathcal{T}\mathrm{op}}$ for the category of topological spaces, we denote the geometric realization functor by
$|-|\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{T}\mathrm{op}}.$
Conversely, for a topological space $Y$, its singular complex $\operatorname{Sing}(Y)$ is the simplicial set with $n$-simplices
$\operatorname{Sing}(Y)_{n}=\operatorname{Hom}_{\mathcal{T}\mathrm{op}}\left(|\Delta[n]|,Y\right).$
This pair of functors forms an adjunction that we use to oscillate between simplicial sets and topological spaces.
Proposition 2.7 ([GJ99, Chapter I, Proposition 2.2]).
The geometric realization functor $|-|\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{T}\mathrm{op}}$ is left adjoint to the singular complex functor $\operatorname{Sing}\colon{\mathcal{T}\mathrm{op}}\rightarrow{\mathcal{SS}\mathrm{ets}}$.
Our work largely involves pointed simplicial sets, which are simplicial sets with a designated 00-simplex as their basepoint. We write ${\mathcal{SS}\mathrm{ets}}_{\ast}$ for the category of pointed simplicial sets and pointed simplicial maps between them. The following adjunction is verified just like its analog in topological spaces.
Lemma 2.8.
The forgetful functor $V\colon{\mathcal{SS}\mathrm{ets}}_{\ast}\rightarrow{\mathcal{SS}\mathrm{ets}}$ has a left adjoint
$(-)_{+}\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$
defined by attaching a disjoint basepoint.
2.2. The $1$-skeleton and the $1$-coskeleton of a simplicial set
To define $1$-reduced simplicial sets, we recall the $1$-skeleton of a simplicial set. We also recall the $1$-coskeleton of a simplicial set, which we refer to in the definition of the $1$st Eilenberg subcomplex of a pointed simplicial set. Our reference is [Ber18, Chapter 2.6].
Let $\Delta_{1}$ denote the full subcategory of the simplex category $\Delta$ with objects $[0]$ and $[1]$. The inclusion
$\Delta_{1}\rightarrow\Delta$
induces a truncation functor
$\operatorname{tr}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{S}\mathrm{ets}}^{\Delta_{1}^{\operatorname{op}}}.$
This truncation functor admits both a left adjoint
$s_{1}\colon{\mathcal{S}\mathrm{ets}}^{\Delta_{1}^{\operatorname{op}}}\rightarrow{\mathcal{SS}\mathrm{ets}}$
and a right adjoint
$c_{1}\colon{\mathcal{S}\mathrm{ets}}^{\Delta_{1}^{\operatorname{op}}}\rightarrow{\mathcal{SS}\mathrm{ets}},$
in terms of which we can define the $1$-skeleton and the $1$-coskeleton of a simplicial set.
Definition 2.9.
Let $X$ be a simplicial set.
  1. (1)
    The $1$-skeleton of $X$ is the simplicial set
    $\operatorname{sk}_{1}(X)=s_{1}\operatorname{tr}_{1}(X).$
  2. (2)
    The $1$-coskeleton of $X$ is the simplicial set
    $\operatorname{cosk}_{1}(X)=c_{1}\operatorname{tr}_{1}(X).$
In detail, both $\operatorname{sk}_{1}(X)$ and $\operatorname{cosk}_{1}(X)$ have set of 00-simplices $X_{0}$ and set of $1$-simplices $X_{1}$, but they differ in their higher-dimensional simplices. The higher-dimensional simplices of $\operatorname{sk}_{1}(X)$ are the iterated degeneracies of its $1$-simplices. On the contrary, $\operatorname{cosk}_{1}(X)$ has an $n$-simplex for every compatible collection of $(n-1)$-simplices. Equivalently, $\operatorname{cosk}_{1}(X)$ has an $n$-simplex for every map $\operatorname{sk}_{1}(\Delta[n])\rightarrow X$.
Remark 2.10.
It follows from this definition that, for a simplicial set $X$, we have an inclusion map
$\operatorname{sk}_{1}(X)\rightarrow X$
and a quotient map
$X\rightarrow\operatorname{cosk}_{1}(X),$
where the latter map identifies the simplices of $X$ that share the same $1$-skeleton.
Example 2.11.
Observe that the $1$-skeleton of the standard $2$-simplex $\Delta[2]$ is its boundary $\partial\Delta[2]$.
2.3. $1$-reduced simplicial sets
We are now ready to define $1$-reduced simplicial sets.
Definition 2.12.
A simplicial set $X$ is $1$-reduced if its $1$-skeleton $\operatorname{sk}_{1}(X)$ is a point. In detail, a simplicial set $X$ is $1$-reduced if it has a unique 00-simplex whose degeneracy is the unique $1$-simplex of $X$.
Remark 2.13.
Note that the geometric realization of a $1$-reduced simplicial set is a $1$-connected CW complex. It is also crucial to observe that a $1$-reduced simplicial set is unambiguously pointed, for it has a unique 00-simplex.
We write ${\mathcal{SS}\mathrm{ets}}_{1}$ for the category of $1$-reduced simplicial sets and simplicial maps between them.
Example 2.14.
For $n\geq 2$, the simplicial $n$-sphere
$S_{\mathrm{simp}}^{n}=\Delta[n]/\partial\Delta[n]$
is a $1$-reduced simplicial set consisting of a unique 00-simplex and a unique non-degenerate $n$-simplex.
The proof of the following lemma is straightforward. We write $\wedge$ for the smash product.
Lemma 2.15.
If $X$ is a $1$-reduced simplicial set and $Z$ is a pointed simplicial set, then $X\wedge Z$ is $1$-reduced.
2.4. The $1$-reduction of a simplicial set
We relate the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets to the category ${\mathcal{SS}\mathrm{ets}}$ of simplicial sets by means of a left adjoint to the forgetful functor $U\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}$. This left adjoint quotients out the $1$-skeleton of a simplicial set to yield a $1$-reduced simplicial set. Although quotients can be homotopically ill-behaved, this adjunction aids our model-categorical arguments later in this paper, for it repeatedly lets us translate lifting problems that we wish to solve into lifting problems that we can solve.
Definition 2.16.
The $1$-reduction of a simplicial set $X$ is the $1$-reduced simplicial set
$R_{1}(X)=X/\operatorname{sk}_{1}(X).$
Example 2.17.
The simplicial $2$-sphere is $S_{\mathrm{simp}}^{2}=R_{1}(\Delta[2])$ because, by Example 2.11, $\operatorname{sk}_{1}(\Delta[2])=\partial\Delta[2]$.
The $1$-reduction of a simplicial set is its largest $1$-reduced quotient in the sense of the following universal property, which is an instance of the universal property of quotients.
Lemma 2.18.
For every map $f\colon X\rightarrow Y$ from a simplicial set $X$ to a $1$-reduced simplicial set $Y$, there exists a unique map $\overline{f}\colon R_{1}(X)\rightarrow Y$ such that the diagram
${X}$${{R_{1}(X)}}$${{Y}}$$\scriptstyle{f}$$\scriptstyle{\overline{f}}$
commutes, where $X\rightarrow R_{1}(X)$ is the quotient map from $X$ to $R_{1}(X)=X/\operatorname{sk}_{1}(X)$.
This universal property translates to the promised adjunction.
Lemma 2.19.
The $1$-reduction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$ is left adjoint to the forgetful functor $U\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}$.
We note a consequence of this adjunction for later use.
Corollary 2.20.
If $X$ is a $1$-reduced simplicial set, then a map $\Delta[2]\rightarrow X$ is adjoint to a map $S_{\mathrm{simp}}^{2}\rightarrow X$.
Proof.
The claim is a consequence of Lemma 2.19 because $S_{\mathrm{simp}}^{2}=R_{1}(\Delta[2])$ by Example 2.17. ∎
2.5. The $1$st Eilenberg subcomplex of a pointed simplicial set
As we shall see, the approach of taking the $1$-reduction of a simplicial set does not play as nicely as we need it to for our homotopy-theoretic purposes, but we can do better. Because a $1$-reduced simplicial set has a unique 00-simplex, there is a forgetful functor
$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$
from the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets to the category ${\mathcal{SS}\mathrm{ets}}_{\ast}$ of pointed simplicial sets. This functor has not only a left adjoint given by $1$-reduction, but also a right adjoint given by the $1$st Eilenberg subcomplex of a pointed simplicial set. We present its definition, though we do not use it beyond this section, as its only feature that we need in this paper is the aforementioned adjunction, which we establish.
Recall from Remark 2.10 that, for a simplicial set $X$, we have a quotient map $X\rightarrow\operatorname{cosk}_{1}(X)$.
Definition 2.21 ([May67, Definition 8.3]).
The $1$st Eilenberg subcomplex $E_{1}(X)$ of a pointed simplicial set $X$ with basepoint $x_{0}$ is defined by the pullback square
${{E_{1}(X)}}$${X}$${{\Delta[0]}}$${{\operatorname{cosk}_{1}(X).}}$$\scriptstyle{x_{0}}$
In detail, $E_{1}(X)$ consists of all simplices $\chi$ of $X$ such that every 00-dimensional face of $\chi$ is $x_{0}$ and every $1$-dimensional face of $\chi$ is the degeneracy of $x_{0}$.
It follows from this definition that the $1$st Eilenberg subcomplex $E_{1}(X)$ of a pointed simplicial set $X$ is $1$-reduced.
The $1$st Eilenberg subcomplex of a pointed simplicial set is its largest $1$-reduced subcomplex in the sense of the following universal property that is straightforward to verify.
Lemma 2.22.
Every pointed map $f\colon X\rightarrow Y$ from a $1$-reduced simplicial set $X$ to a pointed simplicial set $Y$ maps into the $1$st Eilenberg subcomplex $E_{1}(Y)$ of $Y$. Thus, there is a unique map $\hat{f}\colon X\rightarrow E_{1}(Y)$ such that the diagram
${X}$${{E_{1}(Y)}}$${{Y}}$$\scriptstyle{\hat{f}}$$\scriptstyle{f}$
commutes, where $E_{1}(Y)\rightarrow Y$ is the inclusion map.
This universal property amounts to the promised adjunction, which is stated in Part (2) of the following lemma.
Lemma 2.23.
Let $F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$ denote the forgetful functor.
  1. (1)
    The $1$-reduction functor ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}_{\ast}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$ is left adjoint to $F$.
  2. (2)
    The $1$st Eilenberg subcomplex functor $E_{1}\colon{\mathcal{SS}\mathrm{ets}_{\ast}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$ is right adjoint to $F$.
Lemma 2.23 implies that small limits and colimits in ${\mathcal{SS}\mathrm{ets}}_{1}$ are constructed as in ${\mathcal{SS}\mathrm{ets}}_{\ast}$. Given a diagram
$L\colon\mathcal{D}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$
in $1$-reduced simplicial sets, we form its (co)limit by post-composing $L$ with the forgetful functor
$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$
and forming the (co)limit of the diagram
$FL\colon\mathcal{D}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$
in pointed simplicial sets. This (co)limit is a $1$-reduced simplicial set, and all the structure maps are in ${\mathcal{SS}\mathrm{ets}}_{1}$.
In particular, the initial object of ${\mathcal{SS}\mathrm{ets}}_{1}$ is the point $\Delta[0]$, not the empty simplicial set $\emptyset$. Thus, the forgetful functor $U\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}$ does not preserve the initial object, so it cannot admit a right adjoint.
3. The left transfer to $1$-reduced simplicial sets
We transfer the model structure on the category ${\mathcal{SS}\mathrm{ets}}_{\ast}$ of pointed simplicial sets along the adjunction
$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}$
to a model structure on the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets whose weak equivalences are the weak homotopy equivalences. This left transfer is an application of [BHK+15, Theorem 2.23].
3.1. Model categories
We collect some background on model categories from [DS95, §3] and [Hir03, Chapters 7, 11, and 13]. We first recall a concept in the homotopy theory of simplicial sets from [GJ99, Chapter I].
Definition 3.1.
A Kan fibration is a simplicial map $p\colon X\rightarrow Y$ that has the right lifting property with respect to the set of all horn inclusions
$J=\left\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{ and}0\leq k\leq n\right\}.$
In other words, a simplicial map $p\colon X\rightarrow Y$ is a Kan fibration if every lifting problem of the form
${{\Lambda[n,k]}}$${X}$${{\Delta[n]}}$${Y}$$\scriptstyle{j_{n,k}}$$\scriptstyle{p}$
can be solved. A Kan complex is a simplicial set $X$ such that the unique map $X\rightarrow\Delta[0]$ is a Kan fibration.
Example 3.2 ([GJ99, Chapter I, Lemma 3.3]).
If $Y$ is a space, then $\operatorname{Sing}(Y)$ is a Kan complex.
The homotopy theory of simplicial sets can be studied systematically in the framework of model categories.
Definition 3.3.
A model category comprises a category $\mathcal{M}$ that has all finite limits and colimits, as well as three distinguished classes of morphisms in $\mathcal{M}$: a class of weak equivalences, a class of cofibrations, and a class of fibrations. A weak equivalence that is also a cofibration is called an acyclic cofibration. A weak equivalence that is also a fibration is called an acyclic fibration. This data satisfies the axioms in [DS95, Definition 3.3].
Remark 3.4.
In a model category, the weak equivalences and the fibrations together determine the cofibrations. Dually, the weak equivalences and the cofibrations together determine the fibrations.
Remark 3.5.
Let $\mathcal{M}$ be a model category with initial object $\emptyset$ and terminal object $\ast$. An object $X$ of $\mathcal{M}$ is
  1. (1)
    cofibrant if the unique morphism $\emptyset\rightarrow X$ is a cofibration; and
  2. (2)
    fibrant if the unique morphism $X\rightarrow\ast$ is a fibration.
Then, for an object $Y$ of $\mathcal{M}$,
  1. (1)
    a cofibrant replacement of $Y$ is a weak equivalence $A\rightarrow Y$ with cofibrant source $A$; and
  2. (2)
    a fibrant replacement of $Y$ is a weak equivalence $Y\rightarrow B$ with fibrant target $B$.
Example 3.6 ([Hir03, Theorems 7.10.12 and 7.10.13]).
The category ${\mathcal{SS}\mathrm{ets}}$ of simplicial sets has a model structure in which a map $f$ is
  1. (1)
    a weak equivalence if and only if $f$ is a weak homotopy equivalence;
  2. (2)
    a cofibration if and only if $f$ is a monomorphism; and
  3. (3)
    a fibration if and only if $f$ is a Kan fibration.
The category ${\mathcal{SS}\mathrm{ets}}_{\ast}$ of pointed simplicial sets also has a model structure with the same description.
The model categories ${\mathcal{SS}\mathrm{ets}}$ and ${\mathcal{SS}\mathrm{ets}}_{\ast}$ have the additional property of being cofibrantly generated. In practice, the property of being cofibrantly generated facilitates the detection of (acyclic) fibrations in a model category.
Definition 3.7 ([Hir03, Definition 11.1.2]).
A model category $\mathcal{M}$ is cofibrantly generated if there exist
  1. (1)
    a set $I$ of morphisms in $\mathcal{M}$, called a generating set of cofibrations for $\mathcal{M}$; and
  2. (2)
    a set $J$ of morphisms in $\mathcal{M}$, called a generating set of acyclic cofibrations for $\mathcal{M}$,
such that
  1. (1)
    both $I$ and $J$ permit the small object argument [Hir03, Definition 10.5.15];
  2. (2)
    a morphism $p$ is an acyclic fibration if and only if $p$ has the right lifting property with respect to $I$; and
  3. (3)
    a morphism $p$ is a fibration if and only if $p$ has the right lifting property with respect to $J$.
Example 3.8 ([Hir03, Examples 11.1.6 and 11.1.7]).
Both ${\mathcal{SS}\mathrm{ets}}$ and ${\mathcal{SS}\mathrm{ets}}_{\ast}$ are cofibrantly generated.
  1. (1)
    A generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}$ is the set of all boundary inclusions
    $I=\left\{i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]\mid n\geq 0\right\}.$
    Attaching a disjoint basepoint yields a generating set of cofibrations $I_{+}$ for ${\mathcal{SS}\mathrm{ets}}_{\ast}$.
  2. (2)
    A generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}$ is the set of all horn inclusions
    $J=\left\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{ and}0\leq k\leq n\right\}.$
    Attaching a disjoint basepoint yields a generating set of acyclic cofibrations $J_{+}$ for ${\mathcal{SS}\mathrm{ets}}_{\ast}$.
What is more, ${\mathcal{SS}\mathrm{ets}}$ and ${\mathcal{SS}\mathrm{ets}}_{\ast}$ belong in the following class of well-behaved model categories, which is originally due to J.H. Smith and is particularly expedient for our purposes of left Bousfield localization.
Definition 3.9 ([Dug01, Definition 2.1]).
A model category $\mathcal{M}$ is combinatorial if $\mathcal{M}$ is cofibrantly generated and its underlying category is locally presentable [AR94, Definition 1.17].
Example 3.10 ([AR94, Example 1.12]).
Both ${\mathcal{SS}\mathrm{ets}}$ and ${\mathcal{SS}\mathrm{ets}}_{\ast}$ are locally presentable, thus combinatorial.
Lastly, ${\mathcal{SS}\mathrm{ets}}$ and ${\mathcal{SS}\mathrm{ets}}_{\ast}$ satisfy the following model-categorical property that permits left Bousfield localization.
Definition 3.11 ([Hir03, Definition 13.1.1(1)]).
A model category $\mathcal{M}$ is left proper if every pushout of a weak equivalence along a cofibration is a weak equivalence.
The following result guarantees the left properness of almost every model category that appears in this paper.
Proposition 3.12 ([Hir03, Corollary 13.1.3(1)]).
If every object of a model category $\mathcal{M}$ is cofibrant, then $\mathcal{M}$ is left proper.
3.2. The left transfer to $1$-reduced simplicial sets
Firstly, to appreciate the necessity of using pointed simplicial sets and their $1$st Eilenberg subcomplex, we explain why we cannot use the $1$-reduction functor $R_{1}$ to right-induce, in the sense of [Hir03, Theorem 11.3.2], a model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$ along the adjunction
${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$
from Lemma 2.19. We write $J$ for the set of all horn inclusions in ${\mathcal{SS}\mathrm{ets}}$. If such a right transfer were possible, then, in the resulting right-induced model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$,
  1. (1)
    the class of weak equivalences would be the class of weak homotopy equivalences; and
  2. (2)
    the set $R_{1}(J)$ would be a generating set of acyclic cofibrations.
Thus, every map in $R_{1}(J)$ would be a weak homotopy equivalence. We verify that this necessary condition fails.
Lemma 3.13.
For $n\geq 1$ and $0\leq k\leq n$, let $j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]$ denote the horn inclusion in ${\mathcal{SS}\mathrm{ets}}$.
  1. (1)
    If $n=1$, then ${R}_{1}\left(j_{1,k}\right)$ is the identity map of the point.
  2. (2)
    If $n=2$, then ${R}_{1}\left(j_{2,k}\right)\colon\Delta[0]\rightarrow S_{\mathrm{simp}}^{2}$ picks out the unique 00-simplex of the simplicial $2$-sphere.
  3. (3)
    If $n\geq 3$, then $|{R}_{1}\left(j_{n,k}\right)|$ is the inclusion of a strong deformation retract in topological spaces.
Therefore, ${R}_{1}\left(j_{n,k}\right)$ is a weak homotopy equivalence if and only if $n\neq 2$.
Proof.
Parts (1) and (2) are straightforward. For (3), if $n\geq 3$, then $\operatorname{sk}_{1}\left(\Lambda[n,k]\right)=\operatorname{sk}_{1}\left(\Delta[n]\right)$, so the inclusion of a strong deformation retract $|j_{n,k}|$ descends to the inclusion of a strong deformation retract $|{R}_{1}\left(j_{n,k}\right)|$. ∎
On that note, we also record a computation of a similar nature for future reference in this section.
Lemma 3.14.
If a simplicial map $f\colon X\rightarrow Y$ is a monomorphism, then so is its $1$-reduction ${R}_{1}(f)$.
We left-induce a model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$ by transferring the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ along the adjunction
$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}$
from Lemma 2.23(2). To do so, we apply the following general result on left-induced model structures.
Theorem 3.15 ([BHK+15, Theorem 2.23]).
Let $F\colon\mathcal{C}\rightleftarrows\mathcal{M}\colon G$ be an adjunction, where $\mathcal{C}$ is a locally presentable category and $\mathcal{M}$ is a combinatorial model category with class of cofibrations $Q$. Suppose that the following condition holds: if a morphism $g$ in $\mathcal{C}$ has the right lifting property with respect to the class $F^{-1}(Q)$, then $F(g)$ is a weak equivalence in $\mathcal{M}$. Then, there is a combinatorial model structure on $\mathcal{C}$ in which a morphism $f$ is
  1. (1)
    a weak equivalence if and only if $F(f)$ is a weak equivalence in $\mathcal{M}$; and
  2. (2)
    a cofibration if and only if $F(f)$ is a cofibration in $\mathcal{M}$.
As ${\mathcal{SS}\mathrm{ets}}_{1}$ is a locally presentable category [AR94, Example 1.12] and ${\mathcal{SS}\mathrm{ets}}_{\ast}$ is a combinatorial model category by Example 3.10, it remains to verify the condition in Theorem 3.15.
Lemma 3.16.
Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets. If $f$ has the right lifting property with respect to every monomorphism of $1$-reduced simplicial sets, then $f$ is a weak homotopy equivalence.
Proof.
We show that $f$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}$. For $n\geq 0$, let $i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]$ denote the boundary inclusion in ${\mathcal{SS}\mathrm{ets}}$. By Lemma 3.14, ${R}_{1}(i_{n})$ is a monomorphism of $1$-reduced simplicial sets. Therefore, the hypothesis informs us that, for every $n\geq 0$, $f$ has the right lifting property with respect to ${R}_{1}(i_{n})$. Then, the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19 yields that, for every $n\geq 0$, $f$ has the right lifting property with respect to the boundary inclusion $i_{n}$. We conclude that $f$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}$. ∎
We begin our recovery of the family of model structures on $1$-reduced simplicial sets in [Qui69, Part II, Theorem 2.2] and the characterization of fibrant objects in [Qui69, Part II, Corollary 2.6].
Theorem 3.17.
There is a combinatorial and left proper model structure on the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets in which a map $f$ is
  1. (1)
    a weak equivalence if and only if $f$ is a weak homotopy equivalence; and
  2. (2)
    a cofibration if and only if $f$ is a monomorphism.
In this model structure, a $1$-reduced simplicial set $X$ is fibrant if and only if $X$ is a Kan complex.
Proof.
Theorem 3.15 yields every claim other than the characterization of fibrant objects. Every $1$-reduced Kan complex $X$ is fibrant in ${\mathcal{SS}\mathrm{ets}}_{1}$ because, by virtue of being a Kan complex, $X$ has the right lifting property with respect to the class of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$. Conversely, if a $1$-reduced simplicial set $X$ is fibrant in ${\mathcal{SS}\mathrm{ets}}_{1}$, then we show that $X$ is a Kan complex. For $n\geq 1$ and $0\leq k\leq n$, we solve a lifting problem
${{\Lambda[n,k]}}$${X}$${{\Delta[n].}}$$\scriptstyle{j_{n,k}}$
Using the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19, we may equivalently solve a lifting problem
${{{R}_{1}\left(\Lambda[n,k]\right)}}$${X}$${{{R}_{1}\left(\Delta[n]\right).}}$$\scriptstyle{{R}_{1}\left(j_{n,k}\right)}$
In light of Lemma 3.13, we split our solution in two separate cases. If $n\neq 2$, then, by Lemmas 3.13 and 3.14, ${R}_{1}\left(j_{n,k}\right)$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, so our lifting problem has a solution since $X$ is fibrant in ${\mathcal{SS}\mathrm{ets}}_{1}$. If $n=2$, then we solve a lifting problem
[…]
where $x_{0}$ denotes the unique 00-simplex of $X$, by setting $\psi$ to be the constant map at $x_{0}$. ∎
The left transfer in Theorem 3.17 renders the adjunction
$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}$
from Lemma 2.23(2) a Quillen adjunction; in particular, we note that $E_{1}$ preserves fibrant objects.
Corollary 3.18.
If $X$ is a pointed Kan complex, then $E_{1}(X)$ is a Kan complex.
In particular, we occasionally appeal to a special instance of Corollary 3.18.
Corollary 3.19.
If $Y$ is a pointed topological space, then $E_{1}\operatorname{Sing}(Y)$ is a Kan complex.
Proof.
By Example 3.2, $\operatorname{Sing}(Y)$ is a pointed Kan complex, so the claim follows from Corollary 3.18. ∎
4. Fibrations of $1$-reduced simplicial sets
We study the fibrations in the model category ${\mathcal{SS}\mathrm{ets}}_{1}$. Namely, we produce functorial fibrant replacements, a generating set of cofibrations, and a set of acyclic cofibrations that detects fibrant objects. Moreover, we characterize when a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ is a Kan fibration, and we find a counterintuitive fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
4.1. Fibrant replacement of $1$-reduced simplicial sets
We invoke Corollary 3.19 to produce functorial fibrant replacements in ${\mathcal{SS}\mathrm{ets}}_{1}$. Let $X$ be a $1$-reduced simplicial set. The geometric realization-singular complex adjunction from Proposition 2.7 provides the natural map
$X\rightarrow\operatorname{Sing}(|X|),$
which maps into the $1$st Eilenberg subcomplex $E_{1}\operatorname{Sing}(|X|)$. Thus, we obtain the commutative diagram
${X}$${{E_{1}\operatorname{Sing}(|X|)}}$${{\operatorname{Sing}(|X|),}}$
where $E_{1}\operatorname{Sing}(|X|)\rightarrow\operatorname{Sing}(|X|)$ is the inclusion map. We show that the natural map
$X\rightarrow E_{1}\operatorname{Sing}(|X|)$
is a weak homotopy equivalence, thus a functorial fibrant replacement of $X$ in ${\mathcal{SS}\mathrm{ets}}_{1}$. To do so, it suffices to show that the other two maps in the commutative diagram are weak homotopy equivalences. Firstly, Part (1) of the following theorem informs us that the natural map
$X\rightarrow\operatorname{Sing}(|X|)$
is a weak homotopy equivalence; Part (2) is for later use in this argument.
Theorem 4.1 ([GJ99, Chapter I, Theorem 11.4]).
  1. (1)
    For every simplicial set $X$, the natural map
    $X\rightarrow\operatorname{Sing}(|X|)$
    is a weak homotopy equivalence.
  2. (2)
    For every topological space $Y$, the natural map
    $|\operatorname{Sing}(Y)|\rightarrow Y$
    is a weak homotopy equivalence.
Remark 4.2.
Part (1) informs us that, for a simplicial set $X$, the natural map
$X\rightarrow\operatorname{Sing}(|X|)$
is a functorial fibrant replacement of $X$ in ${\mathcal{SS}\mathrm{ets}}$, as $\operatorname{Sing}(|X|)$ is a Kan complex by Example 3.2.
As for the $1$st Eilenberg subcomplex inclusion
$E_{1}\operatorname{Sing}(|X|)\rightarrow\operatorname{Sing}(|X|),$
it is also a weak homotopy equivalence by a special case of the following theorem.
Theorem 4.3 ([May67, Theorem 8.4(iii)]).
If $X$ is a pointed Kan complex, then the $1$st Eilenberg subcomplex inclusion $E_{1}(X)\rightarrow X$ induces, for every $n\geq 2$, the isomorphism
$\pi_{n}\left(E_{1}(X)\right)\xrightarrow{\cong}\pi_{n}(X).$
In particular, if $X$ is a pointed and $1$-connected Kan complex, then the inclusion
$E_{1}(X)\rightarrow X$
is a weak homotopy equivalence.
Remark 4.4.
In other words, if $X$ is a pointed Kan complex, then the $1$st Eilenberg subcomplex inclusion
$E_{1}(X)\rightarrow X$
is a $1$-connected cover of $X$.
As an aside, using the construction of the homotopy groups of Kan complexes in [GJ99, Chapter I.7] together with Lemma 2.22, we can provide an alternative proof of Theorem 4.3 from that in [May67, Theorem 8.4(iii)].
Proof.
Let $X$ be a pointed Kan complex. Let $n\geq 2$. We first show that $\pi_{n}\left(E_{1}(X)\right)\rightarrow\pi_{n}(X)$ is an epimorphism. As $X$ is a Kan complex, every element of $\pi_{n}(X)$ is of the form $[\omega]$ for a map
$\omega\colon S_{\mathrm{simp}}^{n}\rightarrow X.$
As $n\geq 2$, the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ is $1$-reduced, so the universal property of the $1$st Eilenberg subcomplex from Lemma 2.22 yields the commutative diagram
[…]
It remains to show that $\pi_{n}\left(E_{1}(X)\right)\rightarrow\pi_{n}(X)$ is a monomorphism. Given a pointed homotopy
$H\colon S_{\mathrm{simp}}^{n}\wedge\Delta[1]_{+}\rightarrow X,$
we observe that $S_{\mathrm{simp}}^{n}\wedge\Delta[1]_{+}$ is $1$-reduced by Lemma 2.15 because $S_{\mathrm{simp}}^{n}$ is $1$-reduced, so the universal property of the $1$st Eilenberg subcomplex from Lemma 2.22 yields the commutative diagram
[…]
Corollary 4.5.
If $Y$ is a pointed and $1$-connected topological space, then the inclusion
$E_{1}\operatorname{Sing}(Y)\rightarrow\operatorname{Sing}(Y)$
is a weak homotopy equivalence.
Proof.
By Example 3.2, $\operatorname{Sing}(Y)$ is a Kan complex. By Theorem 4.1(2), $\operatorname{Sing}(Y)$ is $1$-connected. ∎
In conclusion, we obtain functorial fibrant replacements in ${\mathcal{SS}\mathrm{ets}}_{1}$.
Proposition 4.6.
Let $X$ be a $1$-reduced simplicial set. Then, the natural map
$X\rightarrow E_{1}\operatorname{Sing}(|X|)$
is a functorial fibrant replacement of $X$ in ${\mathcal{SS}\mathrm{ets}}_{1}$.
4.2. Generating sets of cofibrations of $1$-reduced simplicial sets
Our left transfer from Theorem 3.17 provides us with a combinatorial model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$, but not with explicit generating sets of cofibrations and acyclic cofibrations. As a first remedy, we produce an explicit generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$.
Proposition 4.7.
Let $I=\left\{i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]\mid n\geq 0\right\}$ denote the generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}$. Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets. The following statements are equivalent.
  1. (1)
    The map $f\colon X\rightarrow Y$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}$.
  2. (2)
    The map $f\colon X\rightarrow Y$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
  3. (3)
    The map $f\colon X\rightarrow Y$ has the right lifting property with respect to the set
    ${R}_{1}(I)=\left\{{R}_{1}(i_{n})\colon{R}_{1}(\partial\Delta[n])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 0\right\}.$
Hence, ${R}_{1}(I)$ is a generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$.
Proof.
Firstly, we show that (1) implies (2). If $f$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}$, then $f$ has the right lifting property with respect to all monomorphisms in ${\mathcal{SS}\mathrm{ets}}$. In particular, $f$ has the right lifting property with respect to all monomorphisms in ${\mathcal{SS}\mathrm{ets}}_{1}$, so $f$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
Then, (2) implies (3) as, by Lemma 3.14, ${R}_{1}(I)$ is a set of cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$.
Lastly, we show that (3) implies (1). Suppose $f$ has the right lifting property with respect to ${R}_{1}(I)$. To prove that $f$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}$, we solve, for $n\geq 0$, a lifting problem
${{\partial\Delta[n]}}$${X}$${{\Delta[n]}}$${{Y.}}$$\scriptstyle{i_{n}}$$\scriptstyle{f}$
Under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19, we equivalently solve a lifting problem
${{{R}_{1}(\partial\Delta[n])}}$${X}$${{{R}_{1}(\Delta[n])}}$${{Y.}}$$\scriptstyle{{R}_{1}(i_{n})}$$\scriptstyle{f}$
As ${R}_{1}(i_{n})$ is in ${R}_{1}(I)$ and $f$ has the right lifting property with respect to ${R}_{1}(I)$, a solution exists. ∎
Even though our treatment of Proposition 4.7 is biased towards the equivalence of (2) and (3), which grants us our generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$, we would be remiss if we neglected the equivalence of (1) and (2). The fact that (1) and (2) are equivalent means that ${\mathcal{SS}\mathrm{ets}}_{1}$ has no pathological acyclic fibrations. As we shall see later in this section, Proposition 4.7 finds no analog in the study of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$. The reason is that ${\mathcal{SS}\mathrm{ets}}_{1}$ has fibrations that are dissimilar to Kan fibrations, a phenomenon that we also investigate in this section. Thus, the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ is more than a restriction of the model category ${\mathcal{SS}\mathrm{ets}}$ to $1$-reduced simplicial sets; after all, the model category axioms are generally not amenable to such a restriction. Instead, the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ formalizes a different homotopy theory that is interesting in its own right.
4.3. Acyclic cofibrations of $1$-reduced simplicial sets
Our attempt to emulate Proposition 4.7 towards a generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$ encounters a problem: the $1$-reduction of a boundary inclusion is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, but the $1$-reduction of a horn inclusion need not be an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
Lemma 4.8.
For $n\geq 1$ and $0\leq k\leq n$, let $j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]$ denote the horn inclusion in ${\mathcal{SS}\mathrm{ets}}$. Then, ${R}_{1}\left(j_{n,k}\right)$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $n\neq 2$.
Proof.
The claim follows from Lemmas 3.13 and 3.14 and the definition of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$. ∎
Nevertheless, we can partially work our way around this obstacle. Let
$J=\left\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{ and}0\leq k\leq n\right\}$
denote the generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}$. Write
${R}_{1}(J)=\left\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 1\textrm{ and }0\leq k\leq n\right\},$
which is a set of cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ by Lemma 3.14, but not a set of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ by Lemma 4.8. To extract a candidate generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}$, we use Lemma 3.13 to strike out
  1. (1)
    the maps ${R}_{1}(j_{1,0})$ and ${R}_{1}(j_{1,1})$, which are the identity map of the point; and
  2. (2)
    the maps ${R}_{1}(j_{2,0})$, ${R}_{1}(j_{2,1})$, and ${R}_{1}(j_{2,2})$ which are not weak homotopy equivalences.
Therefore, we write
$\widetilde{J}=\left\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 3\textrm{ and }0\leq k\leq n\right\},$
whose elements are acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}$, in order to write
${R}_{1}(\widetilde{J})=\left\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\right\},$
which is a set of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$. Hence, having the right lifting property with respect to the set ${R}_{1}(\widetilde{J})$ is necessary for a map of $1$-reduced simplicial sets to be a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. We partially address the question of sufficiency. Namely, we establish that the set ${R}_{1}(\widetilde{J})$ detects fibrant objects in ${\mathcal{SS}\mathrm{ets}}_{1}$.
Proposition 4.9.
Let $X$ be a $1$-reduced simplicial set. The following statements are equivalent.
  1. (1)
    The $1$-reduced simplicial set $X$ is a Kan complex.
  2. (2)
    The $1$-reduced simplicial set $X$ is fibrant in ${\mathcal{SS}\mathrm{ets}}_{1}$.
  3. (3)
    The unique map $c\colon X\rightarrow\Delta[0]$ has the right lifting property with respect to the set
    ${R}_{1}(\widetilde{J})=\left\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\right\}.$
Hence, the set ${R}_{1}(\widetilde{J})$ detects fibrant objects in ${\mathcal{SS}\mathrm{ets}}_{1}$.
Proof.
Firstly, (1) and (2) are equivalent by the characterization of fibrant objects in ${\mathcal{SS}\mathrm{ets}}_{1}$ in Theorem 3.17.
If (2) holds, then $c\colon X\rightarrow\Delta[0]$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. Thus, $c\colon X\rightarrow\Delta[0]$ has the right lifting property with respect to the set ${R}_{1}(\widetilde{J})$, since every map in ${R}_{1}(\widetilde{J})$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. Therefore, (3) holds.
Lastly, suppose (3) holds. We prove that (1) holds. For $n\geq 1$ and $0\leq k\leq n$, a lifting problem
${{\Lambda[n,k]}}$${X}$${{\Delta[n]}}$$\scriptstyle{h}$$\scriptstyle{j_{n,k}}$$\scriptstyle{g}$
is equivalent, under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19, to a lifting problem
${{{R}_{1}\left(\Lambda[n,k]\right)}}$${X}$${{{R}_{1}\left(\Delta[n]\right).}}$$\scriptstyle{\eta}$$\scriptstyle{{R}_{1}\left(j_{n,k}\right)}$$\scriptstyle{\psi}$
In light of Lemma 3.13, we split our solution in three separate cases. If $n=1$, then Lemma 3.13(1) says that ${R}_{1}\left(j_{1,k}\right)$ is the identity map of the point, so our lifting problem can be solved by setting $\psi=\eta$. If $n\geq 3$, then our lifting problem can be solved, as we are assuming that $c\colon X\rightarrow\Delta[0]$ has the right lifting property with respect to the set ${R}_{1}(\widetilde{J})$. Lastly, if $n=2$, then Lemma 3.13(2) says that our lifting problem is
[…]
where $x_{0}$ denotes the unique 00-simplex of $X$, so we take $\psi$ to be the constant map at $x_{0}$. ∎
In fact, a different and more elaborate argument yields a stronger result.
Proposition 4.10 ([Qui69, Part II, Proposition 2.12]).
Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets. Suppose that $Y$ is a Kan complex. The following statements are equivalent.
  1. (1)
    The map $f\colon X\rightarrow Y$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
  2. (2)
    The map $f\colon X\rightarrow Y$ has the right lifting property with respect to the set
    ${R}_{1}(\widetilde{J})=\left\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\right\}.$
Hence, the set ${R}_{1}(\widetilde{J})$ detects fibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ whose target is a Kan complex.
The additional fibrancy condition on the target is crucially used in Quillen’s proof, which suggests that ${R}_{1}(\widetilde{J})$ is not a generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$. In his discussion of [Qui69, Part II, Proposition 2.12], Quillen affirms this suggestion and provides a hint towards showing that the fibrancy condition on the target in Proposition 4.10 is necessary. However, to the best of our knowledge, the question of whether ${R}_{1}(\widetilde{J})$ is a generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$ remains open. Our various attempts at producing a map in ${\mathcal{SS}\mathrm{ets}}_{1}$ that has the right lifting property with respect to ${R}_{1}(\widetilde{J})$ but is not a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ have been unsuccessful in the same fashion: our candidate maps are not fibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$, but they also fail to have the right lifting property with respect to ${R}_{1}(\widetilde{J})$. We pose the following conjecture to the community.
Conjecture 4.11.
Let ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$ denote the $1$-reduction functor. The set
${R}_{1}(\widetilde{J})=\left\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\right\}$
is not a generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$.
4.4. Fibrations of $1$-reduced simplicial sets
The issues that we ran into in our study of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ suggest that there exist fibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ that are dissimilar to Kan fibrations. To investigate this phenomenon, we characterize when a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ is a Kan fibration. In fact, we produce two results to that end, which help us find an explicit example of a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ that is not a Kan fibration.
To begin with, we note that every Kan fibration of $1$-reduced simplicial sets is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
Proposition 4.12.
If $f$ is a Kan fibration of $1$-reduced simplicial sets, then $f$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
Proof.
If $f$ is a Kan fibration, then $f$ has the right lifting property with respect to the class of acyclic cofibrations in $\mathcal{SS}\mathrm{ets}$, which contains the class of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$. Thus, $f$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. ∎
We prove a characterization that detects when a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ is a Kan fibration and helps us disprove the converse of Proposition 4.12. Our result is the $1$-reduced analog of [GJ99, Lemma 6.6]. Recall from Lemma 3.13 that, for $0\leq k\leq 2$, the map ${R}_{1}\left(j_{2,k}\right)\colon\Delta[0]\rightarrow S_{\mathrm{simp}}^{2}$ picks the unique 00-simplex of the simplicial $2$-sphere.
Proposition 4.13.
Let $f\colon X\rightarrow Y$ be a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. For $0\leq k\leq 2$, let $j_{2,k}\colon\Lambda[2,k]\rightarrow\Delta[2]$ denote the horn inclusion. The following statements are equivalent.
  1. (1)
    The map $f\colon X\rightarrow Y$ is a Kan fibration.
  2. (2)
    The map $f\colon X\rightarrow Y$ has the right lifting property with respect to ${R}_{1}\left(j_{2,k}\right)$ for $0\leq k\leq 2$.
Proof.
Under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19, a lifting problem of the type
[…]
is equivalent to a lifting problem of the type
(II) ${{\Lambda[2,k]}}$${X}$${{\Delta[2]}}$${{Y.}}$$\scriptstyle{j_{2,k}}$$\scriptstyle{f}$
If $f$ is a Kan fibration, then every lifting problem of type (II) has a solution. Thus, every lifting problem of type (I) has a solution, and $f$ has the right lifting property with respect to ${R}_{1}\left(j_{2,k}\right)$ for $0\leq k\leq 2$.
Conversely, if $f$ has the right lifting property with respect to ${R}_{1}\left(j_{2,k}\right)$ for $0\leq k\leq 2$, then every lifting problem of type (II) has a solution. If $n\neq 2$ and $0\leq t\leq n$, then solving a lifting problem
${{\Lambda[n,t]}}$${X}$${{\Delta[n]}}$${{Y}}$$\scriptstyle{j_{n,t}}$$\scriptstyle{f}$
is equivalent, under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19, to solving a lifting problem
${{{R}_{1}\left(\Lambda[n,t]\right)}}$${X}$${{{R}_{1}\left(\Delta[n]\right)}}$${{Y.}}$$\scriptstyle{{R}_{1}\left(j_{n,t}\right)}$$\scriptstyle{f}$
By Lemma 4.8, ${R}_{1}\left(j_{n,t}\right)$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ because $n\neq 2$, so our lifting problem has a solution because $f$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. Overall, we conclude that $f$ is a Kan fibration. ∎
Proposition 4.13 helps us find a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ that bears no resemblance to a Kan fibration. To do so, we need an auxiliary lemma about maps from a $1$-reduced simplicial set to the simplicial $2$-sphere $S_{\mathrm{simp}}^{2}$.
Lemma 4.14.
Let $B$ be a $1$-reduced simplicial set. Then, a map $h\colon B\rightarrow S_{\mathrm{simp}}^{2}$ is constant if and only if
$h_{\ast}\colon\pi_{2}(B)\rightarrow\pi_{2}(S_{\mathrm{simp}}^{2})$
is not an epimorphism.
Proof.
If $h$ is constant, then so is $h_{\ast}$. If $h$ is not constant, then there is a $2$-simplex $\omega$ of $B$ such that $h(\omega)$ is the unique non-degenerate $2$-simplex $\epsilon$ of $S_{\mathrm{simp}}^{2}$. As $B$ is $1$-reduced, Corollary 2.20 allows us to view $\omega$ as a map
$\omega\colon S_{\mathrm{simp}}^{2}\rightarrow B.$
Then, we compute that
$h_{\ast}\left([\omega]\right)=[h(\omega)]=[\epsilon],$
which is a generator of $\pi_{2}(S_{\mathrm{simp}}^{2})\cong\mathbb{Z}$. Thus, $h_{\ast}$ is an epimorphism. ∎
Proposition 4.15.
For $0\leq k\leq 2$, the map of $1$-reduced simplicial sets ${R}_{1}\left(j_{2,k}\right)\colon\Delta[0]\rightarrow S_{\mathrm{simp}}^{2}$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ but not a Kan fibration.
Proof.
We first show that ${R}_{1}\left(j_{2,k}\right)$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. For an acyclic cofibration $\gamma\colon A\rightarrow B$ in ${\mathcal{SS}\mathrm{ets}}_{1}$ and a lifting problem
${A}$${{\Delta[0]}}$${B}$${{S_{\mathrm{simp}}^{2},}}$$\scriptstyle{c}$$\scriptstyle{\gamma}$$\scriptstyle{{R}_{1}\left(j_{2,k}\right)}$$\scriptstyle{g}$$\scriptstyle{h}$
we claim that the unique option for $g$ solves our lifting problem. The upper left triangle has to commute. To show that the lower right triangle also commutes, we show that $h$ is constant. By Lemma 4.14, we show that
$h_{\ast}\colon\pi_{2}(B)\rightarrow\pi_{2}(S_{\mathrm{simp}}^{2})$
is not an epimorphism; in fact, we show that $h_{\ast}$ is constant. Firstly, as $c_{\ast}$ is constant, so is
$h_{\ast}\gamma_{\ast}=(h\gamma)_{\ast}=({{R}_{1}\left(j_{2,k}\right)}c)_{\ast}={{R}_{1}\left(j_{2,k}\right)}_{\ast}c_{\ast}.$
Because $\gamma$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, we know that $\gamma_{\ast}$ is an isomorphism. Therefore, $h_{\ast}$ is constant.
Now, we apply the characterization from Proposition 4.13 to argue that ${R}_{1}\left(j_{2,k}\right)$ is not a Kan fibration because it does not have the right lifting property with respect to itself. Indeed, the lifting problem
[…]
has no solution because the unique option for $g$ does not make the lower right triangle commute. ∎
In the presence of a fibrancy condition on the target, Proposition 4.13 can be strengthened. Namely, we establish the following characterization result, which is the $1$-reduced analog of [GJ99, Corollary 6.9].
Proposition 4.16.
Let $f\colon X\rightarrow Y$ be a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. Suppose that $Y$ is a Kan complex. Then, $X$ is also a Kan complex, and the following statements are equivalent.
  1. (1)
    The map $f\colon X\rightarrow Y$ is a Kan fibration.
  2. (2)
    The map $f\colon X\rightarrow Y$ has the right lifting property with respect to ${R}_{1}\left(j_{2,k}\right)$ for $0\leq k\leq 2$.
  3. (3)
    The map $f\colon X\rightarrow Y$ induces an epimorphism $f_{\ast}\colon\pi_{2}(X)\rightarrow\pi_{2}(Y)$.
In our proof, we use the construction of the homotopy groups of Kan complexes from [GJ99, Chapter I.7].
Proof.
[…]
Now, Proposition 4.13 says that (1) and (2) are equivalent. We show that (2) and (3) are equivalent. Firstly, suppose that $f$ has the right lifting property with respect to ${R}_{1}\left(j_{2,k}\right)$ for $0\leq k\leq 2$. As $Y$ is a Kan complex, every element of $\pi_{2}(Y)$ is of the form $[\omega]$ for a map
$\omega\colon S_{\mathrm{simp}}^{2}\rightarrow Y.$
We solve the lifting problem
[…]
to produce an element $[\alpha]\in\pi_{2}(X)$ such that
$f_{\ast}([\alpha])=[f\alpha]=[\omega]\in\pi_{2}(Y).$
Conversely, suppose that $f$ induces an epimorphism at $\pi_{2}$. For $0\leq k\leq 2$, we solve a lifting problem
[…]
As $f_{\ast}$ is an epimorphism, there is a commutative diagram
[…]
where
$f_{\ast}([\beta])=[f\beta]=[\eta]\in\pi_{2}(Y)$
by means of the pointed homotopy
$H\colon S_{\mathrm{simp}}^{2}\wedge\Delta[1]_{+}\rightarrow Y$
starting from $Hd^{1}=f\beta$ and ending at $Hd^{0}=\eta$. By Lemma 2.15, $S_{\mathrm{simp}}^{2}\wedge\Delta[1]_{+}$ is $1$-reduced. As $d^{1}$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ and $f$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, there exists a lift
[…]
Then, $\psi=gd^{0}$ solves our lifting problem because
$f\psi=fgd^{0}=Hd^{0}=\eta.$
The fibrancy condition in Proposition 4.16 is the reason why we could not have used it for our example of a pathological fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ from Proposition 4.15. Indeed, we conclude this section by showing that, for $n\geq 2$, the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ is not a Kan complex. To that end, we need a counting lemma.
Lemma 4.17.
Let $x$ be an $n$-simplex of a simplicial set. For $0\leq k\leq n$, let $s_{k}(x)$ be the $k$-th degeneracy of $x$.
  1. (1)
    If $x$ is a non-degenerate $n$-simplex, then $s_{k}(x)$ has exactly $2$ non-degenerate faces.
  2. (2)
    If $x$ is a degenerate $n$-simplex, then $s_{k}(x)$ has no non-degenerate faces.
Hence, in either case, $s_{k}(x)$ has at most $2$ non-degenerate faces.
Proof.
For $10\leq i\leq n+$ 1, let $d_{i}s_{k}(x)$ be the $i$-th face of $s_{k}(x)$. The result follows from the simplicial identities:
$d_{i}s_{k}(x)=\begin{cases}s_{k-1}d_{i}(x),\textrm{ if }0\leq i\leq k-1;\\ x,\textrm{ if }k\leq i\leq k+1;\\ s_{k}d_{i-1}(x),\textrm{ if }k+2\leq i\leq n+1.\end{cases}$
Proposition 4.18.
For $n\geq 2$, the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ is not a Kan complex, thus not fibrant in ${\mathcal{SS}\mathrm{ets}}_{1}$.
Proof.
Write $\epsilon$ for the unique non-degenerate $n$-simplex of $S_{\mathrm{simp}}^{n}$. For $10\leq k\leq n+$ 1, consider the lifting problem
${{\Lambda[n+1,k]}}$${{S_{\mathrm{simp}}^{n}}}$${{\Delta[n+1],}}$$\scriptstyle{h}$$\scriptstyle{j_{n+1,k}}$$\scriptstyle{g}$
where $h$ is the horn that picks out the $n$-simplices $\left(\epsilon,\dots,\epsilon\right)$. Any simplex $g$ that fills in this horn would be a degenerate $(n+1)$-simplex with $n+1$ non-degenerate faces, which is impossible for $n\geq 2$ by Lemma 4.17. ∎
5. The simplicial model structure on $1$-reduced simplicial sets
A convenient feature of the model categories ${\mathcal{SS}\mathrm{ets}}$ and ${\mathcal{SS}\mathrm{ets}}_{\ast}$ is that they are simplicial, so they carry more structure for us to use. This feature motivates us to endow ${\mathcal{SS}\mathrm{ets}}_{1}$ with a simplicial model structure. We modify the simplicial model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ to define function complexes, tensors, and cotensors of $1$-reduced simplicial sets. Then, we verify that our definitions endow ${\mathcal{SS}\mathrm{ets}}_{1}$ with a simplicial model structure.
5.1. Simplicial model categories
A reference on simplicial model categories is [Hir03, Chapter 9].
Definition 5.1 ([Hir03, Definition 9.1.2]).
Let $\mathcal{M}$ be a model category that is enriched over simplicial sets; in particular, for two objects $X$ and $Y$ of $\mathcal{M}$, we have a function complex $\operatorname{Map}_{\mathcal{M}}\left(X,Y\right)$. Then, $\mathcal{M}$ is a simplicial model category if $\mathcal{M}$ satisfies the following axioms.
  1. (1)
    (Adjointness axiom) For two objects $X$ and $Y$ of $\mathcal{M}$ and a simplicial set $Z$, there is a tensor object $X\otimes Z$ of $\mathcal{M}$ and a cotensor object $Y^{Z}$ of $\mathcal{M}$ such that there are isomorphisms of simplicial sets
    $\operatorname{Map}_{\mathcal{M}}\left(X\otimes Z,Y\right)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(Z,\operatorname{Map}_{\mathcal{M}}\left(X,Y\right)\right)\cong\operatorname{Map}_{\mathcal{M}}\left(X,Y^{Z}\right)$
    that are natural in $X$, $Y$, and $Z$. Note that $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(Z,\operatorname{Map}_{\mathcal{M}}\left(X,Y\right)\right)$ is a function complex of simplicial sets; see Example 5.2.
  2. (2)
    (Homotopy lifting-extension axiom) If $i\colon A\rightarrow B$ is a cofibration in $\mathcal{M}$ and $p\colon X\rightarrow Y$ is a fibration in $\mathcal{M}$, then the induced map
    $\operatorname{Map}_{\mathcal{M}}\left(B,X\right)\xrightarrow{(i^{\ast},p_{\ast})}\operatorname{Map}_{\mathcal{M}}\left(A,X\right)\underset{\operatorname{Map}_{\mathcal{M}}\left(A,Y\right)}{\times}\operatorname{Map}_{\mathcal{M}}\left(B,Y\right)$
    is a Kan fibration, which is also a weak homotopy equivalence if $i$ or $p$ is also a weak equivalence in $\mathcal{M}$.
Example 5.2 ([Hir03, Example 9.1.13]).
The category ${\mathcal{SS}\mathrm{ets}}$ is a simplicial model category. Given simplicial sets $X$, $Y$, and $Z$, we define
  1. (1)
    the function complex $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(X,Y\right)$ to be the simplicial set with $n$-simplices $\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}}\left(X\times\Delta[n],Y\right)$;
  2. (2)
    the tensor of $X$ with $Z$ to be the simplicial set $X\times Z$; and
  3. (3)
    the cotensor of $X$ with $Z$ to be the simplicial set $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(Z,X\right)$.
Example 5.3 ([Hir03, Example 9.1.14]).
The category ${\mathcal{SS}\mathrm{ets}}_{\ast}$ is a simplicial model category. For two pointed simplicial sets $X$ and $Y$ and a simplicial set $Z$, we define
  1. (1)
    the function complex $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X,Y\right)$ as the simplicial set with $n$-simplices $\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge\Delta[n]_{+},Y\right)$;
  2. (2)
    the tensor of $X$ with $Z$ to be the pointed simplicial set $X\times Z_{+}$; and
  3. (3)
    the cotensor of $X$ with $Z$ to be the pointed simplicial set $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(Z_{+},X\right)$.
5.2. Function complexes, tensors, and cotensors of $1$-reduced simplicial sets
We modify the simplicial model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ to define function complexes, tensors, and cotensors in ${\mathcal{SS}\mathrm{ets}}_{1}$. We define
  1. (1)
    the function complex of two $1$-reduced simplicial sets $X$ and $Y$ to be $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)$;
  2. (2)
    the tensor of a $1$-reduced simplicial set $X$ with a simplicial set $Z$ to be $X\wedge Z_{+}$; and
  3. (3)
    the cotensor of a $1$-reduced simplicial set $X$ with a simplicial set $Z$ to be $E_{1}\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},X)\right)$.
5.3. An exponential law
The simplicial model structure on ${\mathcal{SS}\mathrm{ets}}$ is underpinned by the exponential law for simplicial sets [GJ99, Chapter I, Proposition 5.1]. We formulate an exponential law for pointed simplicial sets for our simplicial model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$. For two pointed simplicial sets $X$ and $Y$, the evaluation
$\operatorname{ev}\colon X\wedge\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\rightarrow Y$
is defined at the level of $n$-simplices by
$\operatorname{ev}_{n}([x,f])=f_{n}([x,\sigma_{n}]),$
where $\sigma_{n}$ is the unique non-degenerate $n$-simplex of $\Delta[n]$. We establish the following exponential law.
Proposition 5.4.
For pointed simplicial sets $X$, $Y$, and $Z$, the map
$\operatorname{ev}_{\ast}\colon\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(Z,\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right)\rightarrow\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge Z,Y\right)$
sending $f\colon Z\rightarrow\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)$ to the composite
$X\wedge Z\xrightarrow{\operatorname{id}_{X}\wedge f}X\wedge\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\xrightarrow{\operatorname{ev}}Y$
is a bijection that is natural in $X$, $Y$, and $Z$.
Remark 5.5.
This natural bijection amounts to the adjunction $X\wedge-\colon{\mathcal{SS}\mathrm{ets}}_{\ast}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,-)$.
Proof.
The inverse of $\operatorname{ev}_{\ast}$ sends a map $g\colon X\wedge Z\rightarrow Y$ to the map
$\hat{g}\colon Z\rightarrow\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)$
defined by sending an $n$-simplex $z$ of $Z$, which we view as a map $z\colon\Delta[n]\rightarrow Z$, to the composite
[…]
where, since $Z$ is pointed, we use Lemma 2.8 to induce the pointed map $\widetilde{z}\colon\Delta[n]_{+}\rightarrow Z$. ∎
In particular, we use the following consequence of our exponential law.
Corollary 5.6.
For a $1$-reduced simplicial set $X$, there is an adjunction
$X\wedge-\colon{\mathcal{SS}\mathrm{ets}}_{\ast}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,-).$
Proof.
By Lemma 2.15, the functor $X\wedge-$ maps into ${\mathcal{SS}\mathrm{ets}}_{1}$. The adjunction now follows from Remark 5.5. ∎
5.4. Verification of the adjointness axiom
We use our exponential law to verify that the structure defined in Section 5.2 satisfies the adjointness axiom. We need an auxiliary lemma that is straightforward to check.
Lemma 5.7.
Given two simplicial sets $Z$ and $W$, there is an isomorphism of pointed simplicial sets
$\left(Z_{+}\right)\wedge\left(W_{+}\right)\cong\left(Z\times W\right)_{+}$
that is natural in $Z$ and $W$.
We are ready to verify the adjointness axiom for ${\mathcal{SS}\mathrm{ets}}_{1}$.
Proposition 5.8.
Given two $1$-reduced simplicial sets $X$ and $Y$ and a simplicial set $Z$, there are isomorphisms of simplicial sets
$\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge Z_{+},Y\right)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(Z,\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X,E\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)\right)\right)$
that are natural in $X$, $Y$, and $Z$.
Proof.
To prove $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge Z_{+},Y\right)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(Z,\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right)$, we compute for arbitrary $n$ that
$\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(\left(X\wedge Z_{+}\right)\wedge\Delta[n]_{+},Y\right)$ $\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge\l\left(Z_{+}\wedge\Delta[n]_{+}\right),Y\right)$
$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(Z_{+}\wedge\Delta[n]_{+},\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right)$
$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(\left(Z\times\Delta[n]\right)_{+},\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right)$
$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}}\left(Z\times\Delta[n],\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right).$
The second bijection uses Proposition 5.4, the third one uses Lemma 5.7, and the fourth one uses Lemma 2.8.
To prove $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge Z_{+},Y\right)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X,E\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)\right)\right)$, we compute for arbitrary $n$ that
$\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(\left(X\wedge Z_{+}\right)\wedge\Delta[n]_{+},Y\right)$ $\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(Z_{+}\wedge\left(X\wedge\Delta[n]_{+}\right),Y\right)$
$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge\Delta[n]_{+},\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)\right)$
$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{1}}\left(X\wedge\Delta[n]_{+},E\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)\right)\right)$
$=\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge\Delta[n]_{+},E\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)\right)\right).$
The second bijection comes from Proposition 5.4, and the third bijection is a consequence of Lemma 2.23(2). ∎
5.5. Saturated classes
To verify the homotopy lifting-extension axiom, we adapt the proof of the homotopy lifting-extension axiom for simplicial sets in [GJ99, Chapter I, Proposition 5.2]. The proof for simplicial sets uses saturated classes of monomorphisms of simplicial sets, which are defined as follows.
Definition 5.9 ([GZ67, Chapter IV, Definition 2.1]).
A class $S$ of monomorphisms of simplicial sets is saturated if the following conditions are met.
  1. (1)
    All isomorphisms are contained in $S$.
  2. (2)
    In a pushout square
[…]
    if $m$ is contained in $S$, then $m^{\prime}$ is contained in $S$.
  3. (3)
    In a section-retraction commutative diagram
[…]
    if $m$ is contained in $S$, then $m^{\prime}$ is contained in $S$.
  4. (4)
    Given a sequence of composable maps
    $A_{1}\xrightarrow{m_{1}}A_{2}\xrightarrow{m_{2}}\cdots,$
    if $m_{t}$ is contained in $S$ for every positive integer $t$, then so is $A_{1}\rightarrow\operatorname{colim}(A_{n})$.
  5. (5)
    Given an indexing set $L$ and $m_{l}\colon A_{l}\rightarrow B_{l}$ for $l\in L$, if $m_{l}$ is contained in $S$ for every $l\in L$, then
    $\coprod\limits_{l\in L}m_{l}\colon\coprod\limits_{l\in L}A_{l}\rightarrow\coprod\limits_{l\in L}B_{l}$
    is contained in $S$.
As the intersection of saturated classes forms another saturated class, the following definition is well-posed.
Definition 5.10.
The saturation of a class $P$ of monomorphisms of simplicial sets is
$\langle P\rangle=\bigcap\limits_{\begin{subarray}{c}S\supset P\textrm{ and}\\ S\textrm{ is saturated }\end{subarray}}S.$
Saturated classes provide us with descriptions of the class of acyclic cofibrations of simplicial sets.
Proposition 5.11 ([GJ99, Chapter I, Proposition 4.2]).
Consider the monomorphism classes of simplicial sets
  1. (1)
    $P_{1}=\left\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{and }0\leq k\leq n\right\}$; and
  2. (2)
    $P_{2}=\left\{\left(\Delta[1]\times\partial\Delta[n]\right)\cup\left(\{e\}\times\Delta[n]\right)\rightarrow\Delta[1]\times\Delta[n]\mid e\in\{0,1\}\textrm{ and }n\geq 0\right\}$.
Then, $\langle P_{1}\rangle=\langle P_{2}\rangle$ is the class of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}$.
Attaching a disjoint basepoint yields a description of the class of acyclic cofibrations of pointed simplicial sets.
Corollary 5.12.
Consider the monomorphism classes of pointed simplicial sets
  1. (1)
    $(P_{1})_{+}=\left\{(j_{n,k})_{+}\colon\Lambda[n,k]_{+}\rightarrow\Delta[n]_{+}\mid n\geq 1\textrm{ and }0\leq k\leq n\right\}$; and
  2. (2)
    $(P_{2})_{+}=\left\{\left(\left(\Delta[1]\times\partial\Delta[n]\right)\cup\left(\{e\}\times\Delta[n]\right)\right)_{+}\rightarrow\left(\Delta[1]\times\Delta[n]\right)_{+}\mid e\in\{0,1\}\textrm{ and }n\geq 0\right\}$.
Then, $\langle(P_{1})_{+}\rangle=\langle(P_{2})_{+}\rangle$ is the class of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{\ast}$.
We use Corollary 5.12 to prove a lemma that underpins our verification of the homotopy lifting-extension axiom.
Lemma 5.13.
For two pointed simplicial maps $i\colon A\rightarrow B$ and $j\colon X\rightarrow Y$, if both of them are monomorphisms, then the induced map
$\left(A\wedge Y\right)\underset{\left(A\wedge X\right)}{\bigcup}\left(B\wedge X\right)\rightarrow B\wedge Y$
is a monomorphism, which is also a weak homotopy equivalence if $i$ or $j$ is also a weak homotopy equivalence.
Proof.
If both $i$ and $j$ are monomorphisms, then so is the induced map. To prove the second part of the claim, the key idea is to write $P$ for the class of monomorphisms $A^{\prime}\rightarrow B^{\prime}$ of pointed simplicial sets such that
$\left(A^{\prime}\wedge Y\right)\underset{\left(A\wedge X\right)}{\bigcup}\left(B^{\prime}\wedge X\right)\rightarrow B^{\prime}\wedge Y$
is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{\ast}$. We show that $P$ contains the class of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{\ast}$. As $P$ is a saturated class, thus equal to its own saturation, Corollary 5.12 truncates our task to showing that
$(P_{2})_{+}\subset P.$
Thus, for $e\in\{0,1\}$ and $n\geq 0$, we need to show that
$\left(\left(\left(\Delta[1]\times\partial\Delta[n]\right)\cup\left(\{e\}\times\Delta[n]\right)\right)_{+}\wedge Y\right)\cup\left(\left(\Delta[1]\times\Delta[n]\right)_{+}\wedge X\right)\rightarrow\left(\Delta[1]\times\Delta[n]\right)_{+}\wedge Y$
is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{\ast}$. We use Lemma 5.7 to write down the commutative diagram
[…]
where the upper pair of vertical isomorphisms is an instance of Lemma 5.7 and
$\left(\Delta[1]_{+}\wedge\left(\left(\partial\Delta[n]_{+}\wedge Y\right)\cup\left(\Delta[n]_{+}\wedge X\right)\right)\right)\cup\left(\{e\}_{+}\wedge\left(\Delta[n]_{+}\wedge Y\right)\right)\rightarrow\Delta[1]_{+}\wedge\left(\Delta[n]_{+}\wedge Y\right)$
is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{\ast}$ because its geometric realization is the inclusion of a strong deformation retract in pointed topological spaces. Overall, we conclude that $(P_{2})_{+}\subset P$. ∎
5.6. Verification of the homotopy lifting-extension axiom
We are now ready to verify that the structure defined in Section 5.2 also satisfies the homotopy lifting-extension axiom.
Proposition 5.14.
If $i\colon A\rightarrow B$ is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ and $p\colon X\rightarrow Y$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, then the induced map
$\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,X\right)\xrightarrow{(i^{\ast},p_{\ast})}\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,X\right)\underset{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,Y\right)}{\times}\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,Y\right)$
is a Kan fibration, which is also a weak homotopy equivalence if $i$ or $p$ is also a weak homotopy equivalence.
Proof.
To prove that the induced map is a Kan fibration, we solve a lifting problem of the form
${{\Lambda[n,k]}}$${{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,X\right)}}$${{\Delta[n]}}$${{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,X\right)\underset{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,Y\right)}{\times}\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,Y\right).}}$$\scriptstyle{j_{n,k}}$$\scriptstyle{(i^{\ast},p_{\ast})}$
The adjunction $(-)_{+}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon V$ from Lemma 2.8 yields the equivalent lifting problem
[…]
We apply Corollary 5.6 to infer that this new lifting problem is, in turn, equivalent to the lifting problem
${{\left(\Lambda[n,k]_{+}\wedge B\right)\underset{\left(\Lambda[n,k]_{+}\wedge A\right)}{\bigcup}\left(\Delta[n]_{+}\wedge A\right)}}$${X}$${{\Delta[n]_{+}\wedge B}}$${{Y.}}$$\scriptstyle{p}$
By Lemma 5.13, the left-hand map is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ as $\left(j_{n,k}\right)_{+}$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{\ast}$. As $p$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, our lifting problem can be solved, so the induced map is a Kan fibration.
If $i$ or $p$ is also a weak homotopy equivalence, then we prove that the induced map is also a weak homotopy equivalence. To that end, we solve a lifting problem
${{\partial\Delta[n]}}$${{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,X\right)}}$${{\Delta[n]}}$${{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,X\right)\underset{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,Y\right)}{\times}\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,Y\right).}}$$\scriptstyle{i_{n}}$$\scriptstyle{(i^{\ast},p_{\ast})}$
Another consecutive application of the adjunctions from Lemma 2.8 and Corollary 5.6 translates our lifting problem to the equivalent lifting problem
${{\left(\partial\Delta[n]_{+}\wedge B\right)\underset{\left(\partial\Delta[n]_{+}\wedge A\right)}{\bigcup}\left(\Delta[n]_{+}\wedge A\right)}}$${X}$${{\Delta[n]_{+}\wedge B}}$${{Y.}}$$\scriptstyle{p}$
By Lemma 5.13, the left-hand map is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ because $\left(i_{n}\right)_{+}$ is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{\ast}$. Depending on whether $i$ or $p$ is also a weak homotopy equivalence, we have two cases to scrutinize.
  1. (1)
    If $i$ is also a weak homotopy equivalence, then, by Lemma 5.13, the left-hand map is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. As $p$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, our lifting problem can be solved.
  2. (2)
    If $p$ is also a weak homotopy equivalence, then $p$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$. As the left-hand map is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, our lifting problem can be solved.
In either case, the induced map is both a Kan fibration and a weak homotopy equivalence, as required. ∎
By Propositions 5.8 and 5.14, our definitions in Section 5.2 endow ${\mathcal{SS}\mathrm{ets}}_{1}$ with a simplicial model structure.
Theorem 5.15.
Let ${\mathcal{SS}\mathrm{ets}}_{1}$ denote the model category from Theorem 3.17. We define
  1. (1)
    the function complex of two $1$-reduced simplicial sets $X$ and $Y$ to be $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)$;
  2. (2)
    the tensor of a $1$-reduced simplicial set $X$ with a simplicial set $Z$ to be $X\wedge Z_{+}$; and
  3. (3)
    the cotensor of a $1$-reduced simplicial set $X$ with a simplicial set $Z$ to be $E_{1}\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},X)\right)$.
With these definitions, ${\mathcal{SS}\mathrm{ets}}_{1}$ has the structure of a simplicial model category.
Proof.
With our definition of function complexes, ${\mathcal{SS}\mathrm{ets}}_{1}$ is enriched over simplicial sets as ${\mathcal{SS}\mathrm{ets}}_{\ast}$ is enriched over simplicial sets. By Propositions 5.8 and 5.14, our definitions satisfy the simplicial model category axioms. ∎
6. Rational homotopy theory of $1$-connected spaces
From now on, let $M$ be a multiplicative subset of $\mathbb{Z}$. We gather the background and auxiliary results from the rational homotopy theory of $1$-connected topological spaces that we need to recover the entirety of Quillen’s family of model structures on ${\mathcal{SS}\mathrm{ets}}_{1}$. In particular, we characterize $M$-local spaces and $M$-local homotopy equivalences in terms of function complexes, which prepares us for the left Bousfield localization in Section 7. Even though we do not expect these characterizations to be novel to the rational homotopy theory expert, we were unable to find them in the literature, so we provide a full proof.
6.1. Localization of $1$-connected spaces
We first review some concepts from [FHT01, Chapter 9].
Definition 6.1.
Let $M$ be a multiplicative subset of $\mathbb{Z}$.
  1. (1)
    A map of $1$-connected spaces $f\colon X\rightarrow Y$ is an $M$-local homotopy equivalence if, for every $n\geq 2$, the induced map
    $f_{\ast}\underset{\mathbb{Z}}{\otimes}\operatorname{id}_{M^{-1}\mathbb{Z}}\colon\pi_{n}(X)\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}\rightarrow\pi_{n}(Y)\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}$
    is an isomorphism.
  2. (2)
    A $1$-connected space $Y$ is $M$-local if, for every $n\geq 2$, the abelian group $\pi_{n}(Y)$ is an $M^{-1}\mathbb{Z}$-module.
  3. (3)
    An $M$-localization of a $1$-connected space $Y$ is an $M$-local homotopy equivalence $f\colon Y\rightarrow Y_{M^{-1}\mathbb{Z}}$ with $M$-local target.
When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we speak of a rational homotopy equivalence, a rational space, and a rationalization.
Remark 6.2.
If $f\colon Y\rightarrow Y_{\mathbb{Q}}$ is a rationalization of $Y$, then, for every $n\geq 2$, $f$ induces the isomorphism
$f_{\ast}\underset{\mathbb{Z}}{\otimes}\operatorname{id}_{\mathbb{Q}}\colon\pi_{n}(Y)\underset{\mathbb{Z}}{\otimes}\mathbb{Q}\xrightarrow{\cong}\pi_{n}(Y_{\mathbb{Q}})$
that discards all torsion in the $n$-th homotopy group of $Y$ in a purely topological fashion.
Example 6.3.
Every weak homotopy equivalence of $1$-connected spaces is an $M$-local homotopy equivalence. Also, every $M$-local homotopy equivalence of $M$-local spaces is a weak homotopy equivalence.
Example 6.4.
For $n\geq 2$, the $M$-local $n$-sphere $S^{n}_{M^{-1}\mathbb{Z}}$ is an $M$-local CW complex containing the $n$-sphere $S^{n}$ as a CW subcomplex, and the inclusion $S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$ is an $M$-localization of $S^{n}$.
Every $1$-connected space can be $M$-localized uniquely up to homotopy equivalence as follows.
Theorem 6.5 ([FHT01, Theorem 9.7]).
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $Y$ be a $1$-connected space.
  1. (1)
    There exists a relative CW complex $\left(Y_{M^{-1}\mathbb{Z}},Y\right)$ with no 00-cells and no $1$-cells such that the inclusion $Y\rightarrow Y_{M^{-1}\mathbb{Z}}$ is an $M$-localization of $Y$.
  2. (2)
    Every map $f\colon Y\rightarrow Z$ with $M$-local target extends to a map $g\colon Y_{M^{-1}\mathbb{Z}}\rightarrow Z$ such that the diagram
[…]
    commutes, and every homotopy $f\simeq f^{\prime}$ extends to a homotopy $g\simeq g^{\prime}$. Thus, every map $f\colon Y\rightarrow Z$ with $M$-local target extends to a map $g\colon Y_{M^{-1}\mathbb{Z}}\rightarrow Z$ uniquely up to homotopy.
In particular, any other space $Y^{\prime}_{M^{-1}\mathbb{Z}}$ satisfying (1) and (2) is homotopy equivalent (relative to $Y$) to $Y_{M^{-1}\mathbb{Z}}$.
Remark 6.6.
Note that, if $Y$ is a $1$-connected CW complex, then so is $Y_{M^{-1}\mathbb{Z}}$.
6.2. A characterization of local spaces
For a multiplicative subset $M$ of $\mathbb{Z}$, we characterize $M$-local spaces in a way that supports our model-categorical arguments for the left Bousfield localization in Section 7.
For two topological spaces $X$ and $Y$, the function complex
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Y\right)$
is the simplicial set with $n$-simplices
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Y\right)_{n}=\operatorname{Hom}_{\mathcal{T}\mathrm{op}}\left(X\times|\Delta[n]|,Y\right).$
In particular, observe that
$\pi_{0}\left(\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Y\right)\right)=\left[X,Y\right]$
is the set of homotopy classes of maps from $X$ to $Y$.
The following lemma is a straightforward application of the universal property of $M$-localization.
Lemma 6.7.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $X$ and $Y$ be $1$-connected spaces. If $Y$ is $M$-local, then the map
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Y\right)$
induced by the inclusion $X\rightarrow X_{M^{-1}\mathbb{Z}}$ is a weak homotopy equivalence.
We are now ready to establish a characterization of $M$-local spaces in the language of function complexes.
Proposition 6.8.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A $1$-connected space $Y$ is $M$-local if and only if, for every $n\geq 2$, the map
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n},Y\right)$
induced by the inclusion $S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$ is a weak homotopy equivalence.
Proof.
If $Y$ is $M$-local, then Lemma 6.7 implies that, for $n\geq 2$, the map
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n},Y\right)$
is a weak homotopy equivalence. Conversely, suppose that, for $n\geq 2$, the map
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n},Y\right)$
is a weak homotopy equivalence. Let $\tau\colon S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$ denote the inclusion. Then, for every map $f\colon S^{n}\rightarrow Y$, there exists a map $g\colon S^{n}_{M^{-1}\mathbb{Z}}\rightarrow Y$, unique up to homotopy, such that the diagram
[…]
commutes up to homotopy. For $n\geq 2$, let $[f]$ be an element of $\pi_{n}(Y)$ represented by a map $f\colon S^{n}\rightarrow Y$, and let $m\in M$. Then, there exists a map $g\colon S^{n}_{M^{-1}\mathbb{Z}}\rightarrow Y$, unique up to homotopy, such that the diagram
[…]
commutes up to homotopy. In particular, homotopy commutativity implies that
$[f]=f_{\ast}\left([\operatorname{id}_{S^{n}}]\right)=g_{\ast}\tau_{\ast}\left([\operatorname{id}_{S^{n}}]\right)=g_{\ast}\left([\tau]\right)\in\pi_{n}(Y).$
As $S^{n}_{M^{-1}\mathbb{Z}}$ is $M$-local, there exists a unique element
$\dfrac{1}{m}\cdot[\tau]\in\pi_{n}(S^{n}_{M^{-1}\mathbb{Z}})\cong\pi_{n}(S^{n})\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}\cong\mathbb{Z}\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}\cong M^{-1}\mathbb{Z}.$
Therefore, we uniquely define
$\dfrac{1}{m}\cdot[f]=g_{\ast}\left(\dfrac{1}{m}\cdot[\tau]\right)\in\pi_{n}(Y).$
We conclude that the $1$-connected space $Y$ is $M$-local. ∎
As our function complexes of $1$-reduced simplicial sets are pointed, we need the pointed version of Proposition 6.8. We write $\mathcal{T}\mathrm{op}_{\ast}$ for the category of pointed spaces. For two pointed spaces $X$ and $Y$, the function complex
$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(X,Y\right)$
is the simplicial set with $n$-simplices
$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(X,Y\right)_{n}=\operatorname{Hom}_{\mathcal{T}\mathrm{op}_{\ast}}\left(X\wedge|\Delta[n]|_{+},Y\right).$
Observe that
$\pi_{0}(\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(X,Y\right))=\left[X,Y\right]_{\ast}$
is the set of pointed homotopy classes of pointed maps from $X$ to $Y$.
In terms of these function complexes, we state the pointed version of Proposition 6.8, which is proven identically.
Proposition 6.9.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A $1$-connected pointed space $Y$ is $M$-local if and only if, for every $n\geq 2$, the map
$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(S^{n},Y\right)$
induced by the inclusion $S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$ is a weak homotopy equivalence.
6.3. A characterization of local homotopy equivalences
For a multiplicative subset $M$ of $\mathbb{Z}$, we also characterize the $M$-local homotopy equivalences in preparation for the left Bousfield localization in Section 7.
The following lemma, in conjunction with its pointed analog, underpins our arguments.
Lemma 6.10 ([Hir03, Corollary 9.3.3(2)]).
  1. (1)
    If $f\colon X\rightarrow Y$ is a weak homotopy equivalence of simplicial sets and $Z$ is a Kan complex, then
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(Y,Z)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(X,Z)$
    is a weak homotopy equivalence.
  2. (2)
    If $f\colon X\rightarrow Y$ is a weak homotopy equivalence of CW complexes and $Z$ is a topological space, then
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y,Z)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,Z)$
    is a weak homotopy equivalence.
We provide a series of characterizations of $M$-local homotopy equivalences in terms of function complexes.
Proposition 6.11.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $f\colon X\rightarrow Y$ be a map of $1$-connected CW complexes. The following statements are equivalent.
  1. (1)
    The map $f$ is an $M$-local homotopy equivalence.
  2. (2)
    The induced map on $M$-localizations
    $f_{M^{-1}\mathbb{Z}}\colon X_{M^{-1}\mathbb{Z}}\rightarrow Y_{M^{-1}\mathbb{Z}}$
    is a weak homotopy equivalence.
  3. (3)
    The induced map on $M$-localizations
    $f_{M^{-1}\mathbb{Z}}\colon X_{M^{-1}\mathbb{Z}}\rightarrow Y_{M^{-1}\mathbb{Z}}$
    is a homotopy equivalence.
  4. (4)
    For every $M$-local space $Z$, the map
    $f_{M^{-1}\mathbb{Z}}^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(Y_{M^{-1}\mathbb{Z}},Z\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X_{M^{-1}\mathbb{Z}},Z\right)$
    is a weak homotopy equivalence.
  5. (5)
    For every $M$-local space $Z$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(Y,Z\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Z\right)$
    is a weak homotopy equivalence.
Proof.
Firstly, (1) and (2) are equivalent by applying Example 6.3 to the commutative diagram
[…]
whose vertical maps are $M$-local homotopy equivalences. Then, (2) and (3) are equivalent by the Whitehead theorem, as both $X_{M^{-1}\mathbb{Z}}$ and $Y_{M^{-1}\mathbb{Z}}$ are CW complexes by Remark 6.6. Moreover, (4) and (5) are equivalent because, for an $M$-local space $Z$, we have the commutative diagram
${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(Y_{M^{-1}\mathbb{Z}},Z\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X_{M^{-1}\mathbb{Z}},Z\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(Y,Z\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Z\right)}}$$\scriptstyle{f_{M^{-1}\mathbb{Z}}^{\ast}}$$\scriptstyle{f^{\ast}}$
whose vertical maps are weak homotopy equivalences by Lemma 6.7.
It remains to show that (3) and (4) are equivalent. If (3) holds, then Lemma 6.10(2) informs us that, for every ($M$-local) space $Z$, the map
$f_{M^{-1}\mathbb{Z}}^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(Y_{M^{-1}\mathbb{Z}},Z\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X_{M^{-1}\mathbb{Z}},Z\right)$
is a weak homotopy equivalence. Conversely, if (4) holds, then using the induced bijection of path components
$f_{M^{-1}\mathbb{Z}}^{\ast}\colon\left[Y_{M^{-1}\mathbb{Z}},Z\right]\xrightarrow{\cong}\left[X_{M^{-1}\mathbb{Z}},Z\right]$
for $Z=X_{M^{-1}\mathbb{Z}}$ and $Z=Y_{M^{-1}\mathbb{Z}}$ yields a homotopy inverse of $f_{M^{-1}\mathbb{Z}}$. ∎
It is the equivalence of (1) and (5) in Proposition 6.11 that is relevant to our work in Section 7. Again, we need the pointed version of that equivalence, which is proven identically.
Proposition 6.12.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A pointed map of $1$-connected CW complexes $f\colon X\rightarrow Y$ is an $M$-local homotopy equivalence if and only if, for every $M$-local pointed space $Z$, the map
$f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(Y,Z\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(X,Z\right)$
is a weak homotopy equivalence.
7. The family of model structures on $1$-reduced simplicial sets
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We left Bousfield localize the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ to get a model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ whose weak equivalences are the $M$-local homotopy equivalences. In particular, we obtain a model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ whose weak equivalences are the rational homotopy equivalences. Thus, we recover the family of model structures on $1$-reduced simplicial sets from [Qui69, Part II, Theorem 2.2]. These model structures inherit the generating cofibrations and the simplicial structure that we found for ${\mathcal{SS}\mathrm{ets}}_{1}$.
7.1. Left Bousfield localization
We give an overview of left Bousfield localization; a reference is [Hir03, Chapters 1-4]. Let $\mathcal{M}$ be a simplicial model category and $T$ be a set of morphisms in $\mathcal{M}$.
Definition 7.1.
  1. (1)
    An object $K$ of $\mathcal{M}$ is $T$-local if $K$ is fibrant in $\mathcal{M}$ and, for every morphism $\varphi\colon A\rightarrow B$ in $T$, the map
    $\varphi^{\ast}\colon\operatorname{Map}_{\mathcal{M}}\left(B,K\right)\rightarrow\operatorname{Map}_{\mathcal{M}}\left(A,K\right)$
    is a weak homotopy equivalence.
  2. (2)
    A morphism $f\colon X\rightarrow Y$ in $\mathcal{M}$ is a $T$-local equivalence if, for every $T$-local object $K$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{M}}\left(Y,K\right)\rightarrow\operatorname{Map}_{\mathcal{M}}\left(X,K\right)$
    is a weak homotopy equivalence.
  3. (3)
    A $T$-localization of an object $X$ of $\mathcal{M}$ is a $T$-local equivalence $f\colon X\rightarrow X_{T}$ with $T$-local target.
Example 7.2.
Every morphism in $T$ and every weak equivalence in $\mathcal{M}$ is a $T$-local equivalence.
Theorem 7.3 ([Bar10, Theorem 4.7]).
Let $\mathcal{M}$ be a combinatorial, left proper, and simplicial model category. Let $T$ be a set of morphisms in $\mathcal{M}$. There is a model structure $\mathcal{L}_{T}\mathcal{M}$ on the underlying category of $\mathcal{M}$, called the left Bousfield localization of $\mathcal{M}$ at $T$, in which a morphism $f\colon X\rightarrow Y$ is
  1. (1)
    a weak equivalence if and only if $f$ is a $T$-local equivalence; and
  2. (2)
    a cofibration if and only if $f$ is a cofibration in $\mathcal{M}$.
An object $K$ is fibrant in $\mathcal{L}_{T}\mathcal{M}$ if and only if $K$ is $T$-local. The model category $\mathcal{L}_{T}\mathcal{M}$ is combinatorial and left proper, and it inherits the simplicial structure of $\mathcal{M}$.
Remark 7.4.
Note that a $T$-localization in $\mathcal{M}$ is a fibrant replacement in $\mathcal{L}_{T}\mathcal{M}$.
The left Bousfield localization of $\mathcal{M}$ at $T$ enlarges, in a systematic fashion, the class of weak equivalences of $\mathcal{M}$ to encompass the morphisms in the set $T$. Observe that $\mathcal{L}_{T}\mathcal{M}$ inherits the generating cofibrations of $\mathcal{M}$, but it generally does not also inherit the generating acyclic cofibrations of $\mathcal{M}$ as, in general, the class of acyclic cofibrations in $\mathcal{M}$ is properly contained in the class of acyclic cofibrations in $\mathcal{L}_{T}\mathcal{M}$.
7.2. The set for the left Bousfield localization
For a multiplicative subset $M$ of $\mathbb{Z}$, our characterization of $M$-local spaces from Proposition 6.9 suggests that we produce a $1$-reduced simplicial analog of the set of maps
$\left\{S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}\mid n\geq 2\right\}$
in order to left Bousfield localize the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ at that set.
A sensible first suggestion is to start with the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ and use it to emulate the construction of $S^{n}_{M^{-1}\mathbb{Z}}$ in the $1$-reduced setting. This approach runs into a problem: the construction of $S^{n}_{M^{-1}\mathbb{Z}}$ involves self-maps of $S^{n}$ of arbitrary degree, whereas the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ only has the trivial self-maps.
Lemma 7.5.
For $n\geq 2$, the only self-maps of $S_{\mathrm{simp}}^{n}$ are the constant map and the identity map.
Proof.
Let $f\colon S_{\mathrm{simp}}^{n}\rightarrow S_{\mathrm{simp}}^{n}$ be a self-map of $S_{\mathrm{simp}}^{n}$. Let $x$ denote the unique 00-simplex of $S_{\mathrm{simp}}^{n}$ and let $\epsilon$ denote the non-degenerate $n$-simplex of $S_{\mathrm{simp}}^{n}$. There are two possibilities for the $n$-simplex $f_{n}(\epsilon)$ of $S_{\mathrm{simp}}^{n}$.
  1. (1)
    If $f_{n}(\epsilon)=\epsilon$, then $f$ is the identity map of $S_{\mathrm{simp}}^{n}$.
  2. (2)
    If $f_{n}(\epsilon)$ is the iterated degeneracy of $x$, then $f$ is the constant self-map of $S_{\mathrm{simp}}^{n}$.
Thus, we have to settle for an indirect route. We start with the CW inclusion
$S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}.$
We apply the singular complex functor to obtain the inclusion of simplicial sets
Sing($\operatorname{Sing}(S^{n})\rightarrow\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}).$
Then, we apply the $1$st Eilenberg subcomplex functor to obtain the inclusion of $1$-reduced simplicial sets
$E_{1}\operatorname{Sing}(S^{n})\rightarrow E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}).$
To confirm that this map is the analog we are looking for, we study its geometric realization
$|E_{1}\operatorname{Sing}(S^{n})|\rightarrow|E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|.$
In particular, we study the commutative diagram of CW complexes
${{|E_{1}\operatorname{Sing}(S^{n})|}}$${{|\operatorname{Sing}(S^{n})|}}$${{S^{n}}}$${{|E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|}}$|Sing(${{|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|}}$${{S^{n}_{M^{-1}\mathbb{Z}},}}$
where the horizontal maps in the left-hand square are CW subcomplex inclusions and the horizontal maps in the right-hand square are natural maps from the geometric realization-singular complex adjunction. We show that all the horizontal maps are weak homotopy equivalences.
Proposition 7.6.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $n\geq 2$. In the commutative diagram
${{|E_{1}\operatorname{Sing}(S^{n})|}}$${{|\operatorname{Sing}(S^{n})|}}$${{S^{n}}}$${{|E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|}}$|Sing(${{|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|}}$${{S^{n}_{M^{-1}\mathbb{Z}},}}$
all the horizontal maps are weak homotopy equivalences.
Proof.
By Corollary 4.5, the horizontal maps in the left-hand square are weak homotopy equivalences. By Theorem 4.1(2), the horizontal maps in the right-hand square are weak homotopy equivalences, as well. ∎
Proposition 7.6 translates to the success of our approach. Thus, for a multiplicative subset $M$ of $\mathbb{Z}$, we left Bousfield localize ${\mathcal{SS}\mathrm{ets}}_{1}$ at the set of maps
$T_{M^{-1}\mathbb{Z}}=\left\{E_{1}\operatorname{Sing}(S^{n})\rightarrow E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\right\}.$
7.3. The family of model structures on $1$-reduced simplicial sets
For a multiplicative subset $M$ of $\mathbb{Z}$, we describe the left Bousfield localization of the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ at the set of maps
$T_{M^{-1}\mathbb{Z}}=\left\{E_{1}\operatorname{Sing}(S^{n})\rightarrow E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\right\}.$
In particular, we describe the $T_{M^{-1}\mathbb{Z}}$-local objects and the $T_{M^{-1}\mathbb{Z}}$-local equivalences in ${\mathcal{SS}\mathrm{ets}}_{1}$. To state our characterizations, we translate the definitions from Section 6.1 from the language of $1$-connected topological spaces to that of $1$-reduced simplicial sets using geometric realization.
Definition 7.7.
Let $M$ be a multiplicative subset of $\mathbb{Z}$.
  1. (1)
    A map of $1$-reduced simplicial sets $f$ is an $M$-local homotopy equivalence if $|f|$ is an $M$-local homotopy equivalence of $1$-connected spaces.
  2. (2)
    A $1$-reduced simplicial set $K$ is $M$-local if $K$ is a Kan complex and $|K|$ is an $M$-local space.
  3. (3)
    An $M$-localization of a $1$-reduced simplicial set $X$ is an $M$-local homotopy equivalence $f\colon X\rightarrow X_{M^{-1}\mathbb{Z}}$ with $M$-local target.
When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we speak of a rational homotopy equivalence, a rational simplicial set, and a rationalization.
Example 7.8.
By Proposition 7.6, every map in $T_{M^{-1}\mathbb{Z}}$ is an $M$-localization. Note that Corollary 3.19 says that the target of every map in $T_{M^{-1}\mathbb{Z}}$ is a Kan complex.
The following proposition helps us transfer the rational homotopy theory of $1$-connected spaces from Section 6 to $1$-reduced simplicial sets.
Proposition 7.9 ([Hir03, Proposition 1.1.11]).
If $A$ and $X$ are pointed simplicial sets and $X$ is a Kan complex, then there is a weak homotopy equivalence
$\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(A,X\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|A|,|X|\right)$
that is natural in $A$ and in $X$.
By a series of characterizations, we show that a $1$-reduced simplicial set is $T_{M^{-1}\mathbb{Z}}$-local if and only if it is $M$-local.
Proposition 7.10.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $X$ be a $1$-reduced simplicial set. The following statements are equivalent.
  1. (1)
    The $1$-reduced simplicial set $X$ is $T_{M^{-1}\mathbb{Z}}$-local.
  2. (2)
    The $1$-reduced simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
    $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(E_{1}\operatorname{Sing}(S^{n}),X\right)$
    is a weak homotopy equivalence.
  3. (3)
    The $1$-reduced simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
    $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}),X\right)$
    is a weak homotopy equivalence.
  4. (4)
    The $1$-reduced simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
    $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n})|,|X|\right)$
    is a weak homotopy equivalence.
  5. (5)
    The $1$-reduced simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
    $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n}_{M^{-1}\mathbb{Z}},|X|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n},|X|)$
    is a weak homotopy equivalence.
  6. (6)
    The $1$-reduced simplicial set $X$ is a Kan complex such that $|X|$ is an $M$-local space.
  7. (7)
    The $1$-reduced simplicial set $X$ is $M$-local.
Hence, $X$ is $T_{M^{-1}\mathbb{Z}}$-local if and only if $X$ is $M$-local.
Proof.
Firstly, (1) and (2) are equivalent by the definition of a $T_{M^{-1}\mathbb{Z}}$-local object in ${\mathcal{SS}\mathrm{ets}}_{1}$ combined with the characterization of fibrant objects in ${\mathcal{SS}\mathrm{ets}}_{1}$ from Theorem 3.17. Then, (2) and (3) are equivalent by the commutative diagram
${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}),X\right)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(E_{1}\operatorname{Sing}(S^{n}),X\right),}}$
whose horizontal maps are weak homotopy equivalences by Proposition 7.6. Moreover, (3) and (4) are equivalent by the commutative diagram
${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}),X\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n})|,|X|\right),}}$
whose horizontal maps are weak homotopy equivalences by Proposition 7.9. Furthermore, (4) and (5) are equivalent by the commutative diagram
${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n}_{M^{-1}\mathbb{Z}},|X|)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n},|X|)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n})|,|X|\right),}}$
whose horizontal maps are weak homotopy equivalences by Proposition 7.6. Also, (5) and (6) are equivalent by Proposition 6.9. Lastly, (6) and (7) are equivalent by the definition of an $M$-local simplicial set. ∎
By a similar series of characterizations, we establish that a map of $1$-reduced simplicial sets is a $T_{M^{-1}\mathbb{Z}}$-local equivalence if and only if it is an $M$-local homotopy equivalence. To that end, we need two auxiliary lemmas.
Lemma 7.11.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. If $Z$ is an $M$-local pointed space, then $E_{1}\operatorname{Sing}(Z)$ is an $M$-local simplicial set.
Proof.
By Corollary 3.19, $E_{1}\operatorname{Sing}(Z)$ is a Kan complex. Then, Corollary 4.5 and Theorem 4.1(2) provide the weak homotopy equivalences
$|E_{1}\operatorname{Sing}(Z)|\rightarrow|\operatorname{Sing}(Z)|\rightarrow Z,$
so $|E_{1}\operatorname{Sing}(Z)|$ is $M$-local. Overall, $E_{1}\operatorname{Sing}(Z)$ is $M$-local, as required. ∎
Lemma 7.12.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets. The following statements are equivalent.
  1. (1)
    For every $M$-local simplicial set $K$, the map
    $|f|^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,|K|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,|K|)$
    is a weak homotopy equivalence.
  2. (2)
    For every $M$-local pointed space $Z$, the map
    $|f|^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,Z)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,Z)$
    is a weak homotopy equivalence.
Proof.
Suppose (1) holds. Let $Z$ be an $M$-local pointed space. Then, by Lemma 7.11, $E_{1}\operatorname{Sing}(Z)$ is an $M$-local simplicial set. There is a commutative diagram
${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|Y|,|E_{1}\operatorname{Sing}(Z)|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|Y|,|\operatorname{Sing}(Z)|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,Z)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|X|,|E_{1}\operatorname{Sing}(Z)|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|X|,|\operatorname{Sing}(Z)|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,Z)}}$
whose horizontal maps are weak homotopy equivalences by Corollary 4.5 and Theorem 4.1(2) and whose left-hand vertical map is a weak homotopy equivalence by (1). Thus, the right-hand vertical map is a weak homotopy equivalence, which proves (2).
Conversely, suppose (2) holds. Let $K$ be an $M$-local simplicial set. Then, $|K|$ is an $M$-local pointed space, so
$|f|^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|Y|,|K|\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|X|,|K|\right)$
is a weak homotopy equivalence, which proves (1). ∎
We are now ready for the series of characterizations that describes the $T_{M^{-1}\mathbb{Z}}$-local equivalences.
Proposition 7.13.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets. The following statements are equivalent.
  1. (1)
    The map $f$ is a $T_{M^{-1}\mathbb{Z}}$-local equivalence.
  2. (2)
    For every $T_{M^{-1}\mathbb{Z}}$-local simplicial set $K$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(X,K)$
    is a weak homotopy equivalence.
  3. (3)
    For every $M$-local simplicial set $K$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(X,K)$
    is a weak homotopy equivalence.
  4. (4)
    For every $M$-local simplicial set $K$, the map
    $|f|^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,|K|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,|K|)$
    is a weak homotopy equivalence.
  5. (5)
    For every $M$-local pointed space $Z$, the map
    $|f|^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,Z)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,Z)$
    is a weak homotopy equivalence.
  6. (6)
    The map $|f|$ is an $M$-local homotopy equivalence.
  7. (7)
    The map $f$ is an $M$-local homotopy equivalence.
Hence, $f$ is a $T_{M^{-1}\mathbb{Z}}$-local equivalence if and only $f$ is an $M$-local homotopy equivalence.
Proof.
Firstly, (1) and (2) are equivalent by the definition of a $T_{M^{-1}\mathbb{Z}}$-local equivalence. Then, (2) and (3) are equivalent by Proposition 7.10. Now, (3) and (4) are equivalent by the commutative diagram
${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(Y,K)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,|K|)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(X,K)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,|K|),}}$$\scriptstyle{f^{\ast}}$$\scriptstyle{|f|^{\ast}}$
whose horizontal maps are weak homotopy equivalences by Proposition 7.9. Moreover, (4) and (5) are equivalent by Lemma 7.12. Furthermore, (5) and (6) are equivalent by Proposition 6.12. Lastly, (6) and (7) are equivalent by the definition of an $M$-local homotopy equivalence. ∎
Propositions 7.10 and 7.13 together describe the left Bousfield localization of ${\mathcal{SS}\mathrm{ets}}_{1}$ at the set $T_{M^{-1}\mathbb{Z}}$.
Theorem 7.14.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ on the category of $1$-reduced simplicial sets in which a map $f$ is
  1. (1)
    a weak equivalence if and only if $f$ is an $M$-local homotopy equivalence; and
  2. (2)
    a cofibration if and only if $f$ is a monomorphism.
A $1$-reduced simplicial set $X$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $X$ is $M$-local. The model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ is combinatorial, left proper, and simplicial with the structure from Theorem 5.15.
Theorem 7.14 recovers the family of model structures on $1$-reduced simplicial sets in [Qui69, Part II, Theorem 2.2], as well as the characterization of fibrant objects in [Qui69, Part II, Corollary 2.6]. The most prominent model structure in this family is that corresponding to the case $M^{-1}\mathbb{Z}=\mathbb{Q}$, so we present it separately.
Theorem 7.15.
There is a model structure $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ on the category of $1$-reduced simplicial sets in which a map $f$ is
  1. (1)
    a weak equivalence if and only if $f$ is a rational homotopy equivalence; and
  2. (2)
    a cofibration if and only if $f$ is a monomorphism.
A $1$-reduced simplicial set $X$ is fibrant in $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $X$ is rational. The model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ is combinatorial, left proper, and simplicial with the structure from Theorem 5.15.
8. An approach using localization with respect to homology
We give an alternative description of the family of model structures from Theorem 7.14. This description uses Bousfield localization with respect to homology, as developed in [Bou75], together with Theorem 3.15 on left-induced model structures. The idea is that, instead of first left-transferring the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ and then localizing, we can first localize the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ and then left-transfer it. As a pointed simplicial set need not be $1$-connected, we do not localize ${\mathcal{SS}\mathrm{ets}}_{\ast}$ with respect to homotopy, but rather with respect to homology. Because the two notions are in agreement for $1$-connected spaces, the left transfer to ${\mathcal{SS}\mathrm{ets}}_{1}$ yields the model structures from Theorem 7.14, and we extract our alternative description.
8.1. Bousfield localization with respect to homology
Firstly, we introduce the following terminology.
Definition 8.1.
Let $M$ be a multiplicative subset of $\mathbb{Z}$.
  1. (1)
    A map of topological spaces $f\colon X\rightarrow Y$ is an $M$-local homology isomorphism if, for every $n\geq 0$, the induced map
    $f_{\ast}\colon H_{n}(X;M^{-1}\mathbb{Z})\rightarrow H_{n}(Y;M^{-1}\mathbb{Z})$
    is an isomorphism.
  2. (2)
    A simplicial map $g$ is an $M$-local homology isomorphism if $|g|$ is an $M$-local homology isomorphism of topological spaces.
When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we speak of a rational homology isomorphism.
Remark 8.2.
As the abelian group $M^{-1}\mathbb{Z}$ is flat [Rot09, Theorem 4.80], the universal coefficient theorem for homology yields the natural isomorphism
$H_{n}(X;M^{-1}\mathbb{Z})\cong H_{n}(X;\mathbb{Z})\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}.$
Thus, $M$-local homology discards torsion from integral homology in the same fashion $M$-local homotopy discards torsion from higher homotopy groups. When $M^{-1}\mathbb{Z}=\mathbb{Q}$, all torsion is discarded in both cases.
Example 8.3.
By Remark 8.2, the unique map from $\mathbb{R}P^{2}$ to the point is a rational homology isomorphism.
The natural isomorphism from Remark 8.2 also has the following consequence that we note for later use.
Lemma 8.4.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Every integral homology isomorphism is an $M$-local homology isomorphism.
Also for later use, we need to recall the following theorem that relates integral homology and homotopy.
Theorem 8.5 ([DK01, Corollary 6.69(1)]).
Every weak homotopy equivalence is an integral homology isomorphism.
We also need to recall another theorem of a similar nature for future reference. This theorem informs us that, under the hypothesis of $1$-connectedness, $M$-local homology agrees with $M$-local homotopy.
Theorem 8.6 ([FHT01, Theorem 8.6]).
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A map of $1$-connected spaces $f$ is an $M$-local homology isomorphism if and only if $f$ is an $M$-local homotopy equivalence.
For a multiplicative subset $M$ of $\mathbb{Z}$, Bousfield localization with respect to homology endows the category $\mathcal{SS}\mathrm{ets}$ of simplicial sets with a model structure whose weak equivalences are the $M$-local homology isomorphisms.
Theorem 8.7 ([Bou75, Theorem 10.2]).
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}$ on the category of simplicial sets in which a map $f$ is
  1. (1)
    a weak equivalence if and only if $f$ is an $M$-local homology isomorphism; and
  2. (2)
    a cofibration if and only if $f$ is a monomorphism.
This model category is combinatorial, left proper, and simplicial with the structure from Example 5.2.
Remark 8.8.
By [Bou97, §2.5], $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}$ is the left Bousfield localization of $\mathcal{SS}\mathrm{ets}$ at a single map, which allows us to infer that $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}$ inherits the properties of being combinatorial, left proper, and simplicial.
When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we get a model structure whose weak equivalences are the rational homology isomorphisms.
Theorem 8.9 ([Bou75, Theorem 10.2]).
There is a model structure $\mathcal{L}_{H\mathbb{Q}}\mathcal{SS}\mathrm{ets}$ on the category of simplicial sets in which a map $f$ is
  1. (1)
    a weak equivalence if and only if $f$ is a rational homology isomorphism; and
  2. (2)
    a cofibration if and only if $f$ is a monomorphism.
This model category is combinatorial, left proper, and simplicial with the structure from Example 5.2.
We use the pointed version of Theorem 8.7, which holds by the same proof.
Theorem 8.10.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ on the category of pointed simplicial sets in which a map $f$ is
  1. (1)
    a weak equivalence if and only if $f$ is an $M$-local homology isomorphism; and
  2. (2)
    a cofibration if and only if $f$ is a monomorphism.
This model category is combinatorial, left proper, and simplicial with the structure from Example 5.3.
8.2. An alternative description of the family of model structures
For a multiplicative subset $M$ of $\mathbb{Z}$, we show that the model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ from Theorem 7.14 is the left transfer of the model category $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ from Theorem 8.10 along the adjunction
$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}$
from Lemma 2.23(2). For this left transfer, we rely on Theorem 3.15, so we verify the condition in its hypothesis.
Lemma 8.11.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets. If $f$ has the right lifting property with respect to every monomorphism of $1$-reduced simplicial sets, then $f$ is an $M$-local homology isomorphism.
Proof.
Suppose $f$ has the right lifting property with respect to every monomorphism of $1$-reduced simplicial sets. By Lemma 3.16, we know that $f$ is a weak homotopy equivalence. Then, Theorem 8.5 informs us that $f$ is an integral homology isomorphism. Therefore, by Lemma 8.4, $f$ is an $M$-local homology isomorphism. ∎
We invoke Theorem 3.15 to give an alternative approach to the family of model structures from Theorem 7.14.
Proposition 8.12.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. The model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ is the left transfer of the model category $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ along the adjunction
$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}.$
Proof.
Theorem 3.15 says that we can left transfer the model structure $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ along the adjunction
$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}$
to a model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$ in which a map $f$ is
  1. (1)
    a weak equivalence if and only if $f$ is an $M$-local homology isomorphism; and
  2. (2)
    a cofibration if and only if $f$ is a monomorphism.
By Theorem 8.6, a map of $1$-reduced simplicial sets $f$ is an $M$-local homology isomorphism if and only if $f$ is an $M$-local homotopy equivalence. Thus, the left transfer gives us the model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$. ∎
9. Rational homotopy theory beyond $1$-connected spaces
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We study some elements of rational homotopy theory beyond $1$-connected spaces in order to extend Quillen’s family of model structures from $1$-reduced simplicial sets to all simplicial sets. In particular, we characterize $M$-local spaces in terms of function complexes, which prepares us for the left Bousfield localization in Section 10. We expect this characterization to be familiar to the rational homotopy theory community, but we could not find it in the literature, so we provide a self-contained proof.
9.1. Localization of spaces
We adopt the generalization of rational homotopy theory to all spaces from [GTHT00, §6]; in particular, we generalize Definition 6.1 to all spaces as follows.
Definition 9.1.
Let $M$ be a multiplicative subset of $\mathbb{Z}$.
  1. (1)
    A continuous map $f\colon X\rightarrow Y$ is an $M$-local homotopy equivalence if, for every $n\leq 1$, the induced map
    $f_{\ast}\colon\pi_{n}(X)\rightarrow\pi_{n}(Y)$
    is an isomorphism and, for every $n\geq 2$, the induced map
    $f_{\ast}\underset{\mathbb{Z}}{\otimes}\operatorname{id}_{M^{-1}\mathbb{Z}}\colon\pi_{n}(X)\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}\rightarrow\pi_{n}(Y)\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}$
    is an isomorphism.
  2. (2)
    A space $Y$ is $M$-local if, for every $n\geq 2$, the abelian group $\pi_{n}(Y)$ is an $M^{-1}\mathbb{Z}$-module.
  3. (3)
    An $M$-localization of a space $Y$ is an $M$-local homotopy equivalence $f\colon Y\rightarrow Y_{M^{-1}\mathbb{Z}}$ with $M$-local target.
When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we speak of a rational homotopy equivalence, a rational space, and a rationalization.
Our goal of extending Quillen’s family of model structures to all simplicial sets precludes us from imposing any connectedness or nilpotency hypothesis on spaces. The absence of any connectedness hypothesis forces us to keep track of the basepoint across Definition 9.1 and approach disconnected spaces component-by-component. The absence of any nilpotency hypothesis deprives us of any sensible notion of tensoring the fundamental group with $\mathbb{Q}$, so we have to ask that a rational homotopy equivalence induce an isomorphism of fundamental groups.
Example 9.2.
Every weak homotopy equivalence is an $M$-local homotopy equivalence. Also, every $M$-local homotopy equivalence of $M$-local spaces is a weak homotopy equivalence.
Example 9.3.
The spaces $S^{1}$, $S^{1}\vee S^{1}$, and $\mathbb{R}P^{\infty}$ are $M$-local because their higher homotopy groups vanish.
Example 9.3 begs the question of why our definition of an $M$-local space imposes no condition on the fundamental group. In particular, we can ask that, for every element $m$ of $M$, the $m$-th power map $g\mapsto g^{m}$ be a self-bijection of the fundamental group, even though it is generally not a homomorphism for non-abelian groups. The problem with imposing this additional condition is that, with this stricter definition, there are spaces that do not admit any $M$-localization. This striking result is proven by Casacuberta in [Cas93, Proposition 2.1].
On the contrary, in our generalization of rational homotopy theory, every space can be localized, albeit not with the universal property that the localization of $1$-connected spaces enjoys.
Theorem 9.4 ([GTHT00, §6]).
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Every space admits an $M$-localization.
The proof of the theorem is concise and it sheds light on what goes wrong with the universal property of localization, so we present it to provide some commentary afterwards.
Proof.
If $X$ is a connected space, then its $1$st Postnikov approximation yields the fibration
$\widetilde{X}\rightarrow X\rightarrow K(\pi_{1}(X),1),$
where $\widetilde{X}$ is the universal cover of $X$ and $K(\pi_{1}(X),1)$ is an Eilenberg-Mac Lane space. Since $\widetilde{X}$ is a $1$-connected space, we invoke fiberwise $M$-localization, in the sense of [Hir03, Theorem 6.1.3], to obtain a map of fibrations
[…]
where the map on homotopy fibers $\widetilde{X}\rightarrow\widetilde{X}_{M^{-1}\mathbb{Z}}$ is an $M$-localization of the $1$-connected space $\widetilde{X}$ and, by the induced map of long exact sequences of homotopy groups, the map on total spaces $X\rightarrow X_{M^{-1}\mathbb{Z}}$ is the required $M$-localization of $X$. Applying this treatment to each component yields the claim for disconnected spaces. ∎
The downside with fiberwise localization in our setting is that, because it applies to fibrations over a fixed base space $B$, its universal property states that it is initial among fibrations over $B$ with local fibers. Thus, our application of fiberwise localization in the proof of Theorem 9.4 only yields a universal property of $M$-localization over the Eilenberg-Mac Lane space $K(\pi_{1}(X),1)$, and we are neither aware of how to get rid of this restriction nor convinced that we should be able to do so in the first place.
The absence of an analog of the universal property of localization for $1$-connected spaces has the following domino effect. The universal property of localization was instrumental for our characterization of local homotopy equivalences in terms of function complexes, which in turn gave us the description of the weak equivalences in the left Bousfield localization in Section 7. Thus, the left Bousfield localization in Section 10 is only a partial extension of that in Section 7 to all simplicial sets: we produce model structures whose fibrant objects are the local simplicial sets, but we do not know whether the weak equivalences are the local homotopy equivalences.
9.2. The construction of local spheres
Let $M$ be a multiplicative subset of $\mathbb{Z}$ and let $n\geq 2$. To salvage our function complex characterization of $M$-local spaces in the absence of a universal property for $M$-localizatrion, we need the details of the CW complex construction of the $M$-local $n$-sphere $S^{n}_{M^{-1}\mathbb{Z}}$ from [FHT01, Chapter 9(a)].
Let $\mathbb{N}$ denote the set of positive integers. We enumerate
$M\cap\mathbb{N}=\left\{k_{1},k_{2},k_{3},\dots\right\}.$
We denote the $(n+1)$-disk by $D^{n+1}$. The $M$-local $n$-sphere $S^{n}_{M^{-1}\mathbb{Z}}$ is constructed as the pushout
[…]
where, for $j\geq 1$, the $j$-th $(n+1)$-cell $(D^{n+1})_{j}$ is attached to $\bigvee\limits_{i=0}^{\infty}(S^{n})_{i}$ by an attaching map $h$ given by
[…]
The map
$(-k_{j})\colon S^{n}\rightarrow(S^{n})_{j}$
is a self-map of the $n$-sphere $S^{n}$ that has degree $(-k_{j})$. In other words, the $1$-connected CW complex
[…]
is the mapping telescope of the sequence of maps
$S^{n}\xrightarrow{k_{1}}S^{n}\xrightarrow{k_{2}}S^{n}\xrightarrow{k_{3}}\cdots.$
9.3. A characterization of local spaces
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We use the construction of $S^{n}_{M^{-1}\mathbb{Z}}$ from Section 9.2, including our notation, to prove the following key lemma.
Lemma 9.5.
Let $M$ be a multiplicative subset of $\mathbb{Z}$ and let $n\geq 2$. Let $Y$ be a space such that $\pi_{n}(Y)$ is an $M^{-1}\mathbb{Z}$-module. Then, every map $f\colon S^{n}\rightarrow Y$ extends uniquely up to homotopy to a map $g\colon S^{n}_{M^{-1}\mathbb{Z}}\rightarrow Y$ such that the diagram
[…]
commutes.
Proof.
Given a map $f\colon S^{n}\rightarrow Y$, we know that $[f]$ is an element of the $M^{-1}\mathbb{Z}$-module $\pi_{n}(Y)$. Thus, there exists a unique element $[u]$ of $\pi_{n}(Y)$ such that
$[f]=k_{1}\cdot[u],$
and the representative $u\colon S^{n}\rightarrow Y$ of $[u]$ is unique up to homotopy. Using the notation in our construction of
[…]
we extend $f$ from $S^{n}=(S^{n})_{0}$ to $(S^{n})_{0}\vee(S^{n})_{1}$ uniquely up to homotopy by setting
$g|_{(S^{n})_{1}}=u.$
As
$[f]-\left(k_{1}\cdot[u]\right)=0\in\pi_{n}(Y),$
we can further extend $f$ to
[…]
uniquely up to homotopy. Continuing further down the telescope in this fashion yields the claim. ∎
In particular, we need the following consequence of Lemma 9.5 to give a characterization of $M$-local spaces.
Corollary 9.6.
Let $M$ be a multiplicative subset of $\mathbb{Z}$ and let $n\geq 2$. Let $Y$ be an $M$-local space. Then, every map $f\colon S^{n}\rightarrow Y$ extends uniquely up to homotopy to a map $g\colon S^{n}_{M^{-1}\mathbb{Z}}\rightarrow Y$ such that the diagram
[…]
commutes.
We salvage our characterization of $M$-local spaces for the left Bousfield localization in the next section.
Proposition 9.7.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A space $Y$ is $M$-local if and only if, for every $n\geq 2$, the map
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n},Y\right)$
induced by the inclusion $S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$ is a weak homotopy equivalence.
Proof.
If $Y$ is $M$-local, then Corollary 9.6 implies that, for every $n\geq 2$, the map
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n},Y\right)$
is a weak homotopy equivalence. The converse implication is established as in the proof of Proposition 6.8. ∎
10. The family of model structures on simplicial sets
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We left Bousfield localize the model category ${\mathcal{SS}\mathrm{ets}}$ of simplicial sets to obtain a model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ whose fibrant objects are the $M$-local ones. In particular, we produce a model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ whose fibrant objects are the rational simplicial sets. Thus, we partially extend Quillen’s family of model structures from $1$-reduced simplicial sets to all simplicial sets. Our family of model structures is meant to axiomatize the generalization of rational homotopy theory that we studied in Section 9.
10.1. The set for the left Bousfield localization
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We left Bousfield localize ${\mathcal{SS}\mathrm{ets}}$ at the set of maps
$T_{M^{-1}\mathbb{Z}}=\left\{E_{1}\operatorname{Sing}(S^{n})\rightarrow E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\right\}$
from Section 7.2. We observe that, by Proposition 7.6, the application of the $1$st Eilenberg subcomplex functor $E_{1}$ is superfluous for this localization, for we are no longer working within the confines of $1$-reduced simplicial sets. In other words, left Bousfield localizing ${\mathcal{SS}\mathrm{ets}}$ at the set of maps
$B_{M^{-1}\mathbb{Z}}=\left\{\operatorname{Sing}(S^{n})\rightarrow\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\right\}$
yields the same model category, so we equivalently localize ${\mathcal{SS}\mathrm{ets}}$ at the set of maps $B_{M^{-1}\mathbb{Z}}$ instead.
10.2. The family of model structures on simplicial sets
For a multiplicative subset $M$ of $\mathbb{Z}$, we describe the left Bousfield localization of the model category ${\mathcal{SS}\mathrm{ets}}$ at the set of maps
$B_{M^{-1}\mathbb{Z}}=\left\{\operatorname{Sing}(S^{n})\rightarrow\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\right\}.$
In particular, we characterize the $B_{M^{-1}\mathbb{Z}}$-local objects in ${\mathcal{SS}\mathrm{ets}}$ in terms of the following definition.
Definition 10.1.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A simplicial set $K$ is $M$-local if $K$ is a Kan complex and $|K|$ is an $M$-local space. When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we speak of a rational simplicial set.
Example 10.2.
By Proposition 7.6, the target of every map in $B_{M^{-1}\mathbb{Z}}$ is an $M$-local simplicial set. Note that Example 3.2 says that the target of every map in $B_{M^{-1}\mathbb{Z}}$ is a Kan complex.
The following proposition helps us transfer the rational homotopy theory from Section 9 to simplicial sets.
Proposition 10.3 ([Hir03, Proposition 1.1.11]).
If $A$ and $X$ are simplicial sets and $X$ is a Kan complex, then there is a weak homotopy equivalence
$\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}\left(A,X\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|A|,|X|\right)$
that is natural in $A$ and in $X$.
By a series of characterizations, we show that a simplicial set is $B_{M^{-1}\mathbb{Z}}$-local if and only if it is $M$-local.
Proposition 10.4.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $X$ be a simplicial set. The following statements are equivalent.
  1. (1)
    The simplicial set $X$ is $B_{M^{-1}\mathbb{Z}}$-local.
  2. (2)
    The simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
    $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}\left(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}\left(\operatorname{Sing}(S^{n}),X\right)$
    is a weak homotopy equivalence.
  3. (3)
    The simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
    $\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n})|,|X|\right)$
    is a weak homotopy equivalence.
  4. (4)
    The simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
    $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},|X|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},|X|)$
    is a weak homotopy equivalence.
  5. (5)
    The simplicial set $X$ is a Kan complex such that $|X|$ is an $M$-local space.
  6. (6)
    The simplicial set $X$ is $M$-local.
Hence, $X$ is $B_{M^{-1}\mathbb{Z}}$-local if and only if $X$ is $M$-local.
Proof.
Firstly, (1) and (2) are equivalent by the definition of a $B_{M^{-1}\mathbb{Z}}$-local object in ${\mathcal{SS}\mathrm{ets}}$. Then, (2) and (3) are equivalent by the commutative diagram
${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}\left(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}\left(\operatorname{Sing}(S^{n}),X\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n})|,|X|\right),}}$
whose horizontal maps are weak homotopy equivalences by Proposition 10.3. Now, (3) and (4) are equivalent by the commutative diagram
${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},|X|)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},|X|)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n})|,|X|\right),}}$
whose horizontal maps are weak homotopy equivalences by Proposition 7.6. Also, (4) and (5) are equivalent by Proposition 9.7. Lastly, (5) and (6) are equivalent by the definition of an $M$-local simplicial set. ∎
Proposition 10.4 helps us describe the left Bousfield localization of ${\mathcal{SS}\mathrm{ets}}$ at the set of maps $B_{M^{-1}\mathbb{Z}}$.
Theorem 10.5.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ on the category ${\mathcal{SS}\mathrm{ets}}$ of simplicial sets in which a map $f\colon X\rightarrow Y$ is
  1. (1)
    a weak equivalence if and only if, for every $M$-local simplicial set $K$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(X,K)$
    is a weak homotopy equivalence; and
  2. (2)
    a cofibration if and only if $f$ is a monomorphism.
A simplicial set $X$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ if and only if $X$ is $M$-local. The model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ is combinatorial, left proper, and simplicial with the structure from Example 5.2.
Theorem 10.5 partially extends Quillen’s family of model structures from $1$-reduced simplicial sets to all simplicial sets. A distinguished model structure in this family is that corresponding to the case $M^{-1}\mathbb{Z}=\mathbb{Q}$.
Theorem 10.6.
There is a model structure $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ on the category ${\mathcal{SS}\mathrm{ets}}$ of simplicial sets in which a map $f\colon X\rightarrow Y$ is
  1. (1)
    a weak equivalence if and only if, for every rational simplicial set $K$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(X,K)$
    is a weak homotopy equivalence; and
  2. (2)
    a cofibration if and only if $f$ is a monomorphism.
A simplicial set $X$ is fibrant in $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ if and only if $X$ is rational. The model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ is combinatorial, left proper, and simplicial with the structure from Example 5.2.
11. The family of model structures on topological spaces
Quillen’s family of model structures on the category of $1$-reduced simplicial sets finds no counterpart for the category of $1$-connected spaces, for the latter category cannot be endowed with any model structure because it does not have all finite limits and colimits; this failure necessitates the use of $1$-reduced simplicial sets in Quillen’s model-categorical approach of rational homotopy theory.
To see that the category of $1$-connected spaces does not have all finite limits, consider the pullback square
[…]
where $f\colon\mathbb{R}^{2}\rightarrow\mathbb{R}$ is given by the formula $f(x,y)=x^{2}+y^{2}$, so $f^{-1}(1)=S^{1}$ is the unit circle in the plane. In this pullback square, we start with the contractible spaces $|\Delta[0]|$, $\mathbb{R}^{2}$, and $\mathbb{R}$, and we obtain $S^{1}$, which is not a $1$-connected space because $\pi_{1}(S^{1})\cong\mathbb{Z}$. Thus, we see that connectedness conditions on spaces do not interact well with limits and colimits, in contrast to skeletal conditions on simplicial sets.
However, now that we have a family of model structures on the category of simplicial sets, we continue our work by presenting its analog for the category $\mathcal{T}\mathrm{op}$ of topological spaces. This analog is produced by left Bousfield localizing the model structure on $\mathcal{T}\mathrm{op}$, which we recall in this section. Moreover, because we first characterized local topological spaces and then transferred this characterization to simplicial sets in this section, there is no additional work that we need to do in order to obtain a family of model structures on topological spaces.
11.1. The model category of topological spaces
We recall the model structure on $\mathcal{T}\mathrm{op}$ from [DS95, §8]. To describe it, we first need to introduce the following notions.
Definition 11.1.
  1. (1)
    Given a sequence of subspace inclusions
    $X_{0}\rightarrow X_{1}\rightarrow\cdots$
    such that, for every $n\geq 0$, the pair $(X_{n+1},X_{n})$ is a relative CW complex, we say that the map
    $X_{0}\rightarrow\operatorname{colim}(X_{n})$
    is a generalized relative CW inclusion.
  2. (2)
    A Serre fibration is a continuous map $p\colon X\rightarrow Y$ that has the right lifting property with respect to the class of inclusions
    $\left\{A\times\{0\}\rightarrow A\times[0,1]\mid A\textrm{ is a CW complex}\right\}.$
    In other words, a Serre fibration is a continuous map that has the homotopy lifting property with respect to every CW complex.
Remark 11.2.
A continuous map is often said to be a fibration if it has the homotopy lifting property with respect to every space. This notion is stronger than that of a Serre fibration, and it is sometimes called a Hurewicz fibration to emphasize this distinction. Every Hurewicz fibration is a Serre fibration; the converse implication finds a pathological counterexample in [Bro66, §1], but it holds for maps of CW complexes [Cau92].
Example 11.3.
Every space $X$ is Serre fibrant: the unique map from $X$ to the point is a Serre fibration.
We are now ready to describe the model category of topological spaces.
Theorem 11.4 ([DS95, Propositions 8.3 and 8.9]).
The category ${\mathcal{T}\mathrm{op}}$ of topological spaces has a model structure in which a map $f$ is
  1. (1)
    a weak equivalence if and only if $f$ is a weak homotopy equivalence;
  2. (2)
    a cofibration if and only if $f$ is a retract of a generalized relative CW inclusion; and
  3. (3)
    a fibration if and only if $f$ is a Serre fibration.
The model category of topological spaces is simplicial with the following structure.
Proposition 11.5 ([Hir03, Example 9.1.15]).
Given two spaces $X$ and $Y$ and a simplicial set $Z$, we define
  1. (1)
    the function complex $\operatorname{Map}_{{\mathcal{T}\mathrm{op}}}\left(X,Y\right)$ to be the simplicial set with $n$-simplices $\operatorname{Hom}_{{\mathcal{T}\mathrm{op}}}\left(X\times|\Delta[n]|,Y\right)$;
  2. (2)
    the tensor of $X$ with $Z$ to be the space $X\times|Z|$; and
  3. (3)
    the cotensor of $X$ with $Z$ to be the space $X^{|Z|}$ of maps from $|Z|$ to $X$ with the compact-open topology.
With these definitions, ${\mathcal{T}\mathrm{op}}$ has the structure of a simplicial model category.
The model category of topological spaces is also cofibrantly generated; its generating cofibrations arise from geometrically realizing those for the model category of simplicial sets.
Proposition 11.6 ([Hir03, Example 11.1.8]).
The model category ${\mathcal{T}\mathrm{op}}$ is cofibrantly generated.
  1. (1)
    A generating set of cofibrations for ${\mathcal{T}\mathrm{op}}$ is
    $|I|=\left\{|i_{n}|\colon|\partial\Delta[n]|\rightarrow|\Delta[n]|\mid n\geq 0\right\}.$
  2. (2)
    A generating set of acyclic cofibrations for ${\mathcal{T}\mathrm{op}}$ is
    $|J|=\left\{|j_{n,k}|\colon|\Lambda[n,k]|\rightarrow|\Delta[n]|\mid n\geq 1\textrm{ and }0\leq k\leq n\right\}.$
The model category of topological spaces differs from that of simplicial sets in two noteworthy respects.
  1. (1)
    It is left proper [DS95, Theorem 13.1.10], but not by Proposition 3.12, for not every space is cofibrant.
  2. (2)
    It is not combinatorial because the category ${\mathcal{T}\mathrm{op}}$ is not locally presentable [AR94, Example 1.18(5)].
The latter difference raises the question of whether ${\mathcal{T}\mathrm{op}}$ can be left Bousfield localized. To that end, we remark that ${\mathcal{T}\mathrm{op}}$ shares with ${\mathcal{SS}\mathrm{ets}}$ the feature of being a cellular model category [Hir03, Proposition 12.1.4]. This property is a different way of strengthening the property of being a cofibrantly generated model category, and it allows us to left Bousfield localize ${\mathcal{T}\mathrm{op}}$, for Theorem 7.3 holds true if we replace “combinatorial” with “cellular” throughout [Hir03, Theorem 4.1.1].
11.2. The family of model structures on topological spaces
Our work in Section 9 yields the promised family of model structures on topological spaces.
Theorem 11.7.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{T}\mathrm{op}}$ on the category ${\mathcal{T}\mathrm{op}}$ of topological spaces in which a map $f\colon X\rightarrow Y$ is
  1. (1)
    a weak equivalence if and only if, for every $M$-local space $K$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,K)$
    is a weak homotopy equivalence; and
  2. (2)
    a cofibration if and only if $f$ is a retract of a generalized relative CW inclusion.
A space $X$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{T}\mathrm{op}}$ if and only if $X$ is $M$-local. The model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{T}\mathrm{op}}$ is cellular, left proper, and simplicial with the structure from Proposition 11.5.
Proof.
Proposition 9.7 informs us that $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{T}\mathrm{op}}$ is the left Bousfield localization of ${\mathcal{T}\mathrm{op}}$ at the set of maps
$\left\{S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}\mid n\geq 2\right\}.$
Remark 11.8.
By Proposition 7.6, $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{T}\mathrm{op}}$ is also the left Bousfield localization of ${\mathcal{T}\mathrm{op}}$ at the set of maps
$|B_{M^{-1}\mathbb{Z}}|=\left\{|\operatorname{Sing}(S^{n})|\rightarrow|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|\mid n\geq 2\right\}.$
Thus, by [Hir03, Theorem 3.3.20], the Quillen equivalence
$|-|\colon\mathcal{SS}\mathrm{ets}\rightleftarrows\mathcal{T}\mathrm{op}\colon\operatorname{Sing}$
induces a Quillen equivalence
$|-|\colon\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}\rightleftarrows\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{T}\mathrm{op}\colon\operatorname{Sing}.$
This Quillen equivalence captures the fact that the model categories $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}$ and $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{T}\mathrm{op}$ both describe the $M$-local homotopy theory beyond $1$-connected spaces from Section 9.
Lastly, we distinguish the model category corresponding to the case $M^{-1}\mathbb{Z}=\mathbb{Q}$.
Theorem 11.9.
There is a model structure $\mathcal{L}_{\mathbb{Q}}{\mathcal{T}\mathrm{op}}$ on the category ${\mathcal{T}\mathrm{op}}$ of topological spaces in which a map $f\colon X\rightarrow Y$ is
  1. (1)
    a weak equivalence if and only if, for every rational space $K$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,K)$
    is a weak homotopy equivalence; and
  2. (2)
    a cofibration if and only if $f$ is a retract of a generalized relative CW inclusion.
A space $X$ is fibrant in $\mathcal{L}_{\mathbb{Q}}{\mathcal{T}\mathrm{op}}$ if and only if $X$ is rational. The model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{T}\mathrm{op}}$ is cellular, left proper, and simplicial with the structure from Proposition 11.5.
12. Rational homology versus rational homotopy
We conclude this paper by contrasting Bousfield’s localization of spaces with respect to rational homology to our localization of spaces with respect to rational homotopy. Even though rational homology and rational homotopy agree for $1$-connected spaces, this agreement does not extend to all spaces. Therefore, we verify that the two localized model structures on spaces are not equal.
12.1. Localization of topological spaces with respect to homology
Our argument is best presented using topological spaces, so we formulate the analog of Theorem 8.7 for the category $\mathcal{T}\mathrm{op}$ of topological spaces.
Theorem 12.1 ([Bou75, Theorem 10.2]).
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{T}\mathrm{op}$ on the category of topological spaces in which a map $f$ is
  1. (1)
    a weak equivalence if and only if $f$ is an $M$-local homology isomorphism; and
  2. (2)
    a cofibration if and only if $f$ is a retract of a generalized relative CW inclusion.
This model category is cellular, left proper, and simplicial with the structure from Proposition 11.5.
Remark 12.2.
By [Bou97, §2.5], $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{T}\mathrm{op}$ is the left Bousfield localization of $\mathcal{T}\mathrm{op}$ at a single map, so it inherits the properties of being cellular, left proper, and simplicial.
When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we get a model structure whose weak equivalences are the rational homology isomorphisms.
Theorem 12.3 ([Bou75, Theorem 10.2]).
There is a model structure $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ on the category of topological spaces in which a map $f$ is
  1. (1)
    a weak equivalence if and only if $f$ is a rational homology isomorphism; and
  2. (2)
    a cofibration if and only if $f$ is a retract of a generalized relative CW inclusion.
This model category is cellular, left proper, and simplicial with the structure from Proposition 11.5.
12.2. Rational homology versus rational homotopy
We show that Bousfield’s localization $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ with respect to rational homology does not coincide with our localization $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$ with respect to rational homotopy from Theorem 11.9. This result stems from the fact that rational homology and rational homotopy cease to agree beyond the realm of $1$-connected spaces, and our proof comprises a particular instance of this phenomenon.
Proposition 12.4.
The model structures $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ and $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$ on topological spaces are not equal.
Proof.
We show that the unique map
$f\colon\mathbb{R}P^{2}\rightarrow|\Delta[0]|$
is a weak equivalence in $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ but not a weak equivalence in $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$.
Firstly, $f$ is a weak equivalence in $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ because, by Example 8.3, $f$ is a rational homology isomorphism.
To show that $f$ is not a weak equivalence in $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$, we recall from Example 9.3 that $\mathbb{R}P^{\infty}$ is a rational space because its higher homotopy groups vanish. This recollection renders it sufficient to show that
$f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(|\Delta[0]|,\mathbb{R}P^{\infty})\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(\mathbb{R}P^{2},\mathbb{R}P^{\infty})$
is not a weak homotopy equivalence. In particular, it suffices to show that the induced map on path components
$f^{\ast}\colon[|\Delta[0]|,\mathbb{R}P^{\infty}]\rightarrow[\mathbb{R}P^{2},\mathbb{R}P^{\infty}]$
is not a bijection. To that end, we find a map from $\mathbb{R}P^{2}$ to $\mathbb{R}P^{\infty}$ that is not null. Such a map is the inclusion
$i\colon\mathbb{R}P^{2}\rightarrow\mathbb{R}P^{\infty},$
for its induced map on cohomology rings with $\mathbb{Z}/2\mathbb{Z}$-coefficients is the quotient map
$i^{\ast}\colon(\mathbb{Z}/2\mathbb{Z})[x]\rightarrow(\mathbb{Z}/2\mathbb{Z})[x]/(x^{3}),$
which is not trivial. Overall, $f$ is a weak equivalence in $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ but not a weak equivalence in $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$. ∎
Thanks to Example 3.2 and Theorem 4.1(2), we can safely apply the singular complex functor throughout our proof of Proposition 12.4 to obtain the analogous result for simplicial sets.
Proposition 12.5.
The model structures $\mathcal{L}_{H\mathbb{Q}}\mathcal{SS}\mathrm{ets}$ and $\mathcal{L}_{\mathbb{Q}}\mathcal{SS}\mathrm{ets}$ on simplicial sets are not equal.

### v2

@@
 Abstract.
-In Quillen’s paper on rational homotopy theory, the category of $1$-reduced simplicial sets is endowed with a family of model structures, the most prominent of which is the one whose weak equivalences are the rational homotopy equivalences.
+In Quillen’s paper on rational homotopy theory, the category of $1$-reduced simplicial sets is endowed with a family of model structures, the most prominent of which is the one in which the weak equivalences are the rational homotopy equivalences and the fibrant objects are the rational Kan complexes.
 In this paper, we give a modern approach to this family of model structures.
+We recover Quillen’s family of model structures by first left-transferring the model structure on pointed simplicial sets and then left Bousfield localizing at the rationalization maps of spheres.
+Applying this localization to the model category of all spaces yields a model category in which the weak equivalences are the rational homotopy equivalences in the extended sense of Gómez-Tato, Halperin, and Tanré and the fibrant objects are the rational spaces.
+Thus, we generalize Quillen’s family of model structures beyond the rational homotopy theory of $1$-connected spaces.
+Key words and phrases:
+rational homotopy theory, model categories, left-induced model structures, left Bousfield localization.
+1991 Mathematics Subject Classification:
+55P62 (Primary), 55P60, 55U10, 55U35, 18N55, 18N50, 18N40 (Secondary).
 1. Introduction
-Rational homotopy theory studies $1$-connected spaces by simplifying the computation of their homotopy groups.
+Classically, rational homotopy theory is the study of $1$-connected spaces up to the torsion in their homotopy groups.
 The non-vanishing homotopy groups of a $1$-connected space are abelian, so their torsion can be discarded by tensoring with the rational numbers.
-Discarding torsion makes homotopy group computations more tractable: in the case of spheres, their rational homotopy groups were computed by Serre in [Ser51].
-The success of rational homotopy theory has led to its extension to nilpotent spaces by Bousfield and Kan in [BK72]; Sullivan in [Sul74] and [Sul77]; Hilton, Mislin, and Roitberg in [HMR75]; Bousfield and Gugenheim in [BG76]; and Neisendorfer in [Nei78].
-Ultimately, rational homotopy theory has been extended to all spaces by Casacuberta and Peschke in [CP93]; Gómez-Tato, Halperin, and Tanré in [GTHT00]; and Bastardas and Casacuberta in [BC01].
-Model categories, as introduced by Quillen in [Qui67], provide a formal framework in which many tools and constructions from homotopy theory are defined.
-For instance, every model category admits a homotopy category, which is the localization of the model category at its class of weak equivalences.
-Thus, a model structure on a category endows the category with a homotopy theory and gives us tools to study that homotopy theory.
-In [Qui69], Quillen applies his recently introduced notion of a model category to study rational homotopy theory, and especially how rational homotopy theory translates difficult problems in geometry and topology to more feasible problems in algebra.
-More specifically, Quillen produces model categories both on the topological side and on the algebraic side, and then proves that specific functorial constructions between them induce equivalences of these model categories.
-The notion of equivalence that we are referring to is that of a Quillen equivalence; in particular, a Quillen equivalence induces an equivalence of homotopy categories.
-On the algebraic side, Quillen produces model structures on rational differential graded (DG) Lie algebras and rational DG coalgebras, which are important algebraic invariants for computations in rational homotopy theory.
-In this paper, we focus on the topological side, where Quillen produces a family of model structures on the category of $1$-reduced simplicial sets: simplicial sets whose $1$-skeleton is a point.
-As the category of $1$-connected spaces is ill-behaved with respect to important types of limits and colimits, such as pullbacks, the passage to $1$-reduced simplicial sets is necessary for Quillen’s model-categorical approach to rational homotopy theory.
-This family of model structures interpolates between ordinary homotopy theory, where the weak equivalences are weak homotopy equivalences, and rational homotopy theory, where the weak equivalences are rational homotopy equivalences.
-Namely, for a multiplicative subset $M$ of $\mathbb{Z}$, Quillen endows the category of $1$-reduced simplicial sets with a model structure in which a map is a weak equivalence if and only if it induces an isomorphism of homotopy groups after tensoring with $M^{-1}\mathbb{Z}$; we call such a map an $M$-local homotopy equivalence.
-Tensoring with $M^{-1}\mathbb{Z}$ for different choices of a multiplicative subset $M$ of $\mathbb{Z}$ allows us to preserve specific types of torsion in homotopy groups.
-For instance, for a prime $p$, taking $M=\mathbb{Z}\setminus p\mathbb{Z}$ and tensoring with the $p$-local integers $\mathbb{Z}_{(p)}$ specifically preserves $p$-torsion.
-Thus, we can study the torsion one prime at a time.
+In this framework, we adjust our notion of equivalence from that of a weak homotopy equivalence, which preserves all homotopy groups on the nose, to that of a rational homotopy equivalence, which induces isomorphisms of higher homotopy groups only after tensoring with the rational numbers.
+Conceptually, a rational homotopy equivalence can be thought of as a weak homotopy equivalence up to an error in torsion.
+In the $1$-connected setting, one can equivalently use rational homology instead; a map of $1$-connected spaces is a rational homotopy equivalence if and only if it is a rational homology isomorphism.
+Accordingly, the corresponding notion of a nice space in rational homotopy theory is a space whose higher homotopy groups are already rational vector spaces, that is, uniquely divisible by every positive integer; such spaces are called rational.
+In [39], Sullivan constructs, for a $1$-connected space $X$, a relative CW complex $(X_{\mathbb{Q}},X)$ such that $X_{\mathbb{Q}}$ is a rational space and the relative CW inclusion is a rational homotopy equivalence.
+This construction is called the rationalization of $X$, and it identifies the rational homotopy groups of $X$ with the higher homotopy groups of $X_{\mathbb{Q}}$ in a topological fashion.
+The geometric construction that underpins the rationalization of spaces is the rationalization of the $n$-sphere $S^{n}$ in dimensions $n\geq 2$ and, more specifically, the construction of the rational $n$-sphere $S^{n}_{\mathbb{Q}}$.
+In essence, the rationalization of a general space is built out of the CW inclusions of $n$-spheres into rational $n$-spheres, hence the resulting relative CW construction.
+Thus, the rationalization maps of spheres are the building blocks for the rational homotopy theory of spaces.
+Working rationally can simplify computations significantly.
+In the case of spheres, though their homotopy groups are notoriously difficult to compute, their rational homotopy groups were computed by Serre in [36].
+This contrast captures the fact that the difficulty in computing the homotopy groups of spheres is concentrated in their torsion.
+It also suggests that, while a direct computation of homotopy groups can be too ambitious an approach, the computation of localizations of homotopy groups can be a more tractable goal.
+1.1. Quillen’s approach to rational homotopy theory
+The computational features of rational homotopy theory extend deeper than the systematic disregard of torsion in homotopy groups.
+Rational homotopy theory can be translated purely into algebraic terms, and rational computations in geometry and topology can be transformed into more feasible ones in algebra.
+Quillen’s paper on rational homotopy theory [33] establishes this advantage of rational homotopy theory, and it informs us about the algebra that fully encodes the rational homotopy theory of $1$-connected spaces.
+Specifically, Quillen’s work turns the study of the rational homotopy theory of $1$-connected spaces into that of rational differential graded Lie algebras and rational differential graded coalgebras up to quasi-isomorphism.
+Quillen’s approach to rational homotopy theory takes place within his framework of model categories from [32].
+A model structure on a category allows one to introduce tools and constructions from homotopy theory in that category, and thus endows the category with a homotopy theory.
+Specifically, a model structure on a category comprises a class of morphisms called weak equivalences together with two other classes of morphisms, cofibrations and fibrations, subject to a list of axioms.
+This structure allows for constructions such as the homotopy category of a model category, which is its localization at the class of weak equivalences.
+In practice, model categories often satisfy additional properties that render the study of their homotopy theory less abstract and more explicit.
+In his paper on rational homotopy theory, Quillen produces a number of model structures, including
+  1. (1)
+    one on $1$-connected spaces whose weak equivalences are the rational homotopy equivalences;
+  2. (2)
+    one on rational differential graded Lie algebras whose weak equivalences are the quasi-isomorphisms; and
+  3. (3)
+    one on rational differential graded coalgebras whose weak equivalences are also the quasi-isomorphisms,
+as well as functorial constructions that induce equivalences of these model categories.
+The notion of equivalence that we are referring to is, in fact, called a Quillen equivalence.
+Thus, Quillen’s work quantifies how rational homotopy theory translates difficult problems in geometry and topology into more feasible problems in algebra.
+In other words, Quillen uses his framework of model categories to model the rational homotopy theory of $1$-connected spaces with the homotopy theory of rational differential graded Lie algebras and coalgebras.
+On the side of $1$-connected spaces, Quillen does not work so much with topological spaces as with simplicial sets.
+Simplicial sets can be thought of as simplicial complexes with degeneracy data that allows us to view lower-dimensional simplices as honorary higher-dimensional simplices, as well as directionality data for all simplices.
+This point of view illustrates that simplicial sets are a combinatorial model for spaces, and the degeneracy and directionality data render them a more faithful model than that of ordinary simplicial complexes.
+An exposition of this approach to simplicial sets and their advantages over simplicial complexes is provided by Bergner in [6].
+Instead of working directly with $1$-connected topological spaces, Quillen produces a family of model structures on the category of $1$-reduced simplicial sets, that is, simplicial sets whose $1$-skeleton is a point.
+As the category of $1$-connected topological spaces is ill-behaved with respect to important limits and colimits, such as pullbacks, the passage to $1$-reduced simplicial sets is necessary for Quillen’s model-categorical approach, which requires a category that has all finite limits and colimits.
+Thus, Quillen uses $1$-reduced simplicial sets as a combinatorial analog of $1$-connected spaces in analogy with how simplicial sets are a combinatorial model for spaces, except that, now, the more rigid structure of a $1$-reduced simplicial set yields a nicer underlying category to work in than that of $1$-connected spaces.
+More specifically, Quillen produces a family of model structures on $1$-reduced simplicial sets that interpolates between ordinary homotopy theory, whose weak equivalences are the weak homotopy equivalences, and rational homotopy theory, whose weak equivalences are the rational homotopy equivalences.
+This interpolation captures the fact that we have options between discarding no torsion at all, as is the case with weak homotopy equivalences, and discarding all torsion, as is the case with rational homotopy equivalences.
+For instance, given a prime number $p$, we can specifically preserve $p$-torsion in higher homotopy groups by tensoring them with the $p$-local integers $\mathbb{Z}_{(p)}$, and we can adapt our notion of equivalence to be that of a $p$-local homotopy equivalence, which induces isomorphisms of higher homotopy groups only after tensoring with the $p$-local integers.
+Such a selective approach allows us to study the torsion in homotopy groups one prime at a time rather than study it in one go or discard it altogether.
+In full detail, for a multiplicative subset $M$ of $\mathbb{Z}$, Quillen endows the category of $1$-reduced simplicial sets with a model structure in which a map is a weak equivalence if and only if it induces isomorphisms of higher homotopy groups after tensoring with $M^{-1}\mathbb{Z}$; we call such a map an $M$-local homotopy equivalence.
+Tensoring with $M^{-1}\mathbb{Z}$ for different choices of $M$ allows us to preserve specific types of torsion in homotopy groups.
+For instance, for a prime number $p$, choosing $M=\mathbb{Z}\setminus p\mathbb{Z}$ and tensoring with $\mathbb{Z}_{(p)}$ only keeps $p$-torsion.
+In this notation, a space is called $M$-local if its higher homotopy groups are $M^{-1}\mathbb{Z}$-modules, that is, uniquely divisible by every positive integer in $M$.
+Moreover, for a $1$-connected space $X$, Sullivan’s geometric construction yields a relative CW complex $(X_{M^{-1}\mathbb{Z}},X)$ such that $X_{M^{-1}\mathbb{Z}}$ is an $M$-local space and the relative CW inclusion is an $M$-local homotopy equivalence.
+This construction is called the $M$-localization of $X$.
+Thus, rational homotopy theory can be developed for any localization of $\mathbb{Z}$, though specific computations, such as Serre’s computation of the rational homotopy groups of spheres, much depend on the choice of such a localization.
 The following table depicts the interpolation given by varying our choice of a multiplicative subset $M$ of $\mathbb{Z}$.
-$M$$\{1\}$ $\mathbb{Z}\setminus 2\mathbb{Z}$$\mathbb{Z}\setminus\{0\}$
-[…]
-torsion discarded by $-\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}$ none all but $2$-torsion all
-$M$-local homotopy equivalence weak $2$-local rational
-To establish this family of model structures, Quillen verifies the model category axioms one-by-one.
-In later years, though, further advancements in the study of model categories have provided us with powerful tools that enable us to induce new model structures from old without needing to perform such a painstaking verification.
-The motivating idea behind our work in this paper is that we can use these modern tools to provide a new approach to Quillen’s family of model structures on $1$-reduced simplicial sets.
-The first tool we employ is that of left transfer, as developed by Bayeh, Hess, Karpova, Kędziorek, Riehl, and Shipley in [BHK+15].
+$M$ $\{1\}$ $\mathbb{Z}\setminus p\mathbb{Z}$ $\mathbb{Z}\setminus\{0\}$
+$M^{-1}\mathbb{Z}$ $\mathbb{Z}$ $\mathbb{Z}_{(p)}$ $\mathbb{Q}$
+torsion discarded by $\otimes M^{-1}\mathbb{Z}$ none all but $p$-torsion all
+$M$-local homotopy equivalence weak $p$-local rational
+1.2. A modern perspective on Quillen’s approach
+To establish his family of model structures on $1$-reduced simplicial sets, Quillen painstakingly verifies the model category axioms one-by-one.
+The motivating idea behind our work in this paper is that we can use modern tools to provide a streamlined approach to this family of model structures.
+The first tool we employ is that of left transfer, as developed by Bayeh, Hess, Karpova, Kędziorek, Riehl, and Shipley in [4].
 This tool allows us to transfer the model structure on pointed simplicial sets to a model structure on $1$-reduced simplicial sets whose weak equivalences are the weak homotopy equivalences.
-Then, we left Bousfield localize the latter model structure to obtain the rest of Quillen’s family.
-Although the tool of left Bousfield localization dates back to the work of Bousfield in [Bou75] and [Bou77] and Farjoun in [Far92] and [Far95b], we rely on the more recent existence results for left Bousfield localization developed by Hirschhorn in [Hir03] and by Barwick in [Bar10].
-For the left transfer, we observe that a $1$-reduced simplicial set has a unique 00-simplex, so there is a forgetful functor from the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets to the category ${\mathcal{SS}\mathrm{ets}}_{\ast}$ of pointed simplicial sets.
-This functor admits a right adjoint given by the $1$st Eilenberg subcomplex of a pointed simplicial set as defined by May in [May67]; in fact, Quillen uses this construction to establish his family of model structures.
-As the $1$st Eilenberg subcomplex depends on the basepoint, the passage to pointed simplicial sets, rather than unpointed ones, is necessary.
-We use this adjunction to left-transfer the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ to the following model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$.
-Theorem 1.1.
-There is a model structure on the category ${\mathcal{SS}\mathrm{ets}}_{1}$ in which a map $f$ is
+Then, we left Bousfield localize the latter model structure at the $M$-localization maps of spheres to recover the rest of Quillen’s family.
+Although the tool of left Bousfield localization dates back to the work of Bousfield [7, 8] and Farjoun [19, 21], we rely on the more recent existence results for left Bousfield localization developed by Hirschhorn in [26] and by Barwick in [2].
+For the left transfer, we observe that a $1$-reduced simplicial set has a unique 0-simplex, so there is a forgetful functor from the category of $1$-reduced simplicial sets to the category of pointed simplicial sets.
+This functor admits a right adjoint given by the 11st Eilenberg subcomplex of a pointed simplicial set, as defined by May in [30]; in fact, Quillen uses this construction to give his family of model structures.
+We use the 11st Eilenberg subcomplex functor to left-transfer the model structure on pointed simplicial sets to a model structure on $1$-reduced simplicial sets whose weak equivalences are the weak homotopy equivalences and whose fibrant objects are the Kan complexes.
+Then, for the left Bousfield localization, we characterize, for a multiplicative subset $M$ of $\mathbb{Z}$, the $M$-local spaces and the $M$-local homotopy equivalences using function complexes.
+Our characterizations can be viewed as the enriched-in-simplicial-sets version of the work of Casacuberta, Peschke, and Pfenniger in [15], which involves hom-sets instead.
+Using these characterizations, we left Bousfield localize our left-induced model category at the $M$-localization maps of spheres to obtain a model structure on $1$-reduced simplicial sets whose weak equivalences are the $M$-local homotopy equivalences and whose fibrant objects are the $M$-local Kan complexes.
+Thus, we recover the entirety of Quillen’s family of model structures on $1$-reduced simplicial sets.
+In particular, we obtain a model structure on the category of $1$-reduced simplicial sets in which the weak equivalences are the rational homotopy equivalences and the fibrant objects are the rational Kan complexes.
+This model category encapsulates the rational homotopy theory of $1$-connected spaces; for instance, fibrant replacements in this model category are rationalizations of $1$-connected spaces.
+It is this model category that Quillen establishes to be equivalent to his model structures on rational differential graded Lie algebras and rational differential graded coalgebras.
+We also give an alternative approach using Bousfield localization with respect to homology, as introduced by Bousfield in [7], in conjunction with left transfer.
+The idea is that, for a multiplicative subset $M$ of $\mathbb{Z}$, rather than first left-transfer the model structure on pointed simplicial sets and then localize, we can try to localize the model structure on pointed simplicial sets with respect to homology with $M^{-1}\mathbb{Z}$-coefficients and then left-transfer this localized model structure.
+We show that this approach also recovers Quillen’s family of model structures on $1$-reduced simplicial sets.
+Schematically, for a multiplicative subset $M$ of $\mathbb{Z}$, we denote
   1. (1)
-    a weak equivalence if and only if $f$ is a weak homotopy equivalence; and
+    the model category of pointed simplicial sets by $\mathcal{SS}\mathrm{ets}$$_{\ast}$};
   2. (2)
-    a cofibration if and only if $f$ is a monomorphism.
-Then, for the left Bousfield localization, we characterize, for a multiplicative subset $M$ of $\mathbb{Z}$, the $M$-local spaces and the $M$-local homotopy equivalences in the language of function complexes.
-Our characterizations can be viewed as the enriched-in-simplicial-sets version of the work of Casacuberta, Peschke, and Pfenniger in [CPP92], which uses hom-sets rather than function complexes.
-Using these characterizations, we left Bousfield localize the model category from Theorem 1.1 to recover the entire family of model structures on ${\mathcal{SS}\mathrm{ets}}_{1}$ from [Qui69].
-We say that a simplicial set $X$ is $M$-local if $X$ is a Kan complex whose higher homotopy groups are $M^{-1}\mathbb{Z}$-modules.
-Theorem 1.2.
-Let $M$ be a multiplicative subset of $\mathbb{Z}$.
-There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ on the category of $1$-reduced simplicial sets in which a map $f$ is
+    the model structure on $1$-reduced simplicial sets with weak homotopy equivalences by $\mathcal{SS}\mathrm{ets}_{1}$};
   3. (3)
-    a weak equivalence if and only if $f$ is an $M$-local homotopy equivalence; and
+    the localization of $\mathcal{SS}\mathrm{ets}$$_{\ast}$} with respect to homology with $M^{-1}\mathbb{Z}$-coefficients by $\mathcal{L}_{HM^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{\ast}$}; and
   4. (4)
-    a cofibration if and only if $f$ is a monomorphism.
-A $1$-reduced simplicial set $X$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $X$ is $M$-local.
-We also give an alternative description of this family of model structures using Bousfield localization with respect to homology, as introduced in [Bou75], in conjunction with the result on left-induced model structures from [BHK+15] that we use for Theorem 1.1.
-The idea is that, rather than first left-transfer the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ and then localize, we can try to localize the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ with respect to homology with $M^{-1}\mathbb{Z}$-coefficients and then left-transfer this localized model structure, which we denote by $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$, to ${\mathcal{SS}\mathrm{ets}}_{1}$.
-We show that this approach also yields the family of model structures from Theorem 1.2.
-Schematically, the diagram of model-categorical operations
-[…]
-commutes, thus giving us two modern descriptions of the family of model structures on ${\mathcal{SS}\mathrm{ets}}_{1}$.
-Another motivating goal of ours is to show that these model structures have properties that one generally desires and expects in the context of model categories.
+    the model structure on $1$-reduced simplicial sets with $M$-local homotopy equivalences by $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{1}$},
+and we show that the diagram of model-categorical operations
+$\mathcal{SS}\mathrm{ets}$$_{\ast}$$\mathcal{SS}\mathrm{ets}_{1}$$\mathcal{L}_{HM^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{\ast}$$\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{1}$}}localizationtransferlocalizationtransfer
+commutes.
+As the weak equivalences in $\mathcal{L}_{HM^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{\ast}$} are the $M$-local homology isomorphisms, the commutativity of the diagram stems from the result that a map of $1$-connected spaces is an $M$-local homotopy equivalence if and only if it is an $M$-local homology isomorphism.
+Another motivating goal of ours is to show that these model structures on $1$-reduced simplicial sets have properties that one generally desires and expects in the context of model categories.
 The tools that we invoke guarantee us most of these properties.
-An exception is the property of being a simplicial model category, as defined by Quillen in [Qui67].
-We establish this property explicitly by adapting the techniques of Goerss and Jardine in [GJ99] to show that the model category from Theorem 1.1 is simplicial.
-As left Bousfield localization preserves the simplicial structure, this structure is bestowed upon the entire family of model structures on $1$-reduced simplicial sets from Theorem 1.2.
-We remark that, in [Bur21], Burke proves the analog of our simplicial structure for reduced simplicial sets: simplicial sets with a unique 00-simplex.
-However, Burke’s line of argument is different from that in this paper.
-Also, we partially extend Theorem 1.2 to the category $\mathcal{SS}\mathrm{ets}$ of simplicial sets.
-Namely, we produce a family of model structures on simplicial sets that describes a generalization of rational homotopy theory to all spaces.
-In [Iva22], Ivanov gives an overview of the various approaches to rational homotopy theory beyond the framework of $1$-connectedness or nilpotency.
-We adopt the approach of Gómez-Tato, Halperin, and Tanré from [GTHT00].
-In this approach, one rationalizes a connected space $X$ by applying fiberwise rationalization to the fibration
+An exception is the property of being a simplicial model category, as defined by Quillen in [32].
+We establish this property explicitly by first showing that our left-induced model structure on $1$-reduced simplicial sets is simplicial.
+As left Bousfield localization preserves the simplicial structure, this structure is bestowed upon the entire family of model structures on $1$-reduced simplicial sets.
+1.3. Generalizing Quillen’s family of model structures to all spaces
+The modern tools that we use to recover Quillen’s family of model structures on $1$-reduced simplicial sets allow us to extend it to all spaces.
+This extension encodes the generalization of rational homotopy theory to non-simply connected spaces, which has been accomplished in two stages.
+Firstly, rational homotopy theory was extended to nilpotent spaces by Sullivan [37, 38, 39]; Bousfield and Kan [11]; Hilton, Mislin, and Roitberg [25]; Bousfield and Gugenheim [10]; and Neisendorfer [31].
+Ultimately, it was extended to all spaces by Casacuberta and Peschke [14]; Gómez-Tato, Halperin, and Tanré [24]; and Bastardas and Casacuberta [3].
+At both stages of generalization, rational homotopy theory has been translated to purely algebraic terms, as in the classical case.
+We use our modern perspective to produce a family of model structures on the category of spaces that describes the generalization of rational homotopy theory to non-simply connected spaces due to Gómez-Tato, Halperin, and Tanré.
+In their approach, one rationalizes a connected space $X$ by applying fiberwise rationalization to the fibration
 $\widetilde{X}\rightarrow X\rightarrow K(\pi_{1}(X),1)$
-given by the $1$st Postnikov approximation of $X$; its homotopy fiber $\widetilde{X}$ is the universal cover of $X$, which is a $1$-connected space that can be rationalized in the classical sense.
-The tool of fiberwise localization is invoked by Sullivan in [Sul74] and investigated by Llerena in [Lle82] and [Lle85] and Farjoun in [Far95a].
-This rationalization theory for non-nilpotent spaces is also studied in recent work of Rivera, Wierstra, and Zeinalian in [RWZ21].
-We partially extend Theorem 1.2 to all simplicial sets as follows.
-Theorem 1.3.
-Let $M$ be a multiplicative subset of $\mathbb{Z}$.
-There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ on the category of simplicial sets in which a simplicial set $X$ is fibrant if and only if $X$ is $M$-local.
-We also prove the counterpart of this theorem for the category ${\mathcal{T}\mathrm{op}}$ of topological spaces.
-The reason why Theorem 1.3 only partially extends Theorem 1.2 to all simplicial sets is that we lack a description of the weak equivalences other than the standard one for a left Bousfield localization.
-This deficiency is due to the fact that fiberwise rationalization satisfies a more intricate universal property than that of the rationalization of $1$-connected spaces, which causes our characterization of the weak equivalences in Theorem 1.2 to break down.
-Lastly, we verify that our localization $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ of spaces with respect to rational homotopy from Theorem 1.3 does not coincide with Bousfield’s localization of spaces with respect to rational homology from [Bou75].
-This result captures the fact that rational homotopy and rational homology do not agree beyond the framework of $1$-connected spaces; in fact, our proof of the result involves a specific instance of this general phenomenon.
-Organization of the paper
+given by the 11st Postnikov approximation of $X$, where $K(\pi_{1}(X),1)$ is an Eilenberg-Mac Lane space.
+The fiber $\widetilde{X}$ is the universal cover of $X$, which is a $1$-connected space that can be rationalized in the classical sense.
+This technique generalizes to disconnected spaces by fiberwise rationalizing every component.
+The tool of fiberwise localization is invoked by Sullivan [37] and investigated by Llerena [28, 29] and Farjoun [20], and the specific case of fiberwise rationalization is treated in recent work of Ivanov [27].
+In general, for a multiplicative subset $M$ of $\mathbb{Z}$, the fiberwise $M$-localization of a connected space $X$ entails, at the level of total spaces, a map $X\rightarrow X_{M^{-1}\mathbb{Z}}$ that induces isomorphisms on $\pi_{0}$ and $\pi_{1}$ and isomorphisms of higher homotopy groups after tensoring with $M^{-1}\mathbb{Z}$.
+We call such a map an $M$-local homotopy equivalence, for it extends the notion of an $M$-local homotopy equivalence beyond maps of $1$-connected spaces.
+In [34], Rivera, Wierstra, and Zeinalian give algebraic characterizations for rational homotopy equivalences in this general sense, which they refer to as $\pi_{1}$-rational homotopy equivalences.
+Also, as in the $1$-connected case, a space is called $M$-local if its higher homotopy groups are $M^{-1}\mathbb{Z}$-modules; note that no conditions are imposed on $\pi_{0}$ or $\pi_{1}$.
+In this terminology, we endow the category of spaces with a model structure whose weak equivalences are the $M$-local homotopy equivalences and whose fibrant objects are the $M$-local spaces.
+This model category is the left Bousfield localization of the model category of spaces at the $M$-localization maps of spheres.
+In particular, we get a model structure on the category of spaces in which the weak equivalences are the rational homotopy equivalences and the fibrant objects are the rational spaces.
+In [24], Gómez-Tato, Halperin, and Tanré produce an algebraic category with a notion of homotopy between its morphisms whose homotopy category is equivalent to the homotopy category of rational spaces in their extended sense.
+By our theorem, the latter homotopy category arises from a model category, so it can be studied using the structured toolkit of model categories.
+Lastly, we note that our localization of all spaces at rational homotopy does not coincide with Bousfield’s localization of spaces at rational homology; the unique map from $\mathbb{R}P^{2}$ to the point is a rational homology isomorphism, but not a rational homotopy equivalence in the extended sense, for it does not induce an isomorphism of fundamental groups.
+This result captures the fact that rational homotopy and rational homology do not agree for non-simply connected spaces.
+1.4. Organization of the paper
 Section 2 comprises background on $1$-reduced simplicial sets.
-In Section 3, we perform the left transfer to obtain Theorem 1.1.
-In Section 4, we prove some results about the fibrations in this left-induced model structure.
-In Section 5, we establish that the model category from Theorem 1.1 is simplicial.
+In Section 3, we perform the left transfer.
+In Section 4, we investigate the fibrations in our left-induced model category.
+In Section 5, we establish that our left-induced model category is simplicial.
 Section 6 consists of auxiliary characterization results from the rational homotopy theory of $1$-connected spaces.
-We use these results in Section 7 to perform the left Bousfield localization that yields Theorem 1.2.
-In Section 8, we give an alternative approach to Theorem 1.2 via Bousfield localization with respect to homology.
-In Section 9, we give an overview of rational homotopy theory beyond the setting of $1$-connectedness.
-In Section 10, we perform the left Bousfield localization that yields Theorem 1.3.
-In Section 11, we establish the analog of Theorem 1.3 for topological spaces.
-Lastly, in Section 12, we show that our localization of spaces with respect to rational homotopy is not equal to the Bousfield localization of spaces with respect to rational homology.
-Acknowledgments
+We use these results in Section 7 to perform the left Bousfield localization that recovers Quillen’s family of model structures on $1$-reduced simplicial sets.
+In Section 8, we give an alternative approach via Bousfield localization with respect to homology.
+In Section 9, we give an overview of the rational homotopy theory of non-simply connected spaces.
+Lastly, in Section 10, we perform our left Bousfield localization to all spaces to generalize Quillen’s family of model structures.
+1.5. Acknowledgments
 I am grateful to my PhD advisor at the University of Virginia, Julie Bergner, for her guidance, input, feedback, and support that made the work in this paper possible.
@@
 We begin with some background on $1$-reduced simplicial sets.
-We define the $1$-skeleton and the $1$-coskeleton of a simplicial set.
+We first give background on simplicial sets; in particular, we define the $1$-skeleton and the $1$-coskeleton of a simplicial set.
 Using these notions, we define $1$-reduced simplicial sets.
-Then, we define the $1$-reduction of a simplicial set, as well as the $1$st Eilenberg subcomplex of a pointed simplicial set.
-These constructions provide us with adjunctions between $1$-reduced simplicial sets and (pointed) simplicial sets that we rely on in this paper.
+Then, we define the $1$-reduction of a simplicial set, as well as the 11st Eilenberg subcomplex of a pointed simplicial set.
+These constructions provide us with adjunctions between $1$-reduced simplicial sets and (pointed) simplicial sets that we rely on to bridge our investigation of the homotopy theory of $1$-reduced simplicial sets to known results from the homotopy theory of (pointed) simplicial sets.
 2.1. Simplicial sets
-We collect background and notation on simplicial sets; a reference is [GJ99, Chapter I].
+We collect some background and notation on simplicial sets; a reference is [23, Chapter I].
 The combinatorics involved in the definition of a simplicial set can be packaged in the simplex category $\Delta$.
 Definition 2.1.
-The simplex category $\Delta$ is the category whose
-  1. (1)
-    objects are the sets $[n]=\{0,1,\dots,n\}$ for a non-negative integer $n$; and
-  2. (2)
-    morphisms are the order-preserving maps $\theta\colon[n]\rightarrow[m]$.
+The simplex category $\Delta$ is the category whose objects are the sets $n]=\{0,1,\dots,n\}$ for a non-negative integer $n$ and whose morphisms are the order-preserving maps of these sets.
+The order-preserving maps of finite total orders can be built out of special injections and surjections.
 Example 2.2.
-For $0\leq i\leq n$, two special types of maps in $\Delta$ are
-  1. (1)
-    the $i$-th coface map $d^{i}\colon[n-1]\rightarrow[n]$; and
-  2. (2)
-    the $i$-th codegeneracy map $s^{i}\colon[n+1]\rightarrow[n]$.
-These maps generate the simplex category $\Delta$ subject to a list of identities that determine their composites.
-The simplex category $\Delta$ allows for a concise definition of a simplicial set. We write ${\mathcal{S}\mathrm{ets}}$ for the category of sets.
+For $0\leq i\leq n$, two special maps in $\Delta$ are the $i$-th coface map $d^{i}\colon[n-1]\rightarrow[n]$, given by
+$d^{i}(x)=\begin{cases}x,&x<i;\\ x+1,&x\geq i,\end{cases}$
+and the $i$-th codegeneracy map $s^{i}\colon[n+1]\rightarrow[n]$, given by
+$s^{i}(x)=\begin{cases}x,&x\leq i;\\ x-1,&x>i.\end{cases}$
+These maps generate $\Delta$ subject to a list of identities on their composites.
+The simplex category $\Delta$ yields a concise definition of a simplicial set; we write $\mathcal{S}\mathrm{ets}$} for the category of sets.
 Definition 2.3.
-A simplicial set is a functor $X\colon\Delta^{\operatorname{op}}\rightarrow{\mathcal{S}\mathrm{ets}}$. In detail, a simplicial set comprises
-  1. (1)
-    a set of $n$-simplices $X_{n}$ for every non-negative integer $n$;
-  2. (2)
-    face maps $d_{i}\colon X_{n}\rightarrow X_{n-1}$ for $0\leq i\leq n$; and
-  3. (3)
-    degeneracy maps $s_{i}\colon X_{n}\rightarrow X_{n+1}$ for $0\leq i\leq n$;
-subject to a list of simplicial identities that govern how the faces and degeneracies are composed.
-We write ${\mathcal{SS}\mathrm{ets}}$ for the category of simplicial sets and simplicial maps (natural transformations) between them.
+A simplicial set is a functor $X\colon\Delta^{\operatorname{op}}\rightarrow{\mathcal{S}\mathrm{ets}}$.
+In detail, a simplicial set comprises a set of $n$-simplices $X_{n}$ for every non-negative integer $n$, face maps $d_{i}\colon X_{n}\rightarrow X_{n-1}$ for $0\leq i\leq n$, and degeneracy maps $s_{i}\colon X_{n}\rightarrow X_{n+1}$ for $0\leq i\leq n$ subject to a list of identities on how the faces and degeneracies are composed.
+We write $\mathcal{SS}\mathrm{ets}$} for the category of simplicial sets and their simplicial maps (natural transformations).
+We give the fundamental building block examples of simplicial sets, which are reminiscent of simplicial complexes.
 Example 2.4.
-The standard $n$-simplex is the representable simplicial set
-$\Delta[n]=\operatorname{Hom}_{\Delta}\left(-,[n]\right).$
-The identity $\operatorname{id}_{[n]}$ gives an $n$-simplex $\sigma_{n}$ of $\Delta[n]$. For a simplicial set $X$, the Yoneda lemma yields the bijection
-$X_{n}\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}}\left(\Delta[n],X\right)$
-that is natural in $[n]$ and in $X$. Therefore, we occasionally view an $n$-simplex of $X$ as a map $\Delta[n]\rightarrow X$.
-Example 2.5.
-The boundary $\partial\Delta[n]$ is the smallest subcomplex of $\Delta[n]$ containing the faces $d_{i}(\sigma_{n})$ for $0\leq i\leq n$. We write
-$i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]$
-for the boundary inclusion.
-Example 2.6.
-For $n\geq 1$ and $0\leq k\leq n$, the horn $\Lambda[n,k]$ is the smallest subcomplex of $\Delta[n]$ containing $d_{i}(\sigma_{n})$ for $i\neq k$. We write
-$j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]$
-for the horn inclusion. Observe that horns are undefined for $n=0$.
-We can geometrically realize a simplicial set $X$ to obtain a CW complex $|X|$.
-Writing ${\mathcal{T}\mathrm{op}}$ for the category of topological spaces, we denote the geometric realization functor by
-$|-|\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{T}\mathrm{op}}.$
-Conversely, for a topological space $Y$, its singular complex $\operatorname{Sing}(Y)$ is the simplicial set with $n$-simplices
-$\operatorname{Sing}(Y)_{n}=\operatorname{Hom}_{\mathcal{T}\mathrm{op}}\left(|\Delta[n]|,Y\right).$
-This pair of functors forms an adjunction that we use to oscillate between simplicial sets and topological spaces.
-Proposition 2.7 ([GJ99, Chapter I, Proposition 2.2]).
-The geometric realization functor $|-|\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{T}\mathrm{op}}$ is left adjoint to the singular complex functor $\operatorname{Sing}\colon{\mathcal{T}\mathrm{op}}\rightarrow{\mathcal{SS}\mathrm{ets}}$.
-Our work largely involves pointed simplicial sets, which are simplicial sets with a designated 00-simplex as their basepoint.
-We write ${\mathcal{SS}\mathrm{ets}}_{\ast}$ for the category of pointed simplicial sets and pointed simplicial maps between them.
-The following adjunction is verified just like its analog in topological spaces.
-Lemma 2.8.
-The forgetful functor $V\colon{\mathcal{SS}\mathrm{ets}}_{\ast}\rightarrow{\mathcal{SS}\mathrm{ets}}$ has a left adjoint
-$(-)_{+}\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$
-defined by attaching a disjoint basepoint.
+The standard $n$-simplex is $\Delta[n]=\operatorname{Hom}_{\Delta}(-,[n]);$ the identity $\operatorname{id}_{[n]}$ gives an $n$-simplex $\sigma_{n}$.
+The boundary $\partial\Delta[n]$ is the smallest subcomplex of $\Delta[n]$ containing $d_{i}(\sigma_{n})$ for $0\leq i\leq n$.
+For $n\geq 1$ and $0\leq k\leq n$, the horn $\Lambda[n,k]$ is the smallest subcomplex of $\Delta[n]$ containing $d_{i}(\sigma_{n})$ for $i\neq k$.
+As is the case for simplicial complexes, we can geometrically realize a simplicial set $X$ to obtain a CW complex $|X|$.
+Writing $\mathcal{T}\mathrm{op}$} for the category of topological spaces, we denote the geometric realization functor by $|-|\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{T}\mathrm{op}}$.
+Its right adjoint sends a topological space $Y$ to the singular complex $\operatorname{Sing}(Y)$ with $n$-simplices $\operatorname{Sing}(Y)_{n}=\operatorname{Hom}_{\mathcal{T}\mathrm{op}}(|\Delta[n]|,Y)$.
+As the name suggests, there is a connection between the singular complex of a space and its singular homology; the $n$-th abelian group in the singular chain complex of a space is defined to be the free abelian group on the set of $n$-simplices of the singular complex of that space.
 2.2. The $1$-skeleton and the $1$-coskeleton of a simplicial set
 To define $1$-reduced simplicial sets, we recall the $1$-skeleton of a simplicial set.
-We also recall the $1$-coskeleton of a simplicial set, which we refer to in the definition of the $1$st Eilenberg subcomplex of a pointed simplicial set.
-Our reference is [Ber18, Chapter 2.6].
-Let $\Delta_{1}$ denote the full subcategory of the simplex category $\Delta$ with objects $[0]$ and $[1]$. The inclusion
-$\Delta_{1}\rightarrow\Delta$
-induces a truncation functor
-$\operatorname{tr}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{S}\mathrm{ets}}^{\Delta_{1}^{\operatorname{op}}}.$
-This truncation functor admits both a left adjoint
-$s_{1}\colon{\mathcal{S}\mathrm{ets}}^{\Delta_{1}^{\operatorname{op}}}\rightarrow{\mathcal{SS}\mathrm{ets}}$
-and a right adjoint
-$c_{1}\colon{\mathcal{S}\mathrm{ets}}^{\Delta_{1}^{\operatorname{op}}}\rightarrow{\mathcal{SS}\mathrm{ets}},$
-in terms of which we can define the $1$-skeleton and the $1$-coskeleton of a simplicial set.
+We also recall the $1$-coskeleton of a simplicial set, which we refer to in the definition of the 11st Eilenberg subcomplex of a pointed simplicial set.
+Our reference is [5, Chapter 2.6].
+Let $\Delta_{1}$ denote the full subcategory of $\Delta$ with objects $[0]$ and $[1]$.
+The inclusion $\Delta_{1}\rightarrow\Delta$ induces a truncation functor $\operatorname{tr}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{S}\mathrm{ets}}^{\Delta_{1}^{\operatorname{op}}}$ that admits both a left adjoint $s_{1}\colon{\mathcal{S}\mathrm{ets}}^{\Delta_{1}^{\operatorname{op}}}\rightarrow{\mathcal{SS}\mathrm{ets}}$ and a right adjoint $c_{1}\colon{\mathcal{S}\mathrm{ets}}^{\Delta_{1}^{\operatorname{op}}}\rightarrow{\mathcal{SS}\mathrm{ets}}$, in terms of which we can define the $1$-skeleton and the $1$-coskeleton of a simplicial set.
 Definition 2.5.
 Let $X$ be a simplicial set.
-  1. (1)
-    The $1$-skeleton of $X$ is the simplicial set
-    $\operatorname{sk}_{1}(X)=s_{1}\operatorname{tr}_{1}(X).$
-  2. (2)
-    The $1$-coskeleton of $X$ is the simplicial set
-    $\operatorname{cosk}_{1}(X)=c_{1}\operatorname{tr}_{1}(X).$
-In detail, both $\operatorname{sk}_{1}(X)$ and $\operatorname{cosk}_{1}(X)$ have set of 00-simplices $X_{0}$ and set of $1$-simplices $X_{1}$, but they differ in their higher-dimensional simplices.
+The $1$-skeleton of $X$ is $\operatorname{sk}_{1}(X)=s_{1}\operatorname{tr}_{1}(X)$.
+The $1$-coskeleton of $X$ is $\operatorname{cosk}_{1}(X)=c_{1}\operatorname{tr}_{1}(X)$.
+In detail, both $\operatorname{sk}_{1}(X)$ and $\operatorname{cosk}_{1}(X)$ have set of 0-simplices $X_{0}$ and set of $1$-simplices $X_{1}$, but they differ in their higher-dimensional simplices.
 The higher-dimensional simplices of $\operatorname{sk}_{1}(X)$ are the iterated degeneracies of its $1$-simplices.
-On the contrary, $\operatorname{cosk}_{1}(X)$ has an $n$-simplex for every compatible collection of $(n-1)$-simplices.
-Equivalently, $\operatorname{cosk}_{1}(X)$ has an $n$-simplex for every map $\operatorname{sk}_{1}(\Delta[n])\rightarrow X$.
+On the contrary, $\operatorname{cosk}_{1}(X)$ has an $n$-simplex for every compatible collection of (n1)(n-1)-simplices.
+More precisely, $\operatorname{cosk}_{1}(X)$ has an $n$-simplex for every map $\operatorname{sk}_{1}(\Delta[n])\rightarrow X$.
 Remark 2.6.
-It follows from this definition that, for a simplicial set $X$, we have an inclusion map
-$\operatorname{sk}_{1}(X)\rightarrow X$
-and a quotient map
-$X\rightarrow\operatorname{cosk}_{1}(X),$
-where the latter map identifies the simplices of $X$ that share the same $1$-skeleton.
+It follows that, for a simplicial set $X$, we have an inclusion map $\operatorname{sk}_{1}(X)\rightarrow X$ and a canonical map $X\rightarrow\operatorname{cosk}_{1}(X)$, where the latter map identifies the simplices of $X$ that share the same $1$-skeleton.
+Besides the above remark, we also need the following example for later use in this section.
 Example 2.7.
 Observe that the $1$-skeleton of the standard $2$-simplex $\Delta[2]$ is its boundary $\partial\Delta[2]$.
-2.3. $1$-reduced simplicial sets
-We are now ready to define $1$-reduced simplicial sets.
+2.3. Definition of $1$-reduced simplicial sets
+We are now ready to introduce $1$-reduced simplicial sets.
 Definition 2.8.
 A simplicial set $X$ is $1$-reduced if its $1$-skeleton $\operatorname{sk}_{1}(X)$ is a point.
-In detail, a simplicial set $X$ is $1$-reduced if it has a unique 00-simplex whose degeneracy is the unique $1$-simplex of $X$.
+In detail, a simplicial set $X$ is $1$-reduced if it has a unique 0-simplex whose degeneracy is the unique $1$-simplex of $X$.
 Remark 2.9.
 Note that the geometric realization of a $1$-reduced simplicial set is a $1$-connected CW complex.
-It is also crucial to observe that a $1$-reduced simplicial set is unambiguously pointed, for it has a unique 00-simplex.
+It is also crucial to observe that a $1$-reduced simplicial set is unambiguously pointed, for it has a unique 0-simplex.
 We write ${\mathcal{SS}\mathrm{ets}}_{1}$ for the category of $1$-reduced simplicial sets and simplicial maps between them.
+The most important example of a $1$-reduced simplicial set for our work is the following.
 Example 2.10.
-For $n\geq 2$, the simplicial $n$-sphere
-$S_{\mathrm{simp}}^{n}=\Delta[n]/\partial\Delta[n]$
-is a $1$-reduced simplicial set consisting of a unique 00-simplex and a unique non-degenerate $n$-simplex.
-The proof of the following lemma is straightforward. We write $\wedge$ for the smash product.
+For $n\geq 2$, the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}=\Delta[n]/\partial\Delta[n]$ is a $1$-reduced simplicial set consisting of a unique 0-simplex and a unique non-degenerate $n$-simplex.
+We also need the following lemma, whose proof is straightforward; we write $\wedge$ to denote the smash product.
 Lemma 2.11.
@@
 This left adjoint quotients out the $1$-skeleton of a simplicial set to yield a $1$-reduced simplicial set.
-Although quotients can be homotopically ill-behaved, this adjunction aids our model-categorical arguments later in this paper, for it repeatedly lets us translate lifting problems that we wish to solve into lifting problems that we can solve.
+Although quotients can be homotopically ill-behaved, this adjunction aids our model-categorical arguments in this paper, for it lets us translate lifting problems that we wish to solve in ${\mathcal{SS}\mathrm{ets}}_{1}$ into lifting problems that we can solve in $\mathcal{SS}\mathrm{ets}$}.
 Definition 2.12.
-The $1$-reduction of a simplicial set $X$ is the $1$-reduced simplicial set
-$R_{1}(X)=X/\operatorname{sk}_{1}(X).$
+The $1$-reduction of a simplicial set $X$ is the $1$-reduced simplicial set $R_{1}(X)=X/\operatorname{sk}_{1}(X)$.
+The following observation carries useful consequences for later, so we formulate it explicitly.
 Example 2.13.
 The simplicial $2$-sphere is $S_{\mathrm{simp}}^{2}=R_{1}(\Delta[2])$ because, by Example 2.7, $\operatorname{sk}_{1}(\Delta[2])=\partial\Delta[2]$.
-The $1$-reduction of a simplicial set is its largest $1$-reduced quotient in the sense of the following universal property, which is an instance of the universal property of quotients.
-Lemma 2.18.
+The $1$-reduction of a simplicial set is its largest $1$-reduced quotient in the sense of the following universal property, which is a consequence of the universal property of quotients.
+Proposition 2.14.
 For every map $f\colon X\rightarrow Y$ from a simplicial set $X$ to a $1$-reduced simplicial set $Y$, there exists a unique map $\overline{f}\colon R_{1}(X)\rightarrow Y$ such that the diagram
-${X}$${{R_{1}(X)}}$${{Y}}$$\scriptstyle{f}$$\scriptstyle{\overline{f}}$
+[…]
 commutes, where $X\rightarrow R_{1}(X)$ is the quotient map from $X$ to $R_{1}(X)=X/\operatorname{sk}_{1}(X)$.
 This universal property translates to the promised adjunction.
-Lemma 2.19.
-The $1$-reduction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$ is left adjoint to the forgetful functor $U\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}$.
-We note a consequence of this adjunction for later use.
+Proposition 2.15.
+The $1$-reduction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$ is left adjoint to the forgetful functor.
+We invoke Example 2.13 to note a consequence of this adjunction for later use.
 Corollary 2.16.
@@
 Proof.
-The claim is a consequence of Lemma 2.19 because $S_{\mathrm{simp}}^{2}=R_{1}(\Delta[2])$ by Example 2.17. ∎
-2.5. The $1$st Eilenberg subcomplex of a pointed simplicial set
+The claim is a consequence of Proposition 2.15 because $S_{\mathrm{simp}}^{2}=R_{1}(\Delta[2])$ by Example 2.13. ∎
+2.5. The 11st Eilenberg subcomplex of a pointed simplicial set
 As we shall see, the approach of taking the $1$-reduction of a simplicial set does not play as nicely as we need it to for our homotopy-theoretic purposes, but we can do better.
-Because a $1$-reduced simplicial set has a unique 00-simplex, there is a forgetful functor
-$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$
-from the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets to the category ${\mathcal{SS}\mathrm{ets}}_{\ast}$ of pointed simplicial sets.
-This functor has not only a left adjoint given by $1$-reduction, but also a right adjoint given by the $1$st Eilenberg subcomplex of a pointed simplicial set.
-We present its definition, though we do not use it beyond this section, as its only feature that we need in this paper is the aforementioned adjunction, which we establish.
-Recall from Remark 2.10 that, for a simplicial set $X$, we have a quotient map $X\rightarrow\operatorname{cosk}_{1}(X)$.
-Definition 2.21 ([May67, Definition 8.3]).
-The $1$st Eilenberg subcomplex $E_{1}(X)$ of a pointed simplicial set $X$ with basepoint $x_{0}$ is defined by the pullback square
-${{E_{1}(X)}}$${X}$${{\Delta[0]}}$${{\operatorname{cosk}_{1}(X).}}$$\scriptstyle{x_{0}}$
-In detail, $E_{1}(X)$ consists of all simplices $\chi$ of $X$ such that every 00-dimensional face of $\chi$ is $x_{0}$ and every $1$-dimensional face of $\chi$ is the degeneracy of $x_{0}$.
-It follows from this definition that the $1$st Eilenberg subcomplex $E_{1}(X)$ of a pointed simplicial set $X$ is $1$-reduced.
-The $1$st Eilenberg subcomplex of a pointed simplicial set is its largest $1$-reduced subcomplex in the sense of the following universal property that is straightforward to verify.
-Lemma 2.22.
-Every pointed map $f\colon X\rightarrow Y$ from a $1$-reduced simplicial set $X$ to a pointed simplicial set $Y$ maps into the $1$st Eilenberg subcomplex $E_{1}(Y)$ of $Y$.
+Because a $1$-reduced simplicial set has a unique 0-simplex, there is a forgetful functor $F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$ from the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets to the category $\mathcal{SS}\mathrm{ets}$$_{\ast}$ of pointed simplicial sets.
+This functor also has a right adjoint given by the 11st Eilenberg subcomplex of a pointed simplicial set, which is its largest $1$-reduced subcomplex containing the basepoint.
+Note that, in the absence of a basepoint, there is generally no canonical choice of such a subcomplex, so the use of pointed simplicial sets is essential for this construction.
+In fact, we shall see that the forgetful functor from $1$-reduced simplicial sets to simplicial sets has no right adjoint.
+We present the definition of the 11st Eilenberg subcomplex, though the only feature that we need in this paper is the adjunction that it features in.
+Recall from Remark 2.6 that, for a simplicial set $X$, we have a canonical map $X\rightarrow\operatorname{cosk}_{1}(X)$ to its $1$-coskeleton.
+Definition 2.17 ([30, Definition 8.3]).
+The 11st Eilenberg subcomplex $E_{1}(X)$ of a pointed simplicial set $X$ with basepoint $x_{0}$ is defined by the pullback
+[…]
+In detail, $E_{1}(X)$ consists of all simplices $\chi$ of $X$ such that every 0-dimensional face of $\chi$ is $x_{0}$ and every $1$-dimensional face of $\chi$ is the degeneracy of $x_{0}$.
+It follows that the 11st Eilenberg subcomplex $E_{1}(X)$ of a pointed simplicial set $X$ is $1$-reduced.
+The 11st Eilenberg subcomplex of a pointed simplicial set is its largest $1$-reduced subcomplex in the sense of the following universal property that is straightforward to verify.
+Proposition 2.18.
+Every pointed map $f\colon X\rightarrow Y$ from a $1$-reduced simplicial set $X$ to a pointed simplicial set $Y$ maps into the 11st Eilenberg subcomplex $E_{1}(Y)$ of $Y$.
 Thus, there is a unique map $\hat{f}\colon X\rightarrow E_{1}(Y)$ such that the diagram
-${X}$${{E_{1}(Y)}}$${{Y}}$$\scriptstyle{\hat{f}}$$\scriptstyle{f}$
+[…]
 commutes, where $E_{1}(Y)\rightarrow Y$ is the inclusion map.
-This universal property amounts to the promised adjunction, which is stated in Part (2) of the following lemma.
-Lemma 2.23.
+This universal property amounts to the promised adjunction, which is in part (2) of the following lemma.
+Proposition 2.19.
 Let $F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$ denote the forgetful functor.
@@
   2. (2)
-    The $1$st Eilenberg subcomplex functor $E_{1}\colon{\mathcal{SS}\mathrm{ets}_{\ast}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$ is right adjoint to $F$.
-Lemma 2.23 implies that small limits and colimits in ${\mathcal{SS}\mathrm{ets}}_{1}$ are constructed as in ${\mathcal{SS}\mathrm{ets}}_{\ast}$. Given a diagram
-$L\colon\mathcal{D}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$
-in $1$-reduced simplicial sets, we form its (co)limit by post-composing $L$ with the forgetful functor
-$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$
-and forming the (co)limit of the diagram
-$FL\colon\mathcal{D}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$
-in pointed simplicial sets. This (co)limit is a $1$-reduced simplicial set, and all the structure maps are in ${\mathcal{SS}\mathrm{ets}}_{1}$.
-In particular, the initial object of ${\mathcal{SS}\mathrm{ets}}_{1}$ is the point $\Delta[0]$, not the empty simplicial set $\emptyset$.
-Thus, the forgetful functor $U\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}$ does not preserve the initial object, so it cannot admit a right adjoint.
+    The 11st Eilenberg subcomplex functor $E_{1}\colon{\mathcal{SS}\mathrm{ets}_{\ast}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$ is right adjoint to $F$.
+Proposition 2.19 implies that small limits and colimits in ${\mathcal{SS}\mathrm{ets}}_{1}$ are constructed as in $\mathcal{SS}\mathrm{ets}$$_{\ast}$.
+Given a diagram $L\colon\mathcal{D}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$ in $1$-reduced simplicial sets, we form its (co)limit by post-composing $L$ with the forgetful functor $F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$ and forming the (co)limit of the diagram $FL\colon\mathcal{D}\rightarrow{\mathcal{SS}\mathrm{ets}}_{\ast}$ in pointed simplicial sets.
+In particular, coproducts in ${\mathcal{SS}\mathrm{ets}}_{1}$ are given by wedge sums instead of disjoint unions, for the latter do not yield a $1$-reduced simplicial set.
+More fundamentally, the initial object of ${\mathcal{SS}\mathrm{ets}}_{1}$ is the point $\Delta[0]$; the empty set $\emptyset$ is not $1$-reduced.
+Thus, the forgetful functor $U\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightarrow{\mathcal{SS}\mathrm{ets}}$ does not preserve the initial object or coproducts, so it cannot admit a right adjoint.
 3. The left transfer to $1$-reduced simplicial sets
-We transfer the model structure on the category ${\mathcal{SS}\mathrm{ets}}_{\ast}$ of pointed simplicial sets along the adjunction
+We recover the first model structure on $1$-reduced simplicial sets from Quillen’s family, which we then localize to recover the rest of the family.
+Namely, we left-transfer the model structure on the category $\mathcal{SS}\mathrm{ets}$$_{\ast}$ of pointed simplicial sets along the adjunction
 $F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}$
 to a model structure on the category ${\mathcal{SS}\mathrm{ets}}_{1}$ of $1$-reduced simplicial sets whose weak equivalences are the weak homotopy equivalences.
-This left transfer is an application of [BHK+15, Theorem 2.23].
+(Our notation for adjunctions is amenable to a helpful mnemonic that distinguishes a left transfer from a right transfer; we propagate the existence of a model structure from right to left, so it is a left transfer.) This left transfer is an application of [4, Theorem 2.23], which we recall after going over the background on model category theory that we rely on in this paper.
 3.1. Model categories
-We collect some background on model categories from [DS95, §3] and [Hir03, Chapters 7, 11, and 13]. We first recall a concept in the homotopy theory of simplicial sets from [GJ99, Chapter I].
+We collect some background on model categories from [18, §3] and [26, Chapters 7, 11, and 13]. We first recall a concept in the homotopy theory of simplicial sets from [23, Chapter I].
 Definition 3.1.
-A Kan fibration is a simplicial map $p\colon X\rightarrow Y$ that has the right lifting property with respect to the set of all horn inclusions
-$J=\left\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{ and}0\leq k\leq n\right\}.$
+A Kan fibration is a simplicial map $p\colon X\rightarrow Y$ that has the right lifting property with respect to the set of all horn inclusions $J=\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{ and }0\leq k\leq n\}$.
 In other words, a simplicial map $p\colon X\rightarrow Y$ is a Kan fibration if every lifting problem of the form
-${{\Lambda[n,k]}}$${X}$${{\Delta[n]}}$${Y}$$\scriptstyle{j_{n,k}}$$\scriptstyle{p}$
+[…]
 can be solved. A Kan complex is a simplicial set $X$ such that the unique map $X\rightarrow\Delta[0]$ is a Kan fibration.
-Example 3.2 ([GJ99, Chapter I, Lemma 3.3]).
+Example 3.2 ([23, Chapter I, Lemma 3.3]).
 If $Y$ is a space, then $\operatorname{Sing}(Y)$ is a Kan complex.
-The homotopy theory of simplicial sets can be studied systematically in the framework of model categories.
+Example 3.2 follows from the geometric realization-singular complex adjunction combined with the fact that, at the level of topological spaces, geometric realizations of horns can always be filled because they are retracts of geometric realizations of standard simplices.
+However, the directionality of simplices in a simplicial set precludes the existence of the analogous retraction at the level of simplicial sets, so the concepts of a Kan fibration and a Kan complex are not moot.
+The homotopy theory of simplicial sets can be studied systematically in the framework of model categories, which is the framework that we adhere to throughout this paper.
 Definition 3.3.
@@
 A weak equivalence that is also a fibration is called an acyclic fibration.
-This data satisfies the axioms in [DS95, Definition 3.3].
+This data satisfies the axioms in [18, Definition 3.3].
 Remark 3.4.
-In a model category, the weak equivalences and the fibrations together determine the cofibrations. Dually, the weak equivalences and the cofibrations together determine the fibrations.
-Remark 3.5.
-Let $\mathcal{M}$ be a model category with initial object $\emptyset$ and terminal object $\ast$. An object $X$ of $\mathcal{M}$ is
-  1. (1)
-    cofibrant if the unique morphism $\emptyset\rightarrow X$ is a cofibration; and
-  2. (2)
-    fibrant if the unique morphism $X\rightarrow\ast$ is a fibration.
-Then, for an object $Y$ of $\mathcal{M}$,
-  1. (1)
-    a cofibrant replacement of $Y$ is a weak equivalence $A\rightarrow Y$ with cofibrant source $A$; and
-  2. (2)
-    a fibrant replacement of $Y$ is a weak equivalence $Y\rightarrow B$ with fibrant target $B$.
-Example 3.6 ([Hir03, Theorems 7.10.12 and 7.10.13]).
+In a model category, the weak equivalences and the cofibrations together determine the fibrations.
+To recover Quillen’s family of model structures on $1$-reduced simplicial sets, we rely on the standard model structure on (pointed) simplicial sets.
+Example 3.5 ([26, Theorems 7.10.12 and 7.10.13]).
 The category $\mathcal{SS}\mathrm{ets}$} of simplicial sets has a model structure in which a map $f$ is
   1. (1)
-    a weak equivalence if and only if $f$ is a weak homotopy equivalence;
+    a weak equivalence if and only if $|f|$ is a weak homotopy equivalence;
   2. (2)
@@
 The category $\mathcal{SS}\mathrm{ets}$$_{\ast}$ of pointed simplicial sets also has a model structure with the same description.
-The model categories ${\mathcal{SS}\mathrm{ets}}$ and ${\mathcal{SS}\mathrm{ets}}_{\ast}$ have the additional property of being cofibrantly generated.
-In practice, the property of being cofibrantly generated facilitates the detection of (acyclic) fibrations in a model category.
-Definition 3.7 ([Hir03, Definition 11.1.2]).
+The following definition is intended to introduce some terminology for the rest of this paper.
+Definition 3.6.
+Let $\mathcal{M}$ be a model category with initial object $\emptyset$ and terminal object $\ast$.
+An object $X$ of $\mathcal{M}$ is cofibrant if the morphism $\emptyset\rightarrow X$ is a cofibration, and $X$ is fibrant if the morphism $X\rightarrow\ast$ is a fibration.
+Then, for an object $Y$ of $\mathcal{M}$, a cofibrant replacement of $Y$ is a weak equivalence $A\rightarrow Y$ with cofibrant source $A$, and a fibrant replacement of $Y$ is a weak equivalence $Y\rightarrow B$ with fibrant target $B$.
+The model categories $\mathcal{SS}\mathrm{ets}$} and $\mathcal{SS}\mathrm{ets}$$_{\ast}$ have the property of being cofibrantly generated.
+In practice, this property facilitates the detection of (acyclic) fibrations and fibrant objects in a model category.
+Definition 3.7 ([26, Definition 11.1.2]).
 A model category $\mathcal{M}$ is cofibrantly generated if there exist
@@
   1. (1)
-    both $I$ and $J$ permit the small object argument [Hir03, Definition 10.5.15];
+    both $I$ and $J$ permit the small object argument [26, Definition 10.5.15];
   2. (2)
@@
     a morphism $p$ is a fibration if and only if $p$ has the right lifting property with respect to $J$.
-Example 3.8 ([Hir03, Examples 11.1.6 and 11.1.7]).
+Example 3.8 ([26, Examples 11.1.6 and 11.1.7]).
 Both $\mathcal{SS}\mathrm{ets}$} and $\mathcal{SS}\mathrm{ets}$$_{\ast}$ are cofibrantly generated.
@@
     A generating set of cofibrations for $\mathcal{SS}\mathrm{ets}$} is the set of all boundary inclusions
-    $I=\left\{i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]\mid n\geq 0\right\}.$
-    Attaching a disjoint basepoint yields a generating set of cofibrations $I_{+}$ for ${\mathcal{SS}\mathrm{ets}}_{\ast}$.
+    $I=\{i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]\mid n\geq 0\}.$
   2. (2)
     A generating set of acyclic cofibrations for $\mathcal{SS}\mathrm{ets}$} is the set of all horn inclusions
-    $J=\left\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{ and}0\leq k\leq n\right\}.$
-    Attaching a disjoint basepoint yields a generating set of acyclic cofibrations $J_{+}$ for ${\mathcal{SS}\mathrm{ets}}_{\ast}$.
+    $J=\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{ and }0\leq k\leq n\}.$
+Attaching a disjoint basepoint to the maps in both sets yields the generating (acyclic) cofibrations of $\mathcal{SS}\mathrm{ets}$$_{\ast}$.
 What is more, $\mathcal{SS}\mathrm{ets}$} and $\mathcal{SS}\mathrm{ets}$$_{\ast}$ belong in the following class of well-behaved model categories, which is originally due to J.H.
 Smith and is particularly expedient for our purposes of left Bousfield localization.
-Definition 3.9 ([Dug01, Definition 2.1]).
-A model category $\mathcal{M}$ is combinatorial if $\mathcal{M}$ is cofibrantly generated and its underlying category is locally presentable [AR94, Definition 1.17].
-Example 3.10 ([AR94, Example 1.12]).
+Definition 3.9 ([17, Definition 2.1]).
+A model category $\mathcal{M}$ is combinatorial if $\mathcal{M}$ is cofibrantly generated and its underlying category is locally presentable [1, Definition 1.17].
+In a combinatorial model category, not only do we have a set of building blocks for its (acyclic) cofibrations by virtue of cofibrant generation, but we also have a set of building blocks for its objects by virtue of local presentability.
+Though the category of topological spaces is not locally presentable [1, Example 1.18(5)], the category of simplicial sets does not run into this problem as every simplicial set is built out of standard simplices.
+Example 3.10 ([1, Example 1.12]).
 Both $\mathcal{SS}\mathrm{ets}$} and $\mathcal{SS}\mathrm{ets}$$_{\ast}$ are locally presentable, thus combinatorial.
-Lastly, ${\mathcal{SS}\mathrm{ets}}$ and ${\mathcal{SS}\mathrm{ets}}_{\ast}$ satisfy the following model-categorical property that permits left Bousfield localization.
-Definition 3.11 ([Hir03, Definition 13.1.1(1)]).
+Also, $\mathcal{SS}\mathrm{ets}$} and $\mathcal{SS}\mathrm{ets}$$_{\ast}$ have the following model-categorical property that permits left Bousfield localization.
+Definition 3.11 ([26, Definition 13.1.1(1)]).
 A model category $\mathcal{M}$ is left proper if every pushout of a weak equivalence along a cofibration is a weak equivalence.
-The following result guarantees the left properness of almost every model category that appears in this paper.
-Proposition 3.12 ([Hir03, Corollary 13.1.3(1)]).
+The following result guarantees the left properness of almost every model category that appears in this paper, including every model structure in Quillen’s family of model structures on $1$-reduced simplicial sets.
+Proposition 3.12 ([26, Corollary 13.1.3(1)]).
 If every object of a model category $\mathcal{M}$ is cofibrant, then $\mathcal{M}$ is left proper.
-3.2. The left transfer to $1$-reduced simplicial sets
-Firstly, to appreciate the necessity of using pointed simplicial sets and their $1$st Eilenberg subcomplex, we explain why we cannot use the $1$-reduction functor $R_{1}$ to right-induce, in the sense of [Hir03, Theorem 11.3.2], a model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$ along the adjunction
+In practice, it is common and convenient for the cofibrations in a model category to be the monomorphisms, as is the theme in this paper.
+Then, every object is cofibrant, so we need not worry about cofibrant replacement, and we can use Proposition 3.12 to extract left properness for free.
+3.2. The left transfer
+Before we do the promised left transfer, we explain why we cannot use the $1$-reduction functor $R_{1}$ to right-induce, in the sense of [26, Theorem 11.3.2], a model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$ along the adjunction
 ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$
-from Lemma 2.19.
+from Proposition 2.15.
 We write $J$ for the set of all horn inclusions in $\mathcal{SS}\mathrm{ets}$}.
@@
     the set $R_{1}(J)$ would be a generating set of acyclic cofibrations.
-Thus, every map in $R_{1}(J)$ would be a weak homotopy equivalence. We verify that this necessary condition fails.
+In particular, the set $|R_{1}(J)|$ would comprise weak homotopy equivalences, which is false by part (2) of the following lemma; conceptually, we find an instance of the $1$-reduction quotient being homotopically ill-behaved.
 Lemma 3.13.
@@
   1. (1)
-    If $n=1$, then ${R}_{1}\left(j_{1,k}\right)$ is the identity map of the point.
+    If $n=1$, then ${R}_{1}(j_{1,k})$ is the identity map of the point.
   2. (2)
-    If $n=2$, then ${R}_{1}\left(j_{2,k}\right)\colon\Delta[0]\rightarrow S_{\mathrm{simp}}^{2}$ picks out the unique 00-simplex of the simplicial $2$-sphere.
+    If $n=2$, then ${R}_{1}(j_{2,k})\colon\Delta[0]\rightarrow S_{\mathrm{simp}}^{2}$ picks out the unique 0-simplex of the simplicial $2$-sphere.
   3. (3)
-    If $n\geq 3$, then $|{R}_{1}\left(j_{n,k}\right)|$ is the inclusion of a strong deformation retract in topological spaces.
-Therefore, ${R}_{1}\left(j_{n,k}\right)$ is a weak homotopy equivalence if and only if $n\neq 2$.
+    If $n\geq 3$, then $|{R}_{1}(j_{n,k})|$ is the inclusion of a strong deformation retract in topological spaces.
+Thus, $|{R}_{1}(j_{n,k})|$ is a weak homotopy equivalence if and only if $n\neq 2$.
 Proof.
 Parts (1) and (2) are straightforward.
-For (3), if $n\geq 3$, then $\operatorname{sk}_{1}\left(\Lambda[n,k]\right)=\operatorname{sk}_{1}\left(\Delta[n]\right)$, so the inclusion of a strong deformation retract $|j_{n,k}|$ descends to the inclusion of a strong deformation retract $|{R}_{1}\left(j_{n,k}\right)|$.
+For (3), if $n\geq 3$, then $\operatorname{sk}_{1}(\Lambda[n,k])=\operatorname{sk}_{1}(\Delta[n])$, so the inclusion of a strong deformation retract $|j_{n,k}|$ descends to the inclusion of a strong deformation retract $|{R}_{1}(j_{n,k})|$.
 ∎
@@
 If a simplicial map $f\colon X\rightarrow Y$ is a monomorphism, then so is its $1$-reduction ${R}_{1}(f)$.
-We left-induce a model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$ by transferring the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ along the adjunction
+Instead, we left-induce a model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$ by transferring that on $\mathcal{SS}\mathrm{ets}$$_{\ast}$ along the adjunction
 $F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}$
-from Lemma 2.23(2). To do so, we apply the following general result on left-induced model structures.
-Theorem 3.15 ([BHK+15, Theorem 2.23]).
+from Proposition 2.19(2). To do so, we apply the following general result on left-induced model structures.
+Theorem 3.15 ([4, Theorem 2.23]).
 Let $F\colon\mathcal{C}\rightleftarrows\mathcal{M}\colon G$ be an adjunction, where $\mathcal{C}$ is a locally presentable category and $\mathcal{M}$ is a combinatorial model category with class of cofibrations $Q$.
-Suppose that the following condition holds: if a morphism $g$ in $\mathcal{C}$ has the right lifting property with respect to the class $F^{-1}(Q)$, then $F(g)$ is a weak equivalence in $\mathcal{M}$.
-Then, there is a combinatorial model structure on $\mathcal{C}$ in which a morphism $f$ is
+Suppose that, if a map $g$ in $\mathcal{C}$ has the right lifting property with respect to the class $F^{-1}(Q)$, then $F(g)$ is a weak equivalence in $\mathcal{M}$.
+Then, there is a combinatorial model structure on $\mathcal{C}$, called the left-induced model structure, in which a map $f$ is
   1. (1)
@@
     a cofibration if and only if $F(f)$ is a cofibration in $\mathcal{M}$.
-As ${\mathcal{SS}\mathrm{ets}}_{1}$ is a locally presentable category [AR94, Example 1.12] and ${\mathcal{SS}\mathrm{ets}}_{\ast}$ is a combinatorial model category by Example 3.10, it remains to verify the condition in Theorem 3.15.
+As ${\mathcal{SS}\mathrm{ets}}_{1}$ is a locally presentable category [1, Example 1.12] and $\mathcal{SS}\mathrm{ets}$$_{\ast}$ is a combinatorial model category by Example 3.10, it remains to verify that our adjunction satisfies the homotopical condition in Theorem 3.15.
 Lemma 3.16.
 Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets.
-If $f$ has the right lifting property with respect to every monomorphism of $1$-reduced simplicial sets, then $f$ is a weak homotopy equivalence.
+If $f$ has the right lifting property with respect to every monomorphism of $1$-reduced simplicial sets, then $|f|$ is a weak homotopy equivalence.
 Proof.
@@
 Therefore, the hypothesis informs us that, for every $n\geq 0$, $f$ has the right lifting property with respect to ${R}_{1}(i_{n})$.
-Then, the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19 yields that, for every $n\geq 0$, $f$ has the right lifting property with respect to the boundary inclusion $i_{n}$.
+Then, the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Proposition 2.15 yields that, for every $n\geq 0$, $f$ has the right lifting property with respect to the boundary inclusion $i_{n}$.
 We conclude that $f$ is an acyclic fibration in $\mathcal{SS}\mathrm{ets}$}.
 ∎
-We begin our recovery of the family of model structures on $1$-reduced simplicial sets in [Qui69, Part II, Theorem 2.2] and the characterization of fibrant objects in [Qui69, Part II, Corollary 2.6].
+We begin recovering the family of model structures on $1$-reduced simplicial sets in [33, Part II, Theorem 2.2].
 Theorem 3.17.
@@
   1. (1)
-    a weak equivalence if and only if $f$ is a weak homotopy equivalence; and
+    a weak equivalence if and only if $|f|$ is a weak homotopy equivalence; and
   2. (2)
@@
 For $n\geq 1$ and $0\leq k\leq n$, we solve a lifting problem
-${{\Lambda[n,k]}}$${X}$${{\Delta[n].}}$$\scriptstyle{j_{n,k}}$
-Using the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19, we may equivalently solve a lifting problem
-${{{R}_{1}\left(\Lambda[n,k]\right)}}$${X}$${{{R}_{1}\left(\Delta[n]\right).}}$$\scriptstyle{{R}_{1}\left(j_{n,k}\right)}$
+[…]
+Using the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Proposition 2.15, we equivalently solve a lifting problem
+[…]
 In light of Lemma 3.13, we split our solution in two separate cases.
-If $n\neq 2$, then, by Lemmas 3.13 and 3.14, ${R}_{1}\left(j_{n,k}\right)$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, so our lifting problem has a solution since $X$ is fibrant in ${\mathcal{SS}\mathrm{ets}}_{1}$.
+If $n\neq 2$, then, by Lemmas 3.13 and 3.14, ${R}_{1}(j_{n,k})$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, so our lifting problem has a solution since $X$ is fibrant in ${\mathcal{SS}\mathrm{ets}}_{1}$.
 If $n=2$, then we solve a lifting problem
 […]
-where $x_{0}$ denotes the unique 00-simplex of $X$, by setting $\psi$ to be the constant map at $x_{0}$. ∎
-The left transfer in Theorem 3.17 renders the adjunction
-$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}$
-from Lemma 2.23(2) a Quillen adjunction; in particular, we note that $E_{1}$ preserves fibrant objects.
-Corollary 3.18.
-If $X$ is a pointed Kan complex, then $E_{1}(X)$ is a Kan complex.
-In particular, we occasionally appeal to a special instance of Corollary 3.18.
+where $x_{0}$ denotes the unique 0-simplex of $X$, by setting $\psi$ to be the constant map at $x_{0}$. ∎
+The left transfer renders $F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}$ a Quillen adjunction [26, Definition 8.5.2], so $E_{1}$ preserves fibrant objects; an alternative proof of this preservation uses the pullback definition of the 11st Eilenberg subcomplex together with the fact that, if $X$ is a Kan complex, then $X\rightarrow\operatorname{cosk}_{1}(X)$ is a Kan fibration.
+We record the following consequence for the next section.
 Corollary 3.18.
 If $Y$ is a pointed topological space, then $E_{1}\operatorname{Sing}(Y)$ is a Kan complex.
-Proof.
-By Example 3.2, $\operatorname{Sing}(Y)$ is a pointed Kan complex, so the claim follows from Corollary 3.18. ∎
 4. Fibrations of $1$-reduced simplicial sets
@@
 Moreover, we characterize when a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ is a Kan fibration, and we find a counterintuitive fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
+The existence of such counterintuitive fibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ serves not only as an impediment toward finding a generating set of acyclic cofibrations, but also as evidence that the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ should not be thought of as a restriction of the model structure on pointed simplicial sets.
+Instead, ${\mathcal{SS}\mathrm{ets}}_{1}$ encodes its own homotopy theory that has its individual intricacies, some of which we shed light on in this section.
 4.1. Fibrant replacement of $1$-reduced simplicial sets
@@
 Let $X$ be a $1$-reduced simplicial set.
-The geometric realization-singular complex adjunction from Proposition 2.7 provides the natural map
-$X\rightarrow\operatorname{Sing}(|X|),$
-which maps into the $1$st Eilenberg subcomplex $E_{1}\operatorname{Sing}(|X|)$. Thus, we obtain the commutative diagram
-${X}$${{E_{1}\operatorname{Sing}(|X|)}}$${{\operatorname{Sing}(|X|),}}$
+The geometric realization-singular complex adjunction provides the map $X\rightarrow\operatorname{Sing}(|X|)$, which lands into $E_{1}\operatorname{Sing}(|X|)$.
+We get the commutative diagram
+[…]
 where $E_{1}\operatorname{Sing}(|X|)\rightarrow\operatorname{Sing}(|X|)$ is the inclusion map. We show that the natural map
 $X\rightarrow E_{1}\operatorname{Sing}(|X|)$
-is a weak homotopy equivalence, thus a functorial fibrant replacement of $X$ in ${\mathcal{SS}\mathrm{ets}}_{1}$.
-To do so, it suffices to show that the other two maps in the commutative diagram are weak homotopy equivalences.
-Firstly, Part (1) of the following theorem informs us that the natural map
-$X\rightarrow\operatorname{Sing}(|X|)$
-is a weak homotopy equivalence; Part (2) is for later use in this argument.
-Theorem 4.1 ([GJ99, Chapter I, Theorem 11.4]).
+is a weak equivalence, thus a functorial fibrant replacement of $X$ in ${\mathcal{SS}\mathrm{ets}}_{1}$.
+To do so, it suffices to show that the other two maps in the commutative diagram are weak equivalences.
+Firstly, part (1) of the following theorem says that $X\rightarrow\operatorname{Sing}(|X|)$ is a weak equivalence; part (2) is for later use in our argument.
+Theorem 4.1 ([23, Chapter I, Theorem 11.4]).
   1. (1)
-    For every simplicial set $X$, the natural map
-    $X\rightarrow\operatorname{Sing}(|X|)$
-    is a weak homotopy equivalence.
+    For every simplicial set $X$, the map $X\rightarrow\operatorname{Sing}(|X|)$ is a weak equivalence.
   2. (2)
-    For every topological space $Y$, the natural map
-    $|\operatorname{Sing}(Y)|\rightarrow Y$
-    is a weak homotopy equivalence.
+    For every topological space $Y$, the map $|\operatorname{Sing}(Y)|\rightarrow Y$ is a weak homotopy equivalence.
 Remark 4.2.
-Part (1) informs us that, for a simplicial set $X$, the natural map
-$X\rightarrow\operatorname{Sing}(|X|)$
-is a functorial fibrant replacement of $X$ in ${\mathcal{SS}\mathrm{ets}}$, as $\operatorname{Sing}(|X|)$ is a Kan complex by Example 3.2.
-As for the $1$st Eilenberg subcomplex inclusion
-$E_{1}\operatorname{Sing}(|X|)\rightarrow\operatorname{Sing}(|X|),$
-it is also a weak homotopy equivalence by a special case of the following theorem.
-Theorem 4.3 ([May67, Theorem 8.4(iii)]).
-If $X$ is a pointed Kan complex, then the $1$st Eilenberg subcomplex inclusion $E_{1}(X)\rightarrow X$ induces, for every $n\geq 2$, the isomorphism
-$\pi_{n}\left(E_{1}(X)\right)\xrightarrow{\cong}\pi_{n}(X).$
-In particular, if $X$ is a pointed and $1$-connected Kan complex, then the inclusion
-$E_{1}(X)\rightarrow X$
-is a weak homotopy equivalence.
-Remark 4.4.
-In other words, if $X$ is a pointed Kan complex, then the $1$st Eilenberg subcomplex inclusion
-$E_{1}(X)\rightarrow X$
-is a $1$-connected cover of $X$.
-As an aside, using the construction of the homotopy groups of Kan complexes in [GJ99, Chapter I.7] together with Lemma 2.22, we can provide an alternative proof of Theorem 4.3 from that in [May67, Theorem 8.4(iii)].
+By Example 3.2, part (1) defines a functorial fibrant replacement in $\mathcal{SS}\mathrm{ets}$.
+The inclusion $E_{1}\operatorname{Sing}(|X|)\rightarrow\operatorname{Sing}(|X|)$ is also a weak equivalence by the following theorem.
+Theorem 4.3 ([30, Theorem 8.4(iii)]).
+If $X$ is a pointed Kan complex, then the 11st Eilenberg subcomplex inclusion $E_{1}(X)\rightarrow X$ induces, for every $n\geq 2$, an isomorphism $\pi_{n}(E_{1}(X))\xrightarrow{\cong}\pi_{n}(X)$.
+In particular, if $X$ is a pointed and $1$-connected Kan complex, then the inclusion $E_{1}(X)\rightarrow X$ is a weak equivalence.
+In other words, if $X$ is a pointed Kan complex, then the inclusion $E_{1}(X)\rightarrow X$ is a $1$-connected cover of $X$.
+As an aside, using the construction of the homotopy groups of Kan complexes from [23, Chapter I.7] together with Proposition 2.18, we give an alternative proof of Theorem 4.3 to that in [30, Theorem 8.4(iii)].
 Proof.
-Let $X$ be a pointed Kan complex.
-Let $n\geq 2$.
-We first show that $\pi_{n}\left(E_{1}(X)\right)\rightarrow\pi_{n}(X)$ is an epimorphism.
-As $X$ is a Kan complex, every element of $\pi_{n}(X)$ is of the form $[\omega]$ for a map
-$\omega\colon S_{\mathrm{simp}}^{n}\rightarrow X.$
-As $n\geq 2$, the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ is $1$-reduced, so the universal property of the $1$st Eilenberg subcomplex from Lemma 2.22 yields the commutative diagram
+We first show that $\pi_{n}(E_{1}(X))\rightarrow\pi_{n}(X)$ is an epimorphism.
+As $X$ is a Kan complex, every element of $\pi_{n}(X)$ is of the form $\omega]$ for a map $\omega\colon S_{\mathrm{simp}}^{n}\rightarrow X$, and the universal property of the 11st Eilenberg subcomplex yields the commutative diagram
 […]
-It remains to show that $\pi_{n}\left(E_{1}(X)\right)\rightarrow\pi_{n}(X)$ is a monomorphism. Given a pointed homotopy
-$H\colon S_{\mathrm{simp}}^{n}\wedge\Delta[1]_{+}\rightarrow X,$
-we observe that $S_{\mathrm{simp}}^{n}\wedge\Delta[1]_{+}$ is $1$-reduced by Lemma 2.15 because $S_{\mathrm{simp}}^{n}$ is $1$-reduced, so the universal property of the $1$st Eilenberg subcomplex from Lemma 2.22 yields the commutative diagram
+It remains to show that $\pi_{n}(E_{1}(X))\rightarrow\pi_{n}(X)$ is a monomorphism.
+Given a homotopy $H\colon S_{\mathrm{simp}}^{n}\wedge\Delta[1]_{+}\rightarrow X$, we observe that $S_{\mathrm{simp}}^{n}\wedge\Delta[1]_{+}$ is $1$-reduced by Lemma 2.11, so we obtain the commutative diagram
 […]
+We infer that the inclusion $E_{1}\operatorname{Sing}(|X|)\rightarrow\operatorname{Sing}(|X|)$ is also a weak equivalence.
 Corollary 4.4.
-If $Y$ is a pointed and $1$-connected topological space, then the inclusion
-$E_{1}\operatorname{Sing}(Y)\rightarrow\operatorname{Sing}(Y)$
-is a weak homotopy equivalence.
+If $Y$ is a pointed and $1$-connected space, then $E_{1}\operatorname{Sing}(Y)\rightarrow\operatorname{Sing}(Y)$ is a weak equivalence.
 Proof.
-By Example 3.2, $\operatorname{Sing}(Y)$ is a Kan complex. By Theorem 4.1(2), $\operatorname{Sing}(Y)$ is $1$-connected. ∎
+By Example 3.2, $\operatorname{Sing}(Y)$ is a Kan complex. By Theorem 4.1(2), $\operatorname{Sing}(Y)$ is 11-connected. ∎
 In conclusion, we obtain functorial fibrant replacements in ${\mathcal{SS}\mathrm{ets}}_{1}$.
 Proposition 4.5.
-Let $X$ be a $1$-reduced simplicial set. Then, the natural map
-$X\rightarrow E_{1}\operatorname{Sing}(|X|)$
-is a functorial fibrant replacement of $X$ in ${\mathcal{SS}\mathrm{ets}}_{1}$.
-4.2. Generating sets of cofibrations of $1$-reduced simplicial sets
+Let $X$ be a $1$-reduced simplicial set. Then, the natural map $X\rightarrow E_{1}\operatorname{Sing}(|X|)$ is a functorial fibrant replacement of $X$ in ${\mathcal{SS}\mathrm{ets}}_{1}$.
+4.2. Generating cofibrations of $1$-reduced simplicial sets
 Our left transfer from Theorem 3.17 provides us with a combinatorial model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$, but not with explicit generating sets of cofibrations and acyclic cofibrations.
-As a first remedy, we produce an explicit generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$.
+As a first remedy, we extract an explicit generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$ from the set of boundary inclusions for $\mathcal{SS}\mathrm{ets}$} using the $1$-reduction functor $R_{1}$.
 Proposition 4.6.
-Let $I=\left\{i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]\mid n\geq 0\right\}$ denote the generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}$.
+Let $I=\{i_{n}\colon\partial\Delta[n]\rightarrow\Delta[n]\mid n\geq 0\}$ denote the generating set of cofibrations for $\mathcal{SS}\mathrm{ets}$}.
 Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets.
@@
     The map $f\colon X\rightarrow Y$ has the right lifting property with respect to the set
-    ${R}_{1}(I)=\left\{{R}_{1}(i_{n})\colon{R}_{1}(\partial\Delta[n])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 0\right\}.$
+    ${R}_{1}(I)=\{{R}_{1}(i_{n})\colon{R}_{1}(\partial\Delta[n])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 0\}.$
 Hence, ${R}_{1}(I)$ is a generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$.
@@
 In particular, $f$ has the right lifting property with respect to all monomorphisms in ${\mathcal{SS}\mathrm{ets}}_{1}$, so $f$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
-Then, (2) implies (3) as, by Lemma 3.14, ${R}_{1}(I)$ is a set of cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$.
+Also, (2) implies (3) as, by Lemma 3.14, ${R}_{1}(I)$ is a set of cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$.
 Lastly, we show that (3) implies (1).
@@
 To prove that $f$ is an acyclic fibration in $\mathcal{SS}\mathrm{ets}$}, we solve, for $n\geq 0$, a lifting problem
-${{\partial\Delta[n]}}$${X}$${{\Delta[n]}}$${{Y.}}$$\scriptstyle{i_{n}}$$\scriptstyle{f}$
-Under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19, we equivalently solve a lifting problem
-${{{R}_{1}(\partial\Delta[n])}}$${X}$${{{R}_{1}(\Delta[n])}}$${{Y.}}$$\scriptstyle{{R}_{1}(i_{n})}$$\scriptstyle{f}$
+[…]
+Under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Proposition 2.15, we equivalently solve a lifting problem
+[…]
 As ${R}_{1}(i_{n})$ is in ${R}_{1}(I)$ and $f$ has the right lifting property with respect to ${R}_{1}(I)$, a solution exists. ∎
-Even though our treatment of Proposition 4.7 is biased towards the equivalence of (2) and (3), which grants us our generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$, we would be remiss if we neglected the equivalence of (1) and (2).
-The fact that (1) and (2) are equivalent means that ${\mathcal{SS}\mathrm{ets}}_{1}$ has no pathological acyclic fibrations.
-As we shall see later in this section, Proposition 4.7 finds no analog in the study of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$.
-The reason is that ${\mathcal{SS}\mathrm{ets}}_{1}$ has fibrations that are dissimilar to Kan fibrations, a phenomenon that we also investigate in this section.
-Thus, the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ is more than a restriction of the model category ${\mathcal{SS}\mathrm{ets}}$ to $1$-reduced simplicial sets; after all, the model category axioms are generally not amenable to such a restriction.
-Instead, the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ formalizes a different homotopy theory that is interesting in its own right.
+Even though our treatment of Proposition 4.6 is biased toward the equivalence of statements (2) and (3), which grants us our generating set of cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$, we would be remiss if we neglected the equivalence of statements (1) and (2).
+The fact that statements (1) and (2) are equivalent means that ${\mathcal{SS}\mathrm{ets}}_{1}$ has no pathological acyclic fibrations, which follows from the fact that ${\mathcal{SS}\mathrm{ets}}_{1}$ has no pathological cofibrations.
+As we shall see later in this section, Proposition 4.6 finds no analog in the study of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ because ${\mathcal{SS}\mathrm{ets}}_{1}$ has fibrations that are dissimilar to Kan fibrations.
 4.3. Acyclic cofibrations of $1$-reduced simplicial sets
-Our attempt to emulate Proposition 4.7 towards a generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$ encounters a problem: the $1$-reduction of a boundary inclusion is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, but the $1$-reduction of a horn inclusion need not be an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
+Our attempt to emulate Proposition 4.6 toward a generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$ encounters a problem: the $1$-reduction of a boundary inclusion is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, but the $1$-reduction of a horn inclusion need not be an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
 Lemma 4.7.
 For $n\geq 1$ and $0\leq k\leq n$, let $j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]$ denote the horn inclusion in $\mathcal{SS}\mathrm{ets}$}.
-Then, ${R}_{1}\left(j_{n,k}\right)$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $n\neq 2$.
+Then, ${R}_{1}(j_{n,k})$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $n\neq 2$.
 Proof.
 The claim follows from Lemmas 3.13 and 3.14 and the definition of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$. ∎
-Nevertheless, we can partially work our way around this obstacle. Let
-$J=\left\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{ and}0\leq k\leq n\right\}$
-denote the generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}$. Write
-${R}_{1}(J)=\left\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 1\textrm{ and }0\leq k\leq n\right\},$
-which is a set of cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ by Lemma 3.14, but not a set of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ by Lemma 4.8.
+We can partially work our way around this obstacle.
+Let $J=\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{ and }0\leq k\leq n\}$ denote the generating set of acyclic cofibrations for $\mathcal{SS}\mathrm{ets}$}.
+Write
+${R}_{1}(J)=\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 1\textrm{ and }0\leq k\leq n\}.$
 To extract a candidate generating set of acyclic cofibrations for $\mathcal{SS}\mathrm{ets}$}, we use Lemma 3.13 to strike out
@@
   2. (2)
-    the maps ${R}_{1}(j_{2,0})$, ${R}_{1}(j_{2,1})$, and ${R}_{1}(j_{2,2})$ which are not weak homotopy equivalences.
-Therefore, we write
-$\widetilde{J}=\left\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 3\textrm{ and }0\leq k\leq n\right\},$
-whose elements are acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}$, in order to write
-${R}_{1}(\widetilde{J})=\left\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\right\},$
-which is a set of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$.
-Hence, having the right lifting property with respect to the set ${R}_{1}(\widetilde{J})$ is necessary for a map of $1$-reduced simplicial sets to be a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
-We partially address the question of sufficiency.
-Namely, we establish that the set ${R}_{1}(\widetilde{J})$ detects fibrant objects in ${\mathcal{SS}\mathrm{ets}}_{1}$.
+    the maps ${R}_{1}(j_{2,0})$, ${R}_{1}(j_{2,1})$, and ${R}_{1}(j_{2,2})$ which are not weak equivalences.
+Therefore, we write $\widetilde{J}=\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 3\textrm{ and }0\leq k\leq n\}$, whose elements are acyclic cofibrations in $\mathcal{SS}\mathrm{ets}$}, in order to write
+${R}_{1}(\widetilde{J})=\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\},$
+[…]
 Proposition 4.8.
@@
   3. (3)
-    The unique map $c\colon X\rightarrow\Delta[0]$ has the right lifting property with respect to the set
-    ${R}_{1}(\widetilde{J})=\left\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\right\}.$
-Hence, the set ${R}_{1}(\widetilde{J})$ detects fibrant objects in ${\mathcal{SS}\mathrm{ets}}_{1}$.
+    The map $X\rightarrow\Delta[0]$ has the right lifting property with respect to the set
+    ${R}_{1}(\widetilde{J})=\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\}.$
+[…]
 Proof.
-Firstly, (1) and (2) are equivalent by the characterization of fibrant objects in ${\mathcal{SS}\mathrm{ets}}_{1}$ in Theorem 3.17.
-If (2) holds, then $c\colon X\rightarrow\Delta[0]$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
-Thus, $c\colon X\rightarrow\Delta[0]$ has the right lifting property with respect to the set ${R}_{1}(\widetilde{J})$, since every map in ${R}_{1}(\widetilde{J})$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
-Therefore, (3) holds.
-Lastly, suppose (3) holds. We prove that (1) holds. For $n\geq 1$ and $0\leq k\leq n$, a lifting problem
-${{\Lambda[n,k]}}$${X}$${{\Delta[n]}}$$\scriptstyle{h}$$\scriptstyle{j_{n,k}}$$\scriptstyle{g}$
-is equivalent, under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19, to a lifting problem
-${{{R}_{1}\left(\Lambda[n,k]\right)}}$${X}$${{{R}_{1}\left(\Delta[n]\right).}}$$\scriptstyle{\eta}$$\scriptstyle{{R}_{1}\left(j_{n,k}\right)}$$\scriptstyle{\psi}$
-In light of Lemma 3.13, we split our solution in three separate cases.
-If $n=1$, then Lemma 3.13(1) says that ${R}_{1}\left(j_{1,k}\right)$ is the identity map of the point, so our lifting problem can be solved by setting $\psi=\eta$.
-If $n\geq 3$, then our lifting problem can be solved, as we are assuming that $c\colon X\rightarrow\Delta[0]$ has the right lifting property with respect to the set ${R}_{1}(\widetilde{J})$.
-Lastly, if $n=2$, then Lemma 3.13(2) says that our lifting problem is
 […]
-where $x_{0}$ denotes the unique 00-simplex of $X$, so we take $\psi$ to be the constant map at $x_{0}$. ∎
+[…]
+is equivalent, under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Proposition 2.15, to a lifting problem
+[…]
+[…]
+[…]
+where $x_{0}$ denotes the unique 0-simplex of $X$, so we take $\psi$ to be the constant map at $x_{0}$. ∎
 In fact, a different and more elaborate argument yields a stronger result.
-Proposition 4.10 ([Qui69, Part II, Proposition 2.12]).
-Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets. Suppose that $Y$ is a Kan complex. The following statements are equivalent.
-  1. (1)
-    The map $f\colon X\rightarrow Y$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
-  2. (2)
-    The map $f\colon X\rightarrow Y$ has the right lifting property with respect to the set
-    ${R}_{1}(\widetilde{J})=\left\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\right\}.$
-Hence, the set ${R}_{1}(\widetilde{J})$ detects fibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ whose target is a Kan complex.
-The additional fibrancy condition on the target is crucially used in Quillen’s proof, which suggests that ${R}_{1}(\widetilde{J})$ is not a generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$.
-In his discussion of [Qui69, Part II, Proposition 2.12], Quillen affirms this suggestion and provides a hint towards showing that the fibrancy condition on the target in Proposition 4.10 is necessary.
-However, to the best of our knowledge, the question of whether ${R}_{1}(\widetilde{J})$ is a generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$ remains open.
-Our various attempts at producing a map in ${\mathcal{SS}\mathrm{ets}}_{1}$ that has the right lifting property with respect to ${R}_{1}(\widetilde{J})$ but is not a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ have been unsuccessful in the same fashion: our candidate maps are not fibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$, but they also fail to have the right lifting property with respect to ${R}_{1}(\widetilde{J})$.
-We pose the following conjecture to the community.
+Proposition 4.9 ([33, Part II, Proposition 2.12]).
+[…]
+[…]
 Conjecture 4.10.
 Let ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightarrow{\mathcal{SS}\mathrm{ets}}_{1}$ denote the $1$-reduction functor. The set
-${R}_{1}(\widetilde{J})=\left\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\right\}$
+${R}_{1}(\widetilde{J})=\{{R}_{1}(j_{n,k})\colon{R}_{1}(\Lambda[n,k])\rightarrow{R}_{1}(\Delta[n])\mid n\geq 3\textrm{ and }0\leq k\leq n\}$
 is not a generating set of acyclic cofibrations for ${\mathcal{SS}\mathrm{ets}}_{1}$.
-4.4. Fibrations of $1$-reduced simplicial sets
+4.4. Pathological fibrations of $1$-reduced simplicial sets
 The issues that we ran into in our study of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ suggest that there exist fibrations in ${\mathcal{SS}\mathrm{ets}}_{1}$ that are dissimilar to Kan fibrations.
 To investigate this phenomenon, we characterize when a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ is a Kan fibration.
-In fact, we produce two results to that end, which help us find an explicit example of a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ that is not a Kan fibration.
+In fact, we produce two results that help us find an explicit example of a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ that is not a Kan fibration.
 To begin with, we note that every Kan fibration of $1$-reduced simplicial sets is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
@@
 ∎
-We prove a characterization that detects when a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ is a Kan fibration and helps us disprove the converse of Proposition 4.12.
-Our result is the $1$-reduced analog of [GJ99, Lemma 6.6].
-Recall from Lemma 3.13 that, for $0\leq k\leq 2$, the map ${R}_{1}\left(j_{2,k}\right)\colon\Delta[0]\rightarrow S_{\mathrm{simp}}^{2}$ picks the unique 00-simplex of the simplicial $2$-sphere.
+We prove a characterization that detects when a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ is a Kan fibration and helps us disprove the converse of Proposition 4.11; our result is the $1$-reduced analog of [23, Lemma 6.6].
+Recall from Lemma 3.13 that, for $0\leq k\leq 2$, the map ${R}_{1}(j_{2,k})\colon\Delta[0]\rightarrow S_{\mathrm{simp}}^{2}$ picks the unique 0-simplex of the simplicial $2$-sphere.
 Proposition 4.12.
-Let $f\colon X\rightarrow Y$ be a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
+Let $f$ be a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
 For $0\leq k\leq 2$, let $j_{2,k}\colon\Lambda[2,k]\rightarrow\Delta[2]$ denote the horn inclusion.
-The following statements are equivalent.
-  1. (1)
-    The map $f\colon X\rightarrow Y$ is a Kan fibration.
-  2. (2)
-    The map $f\colon X\rightarrow Y$ has the right lifting property with respect to ${R}_{1}\left(j_{2,k}\right)$ for $0\leq k\leq 2$.
+Then, $f$ is a Kan fibration if and only if $f$ has the right lifting property against ${R}_{1}(j_{2,k})$ for $0\leq k\leq 2$.
 Proof.
-Under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19, a lifting problem of the type
+Under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Proposition 2.15, a lifting problem of the type
 […]
 is equivalent to a lifting problem of the type
-(II) ${{\Lambda[2,k]}}$${X}$${{\Delta[2]}}$${{Y.}}$$\scriptstyle{j_{2,k}}$$\scriptstyle{f}$
+[…]
 If $f$ is a Kan fibration, then every lifting problem of type (II) has a solution.
-Thus, every lifting problem of type (I) has a solution, and $f$ has the right lifting property with respect to ${R}_{1}\left(j_{2,k}\right)$ for $0\leq k\leq 2$.
-Conversely, if $f$ has the right lifting property with respect to ${R}_{1}\left(j_{2,k}\right)$ for $0\leq k\leq 2$, then every lifting problem of type (II) has a solution.
+Thus, every lifting problem of type (I) has a solution, and $f$ has the right lifting property with respect to ${R}_{1}(j_{2,k})$ for $0\leq k\leq 2$.
+Conversely, if $f$ has the right lifting property with respect to ${R}_{1}(j_{2,k})$ for $0\leq k\leq 2$, then every lifting problem of type (II) has a solution.
 If $n\neq 2$ and $0\leq t\leq n$, then solving a lifting problem
-${{\Lambda[n,t]}}$${X}$${{\Delta[n]}}$${{Y}}$$\scriptstyle{j_{n,t}}$$\scriptstyle{f}$
-is equivalent, under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Lemma 2.19, to solving a lifting problem
-${{{R}_{1}\left(\Lambda[n,t]\right)}}$${X}$${{{R}_{1}\left(\Delta[n]\right)}}$${{Y.}}$$\scriptstyle{{R}_{1}\left(j_{n,t}\right)}$$\scriptstyle{f}$
-By Lemma 4.8, ${R}_{1}\left(j_{n,t}\right)$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ because $n\neq 2$, so our lifting problem has a solution because $f$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
+[…]
+is equivalent, under the adjunction ${R}_{1}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon U$ from Proposition 2.15, to solving a lifting problem
+[…]
+By Lemma 4.7, ${R}_{1}(j_{n,t})$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ because $n\neq 2$, so our lifting problem has a solution because $f$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
 Overall, we conclude that $f$ is a Kan fibration.
@@
 Lemma 4.13.
-Let $B$ be a $1$-reduced simplicial set. Then, a map $h\colon B\rightarrow S_{\mathrm{simp}}^{2}$ is constant if and only if
-$h_{\ast}\colon\pi_{2}(B)\rightarrow\pi_{2}(S_{\mathrm{simp}}^{2})$
-is not an epimorphism.
+Let $B$ be a $1$-reduced simplicial set.
+Then, a map $h\colon B\rightarrow S_{\mathrm{simp}}^{2}$ is constant if and only if $h_{\ast}\colon\pi_{2}(B)\rightarrow\pi_{2}(S_{\mathrm{simp}}^{2})$ is not an epimorphism.
 Proof.
@@
 If $h$ is not constant, then there is a $2$-simplex $\omega$ of $B$ such that $h(\omega)$ is the unique non-degenerate $2$-simplex $\epsilon$ of $S_{\mathrm{simp}}^{2}$.
-As $B$ is $1$-reduced, Corollary 2.20 allows us to view $\omega$ as a map
-$\omega\colon S_{\mathrm{simp}}^{2}\rightarrow B.$
-Then, we compute that
-$h_{\ast}\left([\omega]\right)=[h(\omega)]=[\epsilon],$
-which is a generator of $\pi_{2}(S_{\mathrm{simp}}^{2})\cong\mathbb{Z}$. Thus, $h_{\ast}$ is an epimorphism. ∎
+As $B$ is $1$-reduced, Corollary 2.16 allows us to view $\omega$ as a map $\omega\colon S_{\mathrm{simp}}^{2}\rightarrow B$.
+Then, we compute that $h_{\ast}([\omega])=[h(\omega)]=[\epsilon]$, which is a generator of $\pi_{2}(S_{\mathrm{simp}}^{2})\cong\mathbb{Z}$.
+∎
 Proposition 4.14.
-For $0\leq k\leq 2$, the map of $1$-reduced simplicial sets ${R}_{1}\left(j_{2,k}\right)\colon\Delta[0]\rightarrow S_{\mathrm{simp}}^{2}$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ but not a Kan fibration.
+For $0\leq k\leq 2$, the map of $1$-reduced simplicial sets ${R}_{1}(j_{2,k})\colon\Delta[0]\rightarrow S_{\mathrm{simp}}^{2}$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ but not a Kan fibration.
 Proof.
-We first show that ${R}_{1}\left(j_{2,k}\right)$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
+We first show that ${R}_{1}(j_{2,k})$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
 For an acyclic cofibration $\gamma\colon A\rightarrow B$ in ${\mathcal{SS}\mathrm{ets}}_{1}$ and a lifting problem
-${A}$${{\Delta[0]}}$${B}$${{S_{\mathrm{simp}}^{2},}}$$\scriptstyle{c}$$\scriptstyle{\gamma}$$\scriptstyle{{R}_{1}\left(j_{2,k}\right)}$$\scriptstyle{g}$$\scriptstyle{h}$
+[…]
 we claim that the unique option for $g$ solves our lifting problem.
@@
 To show that the lower right triangle also commutes, we show that $h$ is constant.
-By Lemma 4.14, we show that
-$h_{\ast}\colon\pi_{2}(B)\rightarrow\pi_{2}(S_{\mathrm{simp}}^{2})$
-is not an epimorphism; in fact, we show that $h_{\ast}$ is constant. Firstly, as $c_{\ast}$ is constant, so is
-$h_{\ast}\gamma_{\ast}=(h\gamma)_{\ast}=({{R}_{1}\left(j_{2,k}\right)}c)_{\ast}={{R}_{1}\left(j_{2,k}\right)}_{\ast}c_{\ast}.$
+By Lemma 4.13, we show that $h_{\ast}\colon\pi_{2}(B)\rightarrow\pi_{2}(S_{\mathrm{simp}}^{2})$ is not an epimorphism; in fact, we show that $h_{\ast}$ is constant.
+As $c_{\ast}$ is constant, so is
+$h_{\ast}\gamma_{\ast}=(h\gamma)_{\ast}=({{R}_{1}(j_{2,k})}c)_{\ast}={{R}_{1}(j_{2,k})}_{\ast}c_{\ast}.$
 Because $\gamma$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, we know that $\gamma_{\ast}$ is an isomorphism. Therefore, $h_{\ast}$ is constant.
-Now, we apply the characterization from Proposition 4.13 to argue that ${R}_{1}\left(j_{2,k}\right)$ is not a Kan fibration because it does not have the right lifting property with respect to itself.
+Now, we apply the characterization from Proposition 4.12 to argue that ${R}_{1}(j_{2,k})$ is not a Kan fibration because it does not have the right lifting property with respect to itself.
 Indeed, the lifting problem
-[…]
+$\Delta[0]$$\Delta[0]$$S_{\mathrm{simp}}^{2}$$S_{\mathrm{simp}}^{2}$$\scriptstyle{{R}_{1}(j_{2,k})}$$\scriptstyle{{R}_{1}(j_{2,k})}$$\scriptstyle{g}$
 has no solution because the unique option for $g$ does not make the lower right triangle commute. ∎
 In the presence of a fibrancy condition on the target, Proposition 4.12 can be strengthened.
-Namely, we establish the following characterization result, which is the $1$-reduced analog of [GJ99, Corollary 6.9].
+Namely, we establish the following characterization result, which is the $1$-reduced analog of [23, Corollary 6.9].
 Proposition 4.15.
@@
   2. (2)
-    The map $f\colon X\rightarrow Y$ has the right lifting property with respect to ${R}_{1}\left(j_{2,k}\right)$ for $0\leq k\leq 2$.
+    The map $f\colon X\rightarrow Y$ has the right lifting property with respect to ${R}_{1}(j_{2,k})$ for $0\leq k\leq 2$.
   3. (3)
     The map $f\colon X\rightarrow Y$ induces an epimorphism $f_{\ast}\colon\pi_{2}(X)\rightarrow\pi_{2}(Y)$.
-In our proof, we use the construction of the homotopy groups of Kan complexes from [GJ99, Chapter I.7].
 Proof.
-[…]
+The unique map $X\rightarrow\Delta[0]$ factors in ${\mathcal{SS}\mathrm{ets}}_{1}$ as the composite of fibrations $X\xrightarrow{f}Y\rightarrow\Delta[0]$.
+Hence, $X$ is fibrant in ${\mathcal{SS}\mathrm{ets}}_{1}$, thus a Kan complex by the characterization of fibrant objects in ${\mathcal{SS}\mathrm{ets}}_{1}$ from Theorem 3.17.
 Now, Proposition 4.12 says that (1) and (2) are equivalent.
 We show that (2) and (3) are equivalent.
-Firstly, suppose that $f$ has the right lifting property with respect to ${R}_{1}\left(j_{2,k}\right)$ for $0\leq k\leq 2$.
-As $Y$ is a Kan complex, every element of $\pi_{2}(Y)$ is of the form $[\omega]$ for a map
-$\omega\colon S_{\mathrm{simp}}^{2}\rightarrow Y.$
+Firstly, suppose that $f$ has the right lifting property with respect to ${R}_{1}(j_{2,k})$ for $0\leq k\leq 2$.
+As $Y$ is a Kan complex, every element of $\pi_{2}(Y)$ is of the form $\omega]$ for a map $\omega\colon S_{\mathrm{simp}}^{2}\rightarrow Y$.
 We solve the lifting problem
 […]
-to produce an element $[\alpha]\in\pi_{2}(X)$ such that
-$f_{\ast}([\alpha])=[f\alpha]=[\omega]\in\pi_{2}(Y).$
+to produce an element $\alpha]\in\pi_{2}(X)$ such that $f_{\ast}([\alpha])=[f\alpha]=[\omega]$.
 Conversely, suppose that $f$ induces an epimorphism at $\pi_{2}$. For $0\leq k\leq 2$, we solve a lifting problem
@@
 […]
-where
-$f_{\ast}([\beta])=[f\beta]=[\eta]\in\pi_{2}(Y)$
-by means of the pointed homotopy
-$H\colon S_{\mathrm{simp}}^{2}\wedge\Delta[1]_{+}\rightarrow Y$
-starting from $Hd^{1}=f\beta$ and ending at $Hd^{0}=\eta$.
+where $f_{\ast}([\beta])=[f\beta]=[\eta]$ by means of the homotopy $H\colon S_{\mathrm{simp}}^{2}\wedge\Delta[1]_{+}\rightarrow Y$ starting from $Hd^{1}=f\beta$ and ending at $Hd^{0}=\eta$.
 By Lemma 2.11, $S_{\mathrm{simp}}^{2}\wedge\Delta[1]_{+}$ is $1$-reduced.
@@
 […]
-Then, $\psi=gd^{0}$ solves our lifting problem because
-$f\psi=fgd^{0}=Hd^{0}=\eta.$
-The fibrancy condition in Proposition 4.16 is the reason why we could not have used it for our example of a pathological fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ from Proposition 4.15.
+Then, $\psi=gd^{0}$ solves our lifting problem because $f\psi=fgd^{0}=Hd^{0}=\eta$. ∎
+The fibrancy condition in Proposition 4.15 is the reason why we could not have used it to streamline our proof of Proposition 4.14.
 Indeed, we conclude this section by showing that, for $n\geq 2$, the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ is not a Kan complex.
@@
     If $x$ is a degenerate $n$-simplex, then $s_{k}(x)$ has no non-degenerate faces.
-Hence, in either case, $s_{k}(x)$ has at most $2$ non-degenerate faces.
+Thus, the degenerate $(n+1)$-simplex $s_{k}(x)$ has at most $2$ non-degenerate faces.
 Proof.
-For $10\leq i\leq n+$ 1, let $d_{i}s_{k}(x)$ be the $i$-th face of $s_{k}(x)$. The result follows from the simplicial identities:
+For $0\leq i\leq n+1$, let $d_{i}s_{k}(x)$ be the $i$-th face of $s_{k}(x)$. The result follows from the simplicial identities
 $d_{i}s_{k}(x)=\begin{cases}s_{k-1}d_{i}(x),\textrm{ if }0\leq i\leq k-1;\\ x,\textrm{ if }k\leq i\leq k+1;\\ s_{k}d_{i-1}(x),\textrm{ if }k+2\leq i\leq n+1.\end{cases}$
+Our counting lemma allows us to prove that, for $n\geq 2$, the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ is not a Kan complex.
 Proposition 4.17.
@@
 Proof.
-Write $\epsilon$ for the unique non-degenerate $n$-simplex of $S_{\mathrm{simp}}^{n}$. For $10\leq k\leq n+$ 1, consider the lifting problem
+Write $\epsilon$ for the unique non-degenerate $n$-simplex of $S_{\mathrm{simp}}^{n}$. For $0\leq k\leq n+1$, consider the lifting problem
 $\Lambda[n+1,k]$$S_{\mathrm{simp}}^{n}$$\Delta[n+1],$$\scriptstyle{h}$$\scriptstyle{j_{n+1,k}}$$\scriptstyle{g}$
-where $h$ is the horn that picks out the $n$-simplices $\left(\epsilon,\dots,\epsilon\right)$.
+where $h$ is the horn that picks out the $n$-simplices $(\epsilon,\dots,\epsilon)$.
 Any simplex $g$ that fills in this horn would be a degenerate $(n+1)$-simplex with $n+1$ non-degenerate faces, which is impossible for $n\geq 2$ by Lemma 4.16.
@@
 A convenient feature of the model categories $\mathcal{SS}\mathrm{ets}$} and $\mathcal{SS}\mathrm{ets}$$_{\ast}$ is that they are simplicial, so they carry more structure for us to use.
-This feature motivates us to endow ${\mathcal{SS}\mathrm{ets}}_{1}$ with a simplicial model structure.
-We modify the simplicial model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ to define function complexes, tensors, and cotensors of $1$-reduced simplicial sets.
-Then, we verify that our definitions endow ${\mathcal{SS}\mathrm{ets}}_{1}$ with a simplicial model structure.
+In particular, between any two objects, we have an explicit construction of a function complex, which facilitates the study of left Bousfield localizations of these model categories.
+In this section, we endow ${\mathcal{SS}\mathrm{ets}}_{1}$ with a simplicial model structure.
+This model structure makes our subsequent left Bousfield localization of ${\mathcal{SS}\mathrm{ets}}_{1}$ more convenient in practice.
+Moreover, as left Bousfield localizations preserve the simplicial structure of a model category, the entire family of model structures on $1$-reduced simplicial sets inherits the simplicial structure in this section, so we get significant traction out of our argument.
 5.1. Simplicial model categories
-A reference on simplicial model categories is [Hir03, Chapter 9].
-Definition 5.1 ([Hir03, Definition 9.1.2]).
-Let $\mathcal{M}$ be a model category that is enriched over simplicial sets; in particular, for two objects $X$ and $Y$ of $\mathcal{M}$, we have a function complex $\operatorname{Map}_{\mathcal{M}}\left(X,Y\right)$.
+A reference on simplicial model categories is [26, Chapter 9]. We reformulate axiom (2) in the definition from [26, Definition 9.1.2] for our purposes using [26, Proposition 9.3.7].
+Definition 5.1.
+Let $\mathcal{M}$ be a model category that is enriched in simplicial sets; in particular, for two objects $X$ and $Y$ of $\mathcal{M}$, we have a function complex $\operatorname{Map}_{\mathcal{M}}(X,Y)$.
 Then, $\mathcal{M}$ is a simplicial model category if $\mathcal{M}$ satisfies the following axioms.
   1. (1)
-    (Adjointness axiom) For two objects $X$ and $Y$ of $\mathcal{M}$ and a simplicial set $Z$, there is a tensor object $X\otimes Z$ of $\mathcal{M}$ and a cotensor object $Y^{Z}$ of $\mathcal{M}$ such that there are isomorphisms of simplicial sets
-    $\operatorname{Map}_{\mathcal{M}}\left(X\otimes Z,Y\right)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(Z,\operatorname{Map}_{\mathcal{M}}\left(X,Y\right)\right)\cong\operatorname{Map}_{\mathcal{M}}\left(X,Y^{Z}\right)$
-    that are natural in $X$, $Y$, and $Z$.
-Note that $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(Z,\operatorname{Map}_{\mathcal{M}}\left(X,Y\right)\right)$ is a function complex of simplicial sets; see Example 5.2.
+    (Adjointness axiom) For two objects $X$ and $Y$ of $\mathcal{M}$ and a simplicial set $Z$, there is a tensor object $X\otimes Z$ of $\mathcal{M}$ and a cotensor object $Y^{Z}$ of $\mathcal{M}$ satisfying the natural isomorphisms of simplicial sets
+    $\operatorname{Map}_{\mathcal{M}}(X\otimes Z,Y)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}(Z,\operatorname{Map}_{\mathcal{M}}(X,Y))\cong\operatorname{Map}_{\mathcal{M}}(X,Y^{Z}).$
+    Note that $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}(Z,\operatorname{Map}_{\mathcal{M}}(X,Y))$ is a function complex of simplicial sets; see Example 5.2.
   2. (2)
-    (Homotopy lifting-extension axiom) If $i\colon A\rightarrow B$ is a cofibration in $\mathcal{M}$ and $p\colon X\rightarrow Y$ is a fibration in $\mathcal{M}$, then the induced map
-    $\operatorname{Map}_{\mathcal{M}}\left(B,X\right)\xrightarrow{(i^{\ast},p_{\ast})}\operatorname{Map}_{\mathcal{M}}\left(A,X\right)\underset{\operatorname{Map}_{\mathcal{M}}\left(A,Y\right)}{\times}\operatorname{Map}_{\mathcal{M}}\left(B,Y\right)$
-    is a Kan fibration, which is also a weak homotopy equivalence if $i$ or $p$ is also a weak equivalence in $\mathcal{M}$.
-Example 5.2 ([Hir03, Example 9.1.13]).
+    (Homotopy lifting-extension axiom) If $i\colon A\rightarrow B$ is a cofibration in $\mathcal{M}$ and $j\colon L\rightarrow K$ is a cofibration of simplicial sets, then the induced morphism
+    $A\otimes K\underset{A\otimes L}{\coprod}B\otimes L\rightarrow B\otimes K$
+    is a cofibration in $\mathcal{M}$ that is acyclic if $i$ or $j$ is acyclic.
+We present the simplicial structure of the model categories $\mathcal{SS}\mathrm{ets}$} and $\mathcal{SS}\mathrm{ets}$$_{\ast}$.
+Example 5.2 ([26, Example 9.1.13]).
 The category $\mathcal{SS}\mathrm{ets}$} is a simplicial model category. Given simplicial sets $X$, $Y$, and $Z$, we define
   1. (1)
-    the function complex $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(X,Y\right)$ to be the simplicial set with $n$-simplices $\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}}\left(X\times\Delta[n],Y\right)$;
+    the function complex $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}(X,Y)$ to be the simplicial set with $n$-simplices $\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}}(X\times\Delta[n],Y)$;
   2. (2)
@@
   3. (3)
-    the cotensor of $X$ with $Z$ to be the simplicial set $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(Z,X\right)$.
-Example 5.3 ([Hir03, Example 9.1.14]).
+    the cotensor of $X$ with $Z$ to be the simplicial set $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}(Z,X)$.
+Example 5.3 ([26, Example 9.1.14]).
 The category $\mathcal{SS}\mathrm{ets}$$_{\ast}$ is a simplicial model category. For two pointed simplicial sets $X$ and $Y$ and a simplicial set $Z$, we define
   1. (1)
-    the function complex $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X,Y\right)$ as the simplicial set with $n$-simplices $\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge\Delta[n]_{+},Y\right)$;
+    the function complex $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)$ as the simplicial set with $n$-simplices $\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X\wedge\Delta[n]_{+},Y)$;
   2. (2)
-    the tensor of $X$ with $Z$ to be the pointed simplicial set $X\times Z_{+}$; and
+    the tensor of $X$ with $Z$ to be the pointed simplicial set $X\wedge Z_{+}$; and
   3. (3)
-    the cotensor of $X$ with $Z$ to be the pointed simplicial set $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(Z_{+},X\right)$.
-5.2. Function complexes, tensors, and cotensors of $1$-reduced simplicial sets
-We modify the simplicial model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ to define function complexes, tensors, and cotensors in ${\mathcal{SS}\mathrm{ets}}_{1}$. We define
+    the cotensor of $X$ with $Z$ to be the pointed simplicial set $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},X)$.
+5.2. The simplicial model structure
+As our left transfer reveals to us, the homotopy theory of $1$-reduced simplicial sets is closer to that of pointed simplicial sets.
+Thus, it is the simplicial structure from Example 5.3 that we modify to get a simplicial structure on ${\mathcal{SS}\mathrm{ets}}_{1}$.
+In fact, our modification guarantees that the homotopy lifting-extension axiom for ${\mathcal{SS}\mathrm{ets}}_{1}$ is identical to that for $\mathcal{SS}\mathrm{ets}$$_{\ast}$, so we circumvent what is often the most difficult part of endowing a model category with a simplicial structure.
+For the simplicial structure on the model category ${\mathcal{SS}\mathrm{ets}}_{1}$, we define
   1. (1)
@@
   3. (3)
-    the cotensor of a $1$-reduced simplicial set $X$ with a simplicial set $Z$ to be $E_{1}\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},X)\right)$.
-5.3. An exponential law
-The simplicial model structure on ${\mathcal{SS}\mathrm{ets}}$ is underpinned by the exponential law for simplicial sets [GJ99, Chapter I, Proposition 5.1].
-We formulate an exponential law for pointed simplicial sets for our simplicial model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$.
-For two pointed simplicial sets $X$ and $Y$, the evaluation
-$\operatorname{ev}\colon X\wedge\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\rightarrow Y$
-is defined at the level of $n$-simplices by
-$\operatorname{ev}_{n}([x,f])=f_{n}([x,\sigma_{n}]),$
-where $\sigma_{n}$ is the unique non-degenerate $n$-simplex of $\Delta[n]$. We establish the following exponential law.
-Proposition 5.4.
-For pointed simplicial sets $X$, $Y$, and $Z$, the map
-$\operatorname{ev}_{\ast}\colon\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(Z,\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right)\rightarrow\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge Z,Y\right)$
-sending $f\colon Z\rightarrow\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)$ to the composite
-$X\wedge Z\xrightarrow{\operatorname{id}_{X}\wedge f}X\wedge\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\xrightarrow{\operatorname{ev}}Y$
-is a bijection that is natural in $X$, $Y$, and $Z$.
-Remark 5.5.
-This natural bijection amounts to the adjunction $X\wedge-\colon{\mathcal{SS}\mathrm{ets}}_{\ast}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,-)$.
-Proof.
-The inverse of $\operatorname{ev}_{\ast}$ sends a map $g\colon X\wedge Z\rightarrow Y$ to the map
-$\hat{g}\colon Z\rightarrow\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)$
-defined by sending an $n$-simplex $z$ of $Z$, which we view as a map $z\colon\Delta[n]\rightarrow Z$, to the composite
-[…]
-where, since $Z$ is pointed, we use Lemma 2.8 to induce the pointed map $\widetilde{z}\colon\Delta[n]_{+}\rightarrow Z$. ∎
-In particular, we use the following consequence of our exponential law.
-Corollary 5.6.
-For a $1$-reduced simplicial set $X$, there is an adjunction
-$X\wedge-\colon{\mathcal{SS}\mathrm{ets}}_{\ast}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{1}\colon\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,-).$
-Proof.
-By Lemma 2.15, the functor $X\wedge-$ maps into ${\mathcal{SS}\mathrm{ets}}_{1}$. The adjunction now follows from Remark 5.5. ∎
-5.4. Verification of the adjointness axiom
-We use our exponential law to verify that the structure defined in Section 5.2 satisfies the adjointness axiom. We need an auxiliary lemma that is straightforward to check.
-Lemma 5.7.
-Given two simplicial sets $Z$ and $W$, there is an isomorphism of pointed simplicial sets
-$\left(Z_{+}\right)\wedge\left(W_{+}\right)\cong\left(Z\times W\right)_{+}$
-that is natural in $Z$ and $W$.
-We are ready to verify the adjointness axiom for ${\mathcal{SS}\mathrm{ets}}_{1}$.
+    the cotensor of a $1$-reduced simplicial set $X$ with a simplicial set $Z$ to be $E_{1}(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},X))$.
+With our definition of function complexes, ${\mathcal{SS}\mathrm{ets}}_{1}$ is enriched in simplicial sets by Example 5.3.
+The homotopy lifting-extension axiom also holds by Example 5.3, so it remains to verify the adjointness axiom.
 Proposition 5.4.
 Given two $1$-reduced simplicial sets $X$ and $Y$ and a simplicial set $Z$, there are isomorphisms of simplicial sets
-$\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge Z_{+},Y\right)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(Z,\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X,E\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)\right)\right)$
+$\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X\wedge Z_{+},Y)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}(Z,\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y))\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,E_{1}(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)))$
 that are natural in $X$, $Y$, and $Z$.
 Proof.
-To prove $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge Z_{+},Y\right)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}\left(Z,\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right)$, we compute for arbitrary $n$ that
-$\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(\left(X\wedge Z_{+}\right)\wedge\Delta[n]_{+},Y\right)$ $\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge\l\left(Z_{+}\wedge\Delta[n]_{+}\right),Y\right)$
-$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(Z_{+}\wedge\Delta[n]_{+},\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right)$
-$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(\left(Z\times\Delta[n]\right)_{+},\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right)$
-$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}}\left(Z\times\Delta[n],\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y)\right).$
-The second bijection uses Proposition 5.4, the third one uses Lemma 5.7, and the fourth one uses Lemma 2.8.
-To prove $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge Z_{+},Y\right)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X,E\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)\right)\right)$, we compute for arbitrary $n$ that
-$\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(\left(X\wedge Z_{+}\right)\wedge\Delta[n]_{+},Y\right)$ $\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(Z_{+}\wedge\left(X\wedge\Delta[n]_{+}\right),Y\right)$
-$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge\Delta[n]_{+},\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)\right)$
-$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{1}}\left(X\wedge\Delta[n]_{+},E\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)\right)\right)$
-$=\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(X\wedge\Delta[n]_{+},E\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)\right)\right).$
-The second bijection comes from Proposition 5.4, and the third bijection is a consequence of Lemma 2.23(2). ∎
-5.5. Saturated classes
-To verify the homotopy lifting-extension axiom, we adapt the proof of the homotopy lifting-extension axiom for simplicial sets in [GJ99, Chapter I, Proposition 5.2].
-The proof for simplicial sets uses saturated classes of monomorphisms of simplicial sets, which are defined as follows.
-Definition 5.9 ([GZ67, Chapter IV, Definition 2.1]).
-A class $S$ of monomorphisms of simplicial sets is saturated if the following conditions are met.
-  1. (1)
-    All isomorphisms are contained in $S$.
-  2. (2)
-    In a pushout square
-[…]
-    if $m$ is contained in $S$, then $m^{\prime}$ is contained in $S$.
-  3. (3)
-    In a section-retraction commutative diagram
-[…]
-    if $m$ is contained in $S$, then $m^{\prime}$ is contained in $S$.
-  4. (4)
-    Given a sequence of composable maps
-    $A_{1}\xrightarrow{m_{1}}A_{2}\xrightarrow{m_{2}}\cdots,$
-    if $m_{t}$ is contained in $S$ for every positive integer $t$, then so is $A_{1}\rightarrow\operatorname{colim}(A_{n})$.
-  5. (5)
-    Given an indexing set $L$ and $m_{l}\colon A_{l}\rightarrow B_{l}$ for $l\in L$, if $m_{l}$ is contained in $S$ for every $l\in L$, then
-    $\coprod\limits_{l\in L}m_{l}\colon\coprod\limits_{l\in L}A_{l}\rightarrow\coprod\limits_{l\in L}B_{l}$
-    is contained in $S$.
-As the intersection of saturated classes forms another saturated class, the following definition is well-posed.
-Definition 5.10.
-The saturation of a class $P$ of monomorphisms of simplicial sets is
-$\langle P\rangle=\bigcap\limits_{\begin{subarray}{c}S\supset P\textrm{ and}\\ S\textrm{ is saturated }\end{subarray}}S.$
-Saturated classes provide us with descriptions of the class of acyclic cofibrations of simplicial sets.
-Proposition 5.11 ([GJ99, Chapter I, Proposition 4.2]).
-Consider the monomorphism classes of simplicial sets
-  1. (1)
-    $P_{1}=\left\{j_{n,k}\colon\Lambda[n,k]\rightarrow\Delta[n]\mid n\geq 1\textrm{and }0\leq k\leq n\right\}$; and
-  2. (2)
-    $P_{2}=\left\{\left(\Delta[1]\times\partial\Delta[n]\right)\cup\left(\{e\}\times\Delta[n]\right)\rightarrow\Delta[1]\times\Delta[n]\mid e\in\{0,1\}\textrm{ and }n\geq 0\right\}$.
-Then, $\langle P_{1}\rangle=\langle P_{2}\rangle$ is the class of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}$.
-Attaching a disjoint basepoint yields a description of the class of acyclic cofibrations of pointed simplicial sets.
-Corollary 5.12.
-Consider the monomorphism classes of pointed simplicial sets
-  1. (1)
-    $(P_{1})_{+}=\left\{(j_{n,k})_{+}\colon\Lambda[n,k]_{+}\rightarrow\Delta[n]_{+}\mid n\geq 1\textrm{ and }0\leq k\leq n\right\}$; and
-  2. (2)
-    $(P_{2})_{+}=\left\{\left(\left(\Delta[1]\times\partial\Delta[n]\right)\cup\left(\{e\}\times\Delta[n]\right)\right)_{+}\rightarrow\left(\Delta[1]\times\Delta[n]\right)_{+}\mid e\in\{0,1\}\textrm{ and }n\geq 0\right\}$.
-Then, $\langle(P_{1})_{+}\rangle=\langle(P_{2})_{+}\rangle$ is the class of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{\ast}$.
-We use Corollary 5.12 to prove a lemma that underpins our verification of the homotopy lifting-extension axiom.
-Lemma 5.13.
-For two pointed simplicial maps $i\colon A\rightarrow B$ and $j\colon X\rightarrow Y$, if both of them are monomorphisms, then the induced map
-$\left(A\wedge Y\right)\underset{\left(A\wedge X\right)}{\bigcup}\left(B\wedge X\right)\rightarrow B\wedge Y$
-is a monomorphism, which is also a weak homotopy equivalence if $i$ or $j$ is also a weak homotopy equivalence.
-Proof.
-If both $i$ and $j$ are monomorphisms, then so is the induced map.
-To prove the second part of the claim, the key idea is to write $P$ for the class of monomorphisms $A^{\prime}\rightarrow B^{\prime}$ of pointed simplicial sets such that
-$\left(A^{\prime}\wedge Y\right)\underset{\left(A\wedge X\right)}{\bigcup}\left(B^{\prime}\wedge X\right)\rightarrow B^{\prime}\wedge Y$
-is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{\ast}$.
-We show that $P$ contains the class of acyclic cofibrations in ${\mathcal{SS}\mathrm{ets}}_{\ast}$.
-As $P$ is a saturated class, thus equal to its own saturation, Corollary 5.12 truncates our task to showing that
-$(P_{2})_{+}\subset P.$
-Thus, for $e\in\{0,1\}$ and $n\geq 0$, we need to show that
-$\left(\left(\left(\Delta[1]\times\partial\Delta[n]\right)\cup\left(\{e\}\times\Delta[n]\right)\right)_{+}\wedge Y\right)\cup\left(\left(\Delta[1]\times\Delta[n]\right)_{+}\wedge X\right)\rightarrow\left(\Delta[1]\times\Delta[n]\right)_{+}\wedge Y$
-is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{\ast}$. We use Lemma 5.7 to write down the commutative diagram
-[…]
-where the upper pair of vertical isomorphisms is an instance of Lemma 5.7 and
-$\left(\Delta[1]_{+}\wedge\left(\left(\partial\Delta[n]_{+}\wedge Y\right)\cup\left(\Delta[n]_{+}\wedge X\right)\right)\right)\cup\left(\{e\}_{+}\wedge\left(\Delta[n]_{+}\wedge Y\right)\right)\rightarrow\Delta[1]_{+}\wedge\left(\Delta[n]_{+}\wedge Y\right)$
-is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{\ast}$ because its geometric realization is the inclusion of a strong deformation retract in pointed topological spaces.
-Overall, we conclude that $(P_{2})_{+}\subset P$.
-∎
-5.6. Verification of the homotopy lifting-extension axiom
-We are now ready to verify that the structure defined in Section 5.2 also satisfies the homotopy lifting-extension axiom.
-Proposition 5.14.
-If $i\colon A\rightarrow B$ is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ and $p\colon X\rightarrow Y$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, then the induced map
-$\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,X\right)\xrightarrow{(i^{\ast},p_{\ast})}\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,X\right)\underset{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,Y\right)}{\times}\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,Y\right)$
-is a Kan fibration, which is also a weak homotopy equivalence if $i$ or $p$ is also a weak homotopy equivalence.
-Proof.
-To prove that the induced map is a Kan fibration, we solve a lifting problem of the form
-${{\Lambda[n,k]}}$${{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,X\right)}}$${{\Delta[n]}}$${{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,X\right)\underset{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,Y\right)}{\times}\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,Y\right).}}$$\scriptstyle{j_{n,k}}$$\scriptstyle{(i^{\ast},p_{\ast})}$
-The adjunction $(-)_{+}\colon{\mathcal{SS}\mathrm{ets}}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon V$ from Lemma 2.8 yields the equivalent lifting problem
-[…]
-We apply Corollary 5.6 to infer that this new lifting problem is, in turn, equivalent to the lifting problem
-${{\left(\Lambda[n,k]_{+}\wedge B\right)\underset{\left(\Lambda[n,k]_{+}\wedge A\right)}{\bigcup}\left(\Delta[n]_{+}\wedge A\right)}}$${X}$${{\Delta[n]_{+}\wedge B}}$${{Y.}}$$\scriptstyle{p}$
-By Lemma 5.13, the left-hand map is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ as $\left(j_{n,k}\right)_{+}$ is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{\ast}$.
-As $p$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, our lifting problem can be solved, so the induced map is a Kan fibration.
-If $i$ or $p$ is also a weak homotopy equivalence, then we prove that the induced map is also a weak homotopy equivalence. To that end, we solve a lifting problem
-${{\partial\Delta[n]}}$${{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,X\right)}}$${{\Delta[n]}}$${{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,X\right)\underset{\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(A,Y\right)}{\times}\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}\left(B,Y\right).}}$$\scriptstyle{i_{n}}$$\scriptstyle{(i^{\ast},p_{\ast})}$
-Another consecutive application of the adjunctions from Lemma 2.8 and Corollary 5.6 translates our lifting problem to the equivalent lifting problem
-${{\left(\partial\Delta[n]_{+}\wedge B\right)\underset{\left(\partial\Delta[n]_{+}\wedge A\right)}{\bigcup}\left(\Delta[n]_{+}\wedge A\right)}}$${X}$${{\Delta[n]_{+}\wedge B}}$${{Y.}}$$\scriptstyle{p}$
-By Lemma 5.13, the left-hand map is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$ because $\left(i_{n}\right)_{+}$ is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{\ast}$.
-Depending on whether $i$ or $p$ is also a weak homotopy equivalence, we have two cases to scrutinize.
-  1. (1)
-    If $i$ is also a weak homotopy equivalence, then, by Lemma 5.13, the left-hand map is an acyclic cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
-As $p$ is a fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, our lifting problem can be solved.
-  2. (2)
-    If $p$ is also a weak homotopy equivalence, then $p$ is an acyclic fibration in ${\mathcal{SS}\mathrm{ets}}_{1}$.
-As the left-hand map is a cofibration in ${\mathcal{SS}\mathrm{ets}}_{1}$, our lifting problem can be solved.
-In either case, the induced map is both a Kan fibration and a weak homotopy equivalence, as required. ∎
-By Propositions 5.8 and 5.14, our definitions in Section 5.2 endow ${\mathcal{SS}\mathrm{ets}}_{1}$ with a simplicial model structure.
+The natural isomorphism $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X\wedge Z_{+},Y)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}(Z,\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,Y))$ is a consequence of Example 5.3.
+To prove that $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X\wedge Z_{+},Y)\cong\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X,E_{1}(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)))$, we compute that
+$\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}((X\wedge Z_{+})\wedge\Delta[n]_{+},Y)$ $\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X\wedge\Delta[n]_{+},\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y))$
+$\cong\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{1}}(X\wedge\Delta[n]_{+},E_{1}(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y)))$
+$=\operatorname{Hom}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(X\wedge\Delta[n]_{+},E_{1}(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},Y))),$
+where the first bijection comes from Example 5.3 and the second bijection follows from Proposition 2.19(2). ∎
+Overall, our modification of Example 5.3 endows the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ with a simplicial structure.
 Theorem 5.5.
@@
   3. (3)
-    the cotensor of a $1$-reduced simplicial set $X$ with a simplicial set $Z$ to be $E_{1}\left(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},X)\right)$.
+    the cotensor of a $1$-reduced simplicial set $X$ with a simplicial set $Z$ to be $E_{1}(\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}_{\ast}}(Z_{+},X))$.
 With these definitions, ${\mathcal{SS}\mathrm{ets}}_{1}$ has the structure of a simplicial model category.
-Proof.
-With our definition of function complexes, ${\mathcal{SS}\mathrm{ets}}_{1}$ is enriched over simplicial sets as ${\mathcal{SS}\mathrm{ets}}_{\ast}$ is enriched over simplicial sets.
-By Propositions 5.8 and 5.14, our definitions satisfy the simplicial model category axioms.
-∎
+Lastly, we note that, in [12], Burke establishes an analogous simplicial structure for reduced simplicial sets, that is, simplicial sets with a unique 0-simplex; Burke’s line of argument is different from ours, though.
 6. Rational homotopy theory of $1$-connected spaces
@@
 We gather the background and auxiliary results from the rational homotopy theory of $1$-connected topological spaces that we need to recover the entirety of Quillen’s family of model structures on ${\mathcal{SS}\mathrm{ets}}_{1}$.
-In particular, we characterize $M$-local spaces and $M$-local homotopy equivalences in terms of function complexes, which prepares us for the left Bousfield localization in Section 7.
-Even though we do not expect these characterizations to be novel to the rational homotopy theory expert, we were unable to find them in the literature, so we provide a full proof.
-6.1. Localization of $1$-connected spaces
-We first review some concepts from [FHT01, Chapter 9].
+In particular, we describe the $M$-local spaces and $M$-local homotopy equivalences in terms of function complexes, which prepares us for the left Bousfield localization in Section 7.
+6.1. Rationalization of $1$-connected spaces
+We give a review from [39, Chapter 2] and [22, Chapter 9].
 Definition 6.1.
@@
   1. (1)
-    A map of $1$-connected spaces $f\colon X\rightarrow Y$ is an $M$-local homotopy equivalence if, for every $n\geq 2$, the induced map
-    $f_{\ast}\underset{\mathbb{Z}}{\otimes}\operatorname{id}_{M^{-1}\mathbb{Z}}\colon\pi_{n}(X)\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}\rightarrow\pi_{n}(Y)\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}$
-    is an isomorphism.
+    A map of $1$-connected spaces $f\colon X\rightarrow Y$ is an $M$-local homotopy equivalence if, for every $n\geq 2$, the map $f_{\ast}\otimes\operatorname{id}_{M^{-1}\mathbb{Z}}\colon\pi_{n}(X)\otimes M^{-1}\mathbb{Z}\rightarrow\pi_{n}(Y)\otimes M^{-1}\mathbb{Z}$ is an isomorphism.
   2. (2)
@@
   3. (3)
-    An $M$-localization of a $1$-connected space $Y$ is an $M$-local homotopy equivalence $f\colon Y\rightarrow Y_{M^{-1}\mathbb{Z}}$ with $M$-local target.
+    An $M$-localization of a $1$-connected space $Y$ is an $M$-local homotopy equivalence with $M$-local target $f\colon Y\rightarrow Y_{M^{-1}\mathbb{Z}}$.
 When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we speak of a rational homotopy equivalence, a rational space, and a rationalization.
 Remark 6.2.
-If $f\colon Y\rightarrow Y_{\mathbb{Q}}$ is a rationalization of $Y$, then, for every $n\geq 2$, $f$ induces the isomorphism
-$f_{\ast}\underset{\mathbb{Z}}{\otimes}\operatorname{id}_{\mathbb{Q}}\colon\pi_{n}(Y)\underset{\mathbb{Z}}{\otimes}\mathbb{Q}\xrightarrow{\cong}\pi_{n}(Y_{\mathbb{Q}})$
-that discards all torsion in the $n$-th homotopy group of $Y$ in a purely topological fashion.
+If $f\colon Y\rightarrow Y_{\mathbb{Q}}$ is a rationalization of $Y$, then, for every $n\geq 2$, $f$ induces the isomorphism $f_{\ast}\otimes\operatorname{id}_{\mathbb{Q}}\colon\pi_{n}(Y)\otimes\mathbb{Q}\xrightarrow{\cong}\pi_{n}(Y_{\mathbb{Q}})$ that discards torsion in the $n$-th homotopy group of $Y$ in a topological fashion.
+The following example features often in our arguments.
 Example 6.3.
 Every weak homotopy equivalence of $1$-connected spaces is an $M$-local homotopy equivalence. Also, every $M$-local homotopy equivalence of $M$-local spaces is a weak homotopy equivalence.
+What is more, the rationalization of the $n$-sphere in dimensions $n\geq 2$ is a building block for the rationalization of all $1$-connected spaces; as such, it is a cornerstone of our arguments in the remainder of this paper.
+We do not need the details of the geometric construction at the moment, though we present them in Section 9.2.
 Example 6.4.
 For $n\geq 2$, the $M$-local $n$-sphere $S^{n}_{M^{-1}\mathbb{Z}}$ is an $M$-local CW complex containing the $n$-sphere $S^{n}$ as a CW subcomplex, and the inclusion $S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$ is an $M$-localization of $S^{n}$.
-Every $1$-connected space can be $M$-localized uniquely up to homotopy equivalence as follows.
-Theorem 6.5 ([FHT01, Theorem 9.7]).
+Every $1$-connected space can be $M$-localized, and thus rationalized, uniquely up to homotopy equivalence.
+Theorem 6.5 ([22, Theorem 9.7]).
 Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $Y$ be a $1$-connected space.
   1. (1)
-    There exists a relative CW complex $\left(Y_{M^{-1}\mathbb{Z}},Y\right)$ with no 00-cells and no $1$-cells such that the inclusion $Y\rightarrow Y_{M^{-1}\mathbb{Z}}$ is an $M$-localization of $Y$.
+    There exists a relative CW complex $(Y_{M^{-1}\mathbb{Z}},Y)$ with no 0-cells and no $1$-cells such that the inclusion $Y\rightarrow Y_{M^{-1}\mathbb{Z}}$ is an $M$-localization of $Y$.
   2. (2)
@@
 Remark 6.6.
-Note that, if $Y$ is a $1$-connected CW complex, then so is $Y_{M^{-1}\mathbb{Z}}$.
-6.2. A characterization of local spaces
-For a multiplicative subset $M$ of $\mathbb{Z}$, we characterize $M$-local spaces in a way that supports our model-categorical arguments for the left Bousfield localization in Section 7.
-For two topological spaces $X$ and $Y$, the function complex
-$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Y\right)$
-is the simplicial set with $n$-simplices
-$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Y\right)_{n}=\operatorname{Hom}_{\mathcal{T}\mathrm{op}}\left(X\times|\Delta[n]|,Y\right).$
-In particular, observe that
-$\pi_{0}\left(\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Y\right)\right)=\left[X,Y\right]$
-is the set of homotopy classes of maps from $X$ to $Y$.
-The following lemma is a straightforward application of the universal property of $M$-localization.
+For future reference, note that, if $Y$ is a $1$-connected CW complex, then so is $Y_{M^{-1}\mathbb{Z}}$.
+6.2. A characterization of rational spaces
+For a multiplicative subset $M$ of $\mathbb{Z}$, we characterize $M$-local spaces, and thus rational spaces, in a way that supports our left Bousfield localization in Section 7.
+Namely, we describe when a $1$-connected space is $M$-local in the language of function complexes.
+For two topological spaces $X$ and $Y$, the function complex $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,Y)$ is the simplicial set with $n$-simplices $\operatorname{Hom}_{\mathcal{T}\mathrm{op}}(X\times|\Delta[n]|,Y)$.
+Observe that
+$\pi_{0}(\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,Y))=[X,Y]$
+is the set of homotopy classes of maps from $X$ to $Y$. The following lemma is a straightforward application of the universal property of $M$-localization.
 Lemma 6.7.
-Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $X$ and $Y$ be $1$-connected spaces. If $Y$ is $M$-local, then the map
-$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Y\right)$
-induced by the inclusion $X\rightarrow X_{M^{-1}\mathbb{Z}}$ is a weak homotopy equivalence.
-We are now ready to establish a characterization of $M$-local spaces in the language of function complexes.
+Let $M$ be a multiplicative subset of $\mathbb{Z}$.
+Let $X$ and $Y$ be $1$-connected spaces.
+If $Y$ is $M$-local, then the restriction map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X_{M^{-1}\mathbb{Z}},Y)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,Y)$ is a weak equivalence.
+We are now ready to characterize $M$-local spaces, and thus rational spaces, using function complexes.
 Proposition 6.8.
-Let $M$ be a multiplicative subset of $\mathbb{Z}$. A $1$-connected space $Y$ is $M$-local if and only if, for every $n\geq 2$, the map
-$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n},Y\right)$
-induced by the inclusion $S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$ is a weak homotopy equivalence.
+Let $M$ be a multiplicative subset of $\mathbb{Z}$.
+A $1$-connected space $Y$ is $M$-local if and only if, for every $n\geq 2$, the restriction map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},Y)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},Y)$ is a weak equivalence.
 Proof.
-If $Y$ is $M$-local, then Lemma 6.7 implies that, for $n\geq 2$, the map
-$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n},Y\right)$
-is a weak homotopy equivalence. Conversely, suppose that, for $n\geq 2$, the map
-$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n},Y\right)$
-is a weak homotopy equivalence.
-Let $\tau\colon S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$ denote the inclusion.
-Then, for every map $f\colon S^{n}\rightarrow Y$, there exists a map $g\colon S^{n}_{M^{-1}\mathbb{Z}}\rightarrow Y$, unique up to homotopy, such that the diagram
-[…]
-commutes up to homotopy.
+If $Y$ is $M$-local, then Lemma 6.7 implies that the restriction map is a weak equivalence.
+Conversely, suppose that, for $n\geq 2$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},Y)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},Y)$ is a weak equivalence.
 For $n\geq 2$, let $[f]$ be an element of $\pi_{n}(Y)$ represented by a map $f\colon S^{n}\rightarrow Y$, and let $m\in M$.
@@
 […]
-commutes up to homotopy. In particular, homotopy commutativity implies that
-$[f]=f_{\ast}\left([\operatorname{id}_{S^{n}}]\right)=g_{\ast}\tau_{\ast}\left([\operatorname{id}_{S^{n}}]\right)=g_{\ast}\left([\tau]\right)\in\pi_{n}(Y).$
-As $S^{n}_{M^{-1}\mathbb{Z}}$ is $M$-local, there exists a unique element
-$\dfrac{1}{m}\cdot[\tau]\in\pi_{n}(S^{n}_{M^{-1}\mathbb{Z}})\cong\pi_{n}(S^{n})\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}\cong\mathbb{Z}\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}\cong M^{-1}\mathbb{Z}.$
-Therefore, we uniquely define
-$\dfrac{1}{m}\cdot[f]=g_{\ast}\left(\dfrac{1}{m}\cdot[\tau]\right)\in\pi_{n}(Y).$
-We conclude that the $1$-connected space $Y$ is $M$-local. ∎
+commutes up to homotopy, where $\tau$ denotes the inclusion. In particular, homotopy commutativity implies that
+$f]=f_{\ast}([\operatorname{id}_{S^{n}}])=g_{\ast}\tau_{\ast}([\operatorname{id}_{S^{n}}])=g_{\ast}([\tau]).$
+As $S^{n}_{M^{-1}\mathbb{Z}}$ is $M$-local, there exists a unique element $\dfrac{1}{m}\cdot[\tau]$, so we uniquely define $\dfrac{1}{m}\cdot[f]=g_{\ast}\left(\dfrac{1}{m}\cdot[\tau]\right)$. ∎
 As our function complexes of $1$-reduced simplicial sets are pointed, we need the pointed version of Proposition 6.8.
 We write $\mathcal{T}\mathrm{op}_{\ast}$ for the category of pointed spaces.
-For two pointed spaces $X$ and $Y$, the function complex
-$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(X,Y\right)$
-is the simplicial set with $n$-simplices
-$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(X,Y\right)_{n}=\operatorname{Hom}_{\mathcal{T}\mathrm{op}_{\ast}}\left(X\wedge|\Delta[n]|_{+},Y\right).$
-Observe that
-$\pi_{0}(\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(X,Y\right))=\left[X,Y\right]_{\ast}$
-is the set of pointed homotopy classes of pointed maps from $X$ to $Y$.
-In terms of these function complexes, we state the pointed version of Proposition 6.8, which is proven identically.
+For two pointed spaces $X$ and $Y$, the function complex $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(X,Y)$ is the simplicial set with $n$-simplices $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(X,Y)_{n}=\operatorname{Hom}_{\mathcal{T}\mathrm{op}_{\ast}}(X\wedge|\Delta[n]|_{+},Y)$.
+Again, $\pi_{0}(\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(X,Y))=[X,Y]_{\ast}$ is the set of pointed homotopy classes of pointed maps from $X$ to $Y$.
+We state the pointed version of Proposition 6.8, which is proven identically.
 Proposition 6.9.
-Let $M$ be a multiplicative subset of $\mathbb{Z}$. A $1$-connected pointed space $Y$ is $M$-local if and only if, for every $n\geq 2$, the map
-$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(S^{n},Y\right)$
-induced by the inclusion $S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$ is a weak homotopy equivalence.
-6.3. A characterization of local homotopy equivalences
-For a multiplicative subset $M$ of $\mathbb{Z}$, we also characterize the $M$-local homotopy equivalences in preparation for the left Bousfield localization in Section 7.
-The following lemma, in conjunction with its pointed analog, underpins our arguments.
-Lemma 6.10 ([Hir03, Corollary 9.3.3(2)]).
+Let $M$ be a multiplicative subset of $\mathbb{Z}$.
+A $1$-connected pointed space $Y$ is $M$-local if and only if, for every $n\geq 2$, the restriction map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n}_{M^{-1}\mathbb{Z}},Y)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n},Y)$ is a weak equivalence.
+6.3. A characterization of rational homotopy equivalences
+For a multiplicative subset $M$ of $\mathbb{Z}$, we also characterize the $M$-local homotopy equivalences using function complexes in preparation for the left Bousfield localization in Section 7.
+The following lemma, together with its pointed analog, underpins our arguments.
+Lemma 6.10 ([26, Corollary 9.3.3(2)]).
   1. (1)
-    If $f\colon X\rightarrow Y$ is a weak homotopy equivalence of simplicial sets and $Z$ is a Kan complex, then
-    $f^{\ast}\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(Y,Z)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(X,Z)$
-    is a weak homotopy equivalence.
+    If $f\colon X\rightarrow Y$ is a weak equivalence of simplicial sets and $Z$ is a Kan complex, then $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(f,Z)\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(Y,Z)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(X,Z)$ is a weak equivalence.
   2. (2)
-    If $f\colon X\rightarrow Y$ is a weak homotopy equivalence of CW complexes and $Z$ is a topological space, then
-    $f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y,Z)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,Z)$
-    is a weak homotopy equivalence.
-We provide a series of characterizations of $M$-local homotopy equivalences in terms of function complexes.
+    If $f\colon X\rightarrow Y$ is a weak homotopy equivalence of CW complexes and $Z$ is a topological space, then $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(f,Z)\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y,Z)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,Z)$ is a weak equivalence.
+We characterize the $M$-local homotopy equivalences, and thus the rational homotopy equivalences, as follows.
 Proposition 6.11.
@@
   2. (2)
-    The induced map on $M$-localizations
-    $f_{M^{-1}\mathbb{Z}}\colon X_{M^{-1}\mathbb{Z}}\rightarrow Y_{M^{-1}\mathbb{Z}}$
-    is a weak homotopy equivalence.
+    The induced map on $M$-localizations $f_{M^{-1}\mathbb{Z}}\colon X_{M^{-1}\mathbb{Z}}\rightarrow Y_{M^{-1}\mathbb{Z}}$ is a weak homotopy equivalence.
   3. (3)
-    The induced map on $M$-localizations
-    $f_{M^{-1}\mathbb{Z}}\colon X_{M^{-1}\mathbb{Z}}\rightarrow Y_{M^{-1}\mathbb{Z}}$
-    is a homotopy equivalence.
+    The induced map on $M$-localizations $f_{M^{-1}\mathbb{Z}}\colon X_{M^{-1}\mathbb{Z}}\rightarrow Y_{M^{-1}\mathbb{Z}}$ is a homotopy equivalence.
   4. (4)
-    For every $M$-local space $Z$, the map
-    $f_{M^{-1}\mathbb{Z}}^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(Y_{M^{-1}\mathbb{Z}},Z\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X_{M^{-1}\mathbb{Z}},Z\right)$
-    is a weak homotopy equivalence.
+    For every $M$-local space $Z$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(f_{M^{-1}\mathbb{Z}},Z)$ is a weak equivalence.
   5. (5)
-    For every $M$-local space $Z$, the map
-    $f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(Y,Z\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Z\right)$
-    is a weak homotopy equivalence.
+    For every $M$-local space $Z$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(f,Z)$ is a weak equivalence.
 Proof.
@@
 Moreover, (4) and (5) are equivalent because, for an $M$-local space $Z$, we have the commutative diagram
-${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(Y_{M^{-1}\mathbb{Z}},Z\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X_{M^{-1}\mathbb{Z}},Z\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(Y,Z\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Z\right)}}$$\scriptstyle{f_{M^{-1}\mathbb{Z}}^{\ast}}$$\scriptstyle{f^{\ast}}$
-whose vertical maps are weak homotopy equivalences by Lemma 6.7.
-It remains to show that (3) and (4) are equivalent. If (3) holds, then Lemma 6.10(2) informs us that, for every ($M$-local) space $Z$, the map
-$f_{M^{-1}\mathbb{Z}}^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(Y_{M^{-1}\mathbb{Z}},Z\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X_{M^{-1}\mathbb{Z}},Z\right)$
-is a weak homotopy equivalence. Conversely, if (4) holds, then using the induced bijection of path components
-$f_{M^{-1}\mathbb{Z}}^{\ast}\colon\left[Y_{M^{-1}\mathbb{Z}},Z\right]\xrightarrow{\cong}\left[X_{M^{-1}\mathbb{Z}},Z\right]$
-for $Z=X_{M^{-1}\mathbb{Z}}$ and $Z=Y_{M^{-1}\mathbb{Z}}$ yields a homotopy inverse of $f_{M^{-1}\mathbb{Z}}$. ∎
-It is the equivalence of (1) and (5) in Proposition 6.11 that is relevant to our work in Section 7. Again, we need the pointed version of that equivalence, which is proven identically.
+$\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y_{M^{-1}\mathbb{Z}},Z)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X_{M^{-1}\mathbb{Z}},Z)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y,Z)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,Z)$}}
+whose vertical maps are weak equivalences by Lemma 6.7.
+It remains to show that (3) and (4) are equivalent.
+If (3) holds, then Lemma 6.10(2) implies that, for every space $Z$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(f_{M^{-1}\mathbb{Z}},Z)$ is a weak equivalence.
+Conversely, if (4) holds, then using the bijection of path components $f_{M^{-1}\mathbb{Z}},Z]\colon[Y_{M^{-1}\mathbb{Z}},Z]\xrightarrow{\cong}[X_{M^{-1}\mathbb{Z}},Z]$ for $Z=X_{M^{-1}\mathbb{Z}}$ and $Z=Y_{M^{-1}\mathbb{Z}}$ yields a homotopy inverse of $f_{M^{-1}\mathbb{Z}}$.
+∎
+It is the equivalence of (1) and (5) that is relevant to our work in Section 7, namely its pointed version.
 Proposition 6.12.
 Let $M$ be a multiplicative subset of $\mathbb{Z}$.
-A pointed map of $1$-connected CW complexes $f\colon X\rightarrow Y$ is an $M$-local homotopy equivalence if and only if, for every $M$-local pointed space $Z$, the map
-$f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(Y,Z\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(X,Z\right)$
-is a weak homotopy equivalence.
+A pointed map of $1$-connected CW complexes $f\colon X\rightarrow Y$ is an $M$-local homotopy equivalence if and only if, for every $M$-local pointed space $Z$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(f,Z)\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(Y,Z)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(X,Z)$ is a weak equivalence.
 7. The family of model structures on $1$-reduced simplicial sets
 Let $M$ be a multiplicative subset of $\mathbb{Z}$.
-We left Bousfield localize the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ to get a model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ whose weak equivalences are the $M$-local homotopy equivalences.
-In particular, we obtain a model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ whose weak equivalences are the rational homotopy equivalences.
-Thus, we recover the family of model structures on $1$-reduced simplicial sets from [Qui69, Part II, Theorem 2.2].
+We left Bousfield localize the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ to get a model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ in which the weak equivalences are the $M$-local homotopy equivalences and the fibrant objects are the $M$-local Kan complexes.
+In particular, we obtain a model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ whose weak equivalences are the rational homotopy equivalences and whose fibrant objects are the rational Kan complexes.
+Thus, we recover the family of model structures on $1$-reduced simplicial sets from [33, Part II, Theorem 2.2] together with the characterization of fibrant objects in [33, Part II, Corollary 2.6].
 These model structures inherit the generating cofibrations and the simplicial structure that we found for ${\mathcal{SS}\mathrm{ets}}_{1}$.
 7.1. Left Bousfield localization
-We give an overview of left Bousfield localization; a reference is [Hir03, Chapters 1-4]. Let $\mathcal{M}$ be a simplicial model category and $T$ be a set of morphisms in $\mathcal{M}$.
+We first provide an overview of left Bousfield localization; a reference is [26, Chapters 1-4].
+Let $\mathcal{M}$ be a simplicial model category; in particular, for two objects $X$ and $Y$, we have a function complex $\operatorname{Map}_{\mathcal{M}}(X,Y)$.
+Though one can left Bousfield localize model categories that are not simplicial using homotopy function complexes [26, Chapter 17], we only localize simplicial model categories in this paper, and our arguments are contingent upon our explicit understanding of the function complexes in our model categories.
+Let $T$ be a set of morphisms in $\mathcal{M}$; in practice, it is important that $T$ is not a proper class.
 Definition 7.1.
   1. (1)
-    An object $K$ of $\mathcal{M}$ is $T$-local if $K$ is fibrant in $\mathcal{M}$ and, for every morphism $\varphi\colon A\rightarrow B$ in $T$, the map
-    $\varphi^{\ast}\colon\operatorname{Map}_{\mathcal{M}}\left(B,K\right)\rightarrow\operatorname{Map}_{\mathcal{M}}\left(A,K\right)$
-    is a weak homotopy equivalence.
+    An object $K$ of $\mathcal{M}$ is $T$-local if $K$ is fibrant in $\mathcal{M}$ and, for every morphism $\lambda\colon A\rightarrow B$ in $T$, the map $\operatorname{Map}_{\mathcal{M}}(\lambda,K)\colon\operatorname{Map}_{\mathcal{M}}(B,K)\rightarrow\operatorname{Map}_{\mathcal{M}}(A,K)$ is a weak equivalence.
   2. (2)
-    A morphism $f\colon X\rightarrow Y$ in $\mathcal{M}$ is a $T$-local equivalence if, for every $T$-local object $K$, the map
-    $f^{\ast}\colon\operatorname{Map}_{\mathcal{M}}\left(Y,K\right)\rightarrow\operatorname{Map}_{\mathcal{M}}\left(X,K\right)$
-    is a weak homotopy equivalence.
+    A morphism $f\colon X\rightarrow Y$ in $\mathcal{M}$ is a $T$-local equivalence if, for every $T$-local object $K$, the map $\operatorname{Map}_{\mathcal{M}}(f,K)\colon\operatorname{Map}_{\mathcal{M}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{M}}(X,K)$ is a weak equivalence.
   3. (3)
-    A $T$-localization of an object $X$ of $\mathcal{M}$ is a $T$-local equivalence $f\colon X\rightarrow X_{T}$ with $T$-local target.
-Example 7.2.
+    A $T$-localization of an object $X$ of $\mathcal{M}$ is a $T$-local equivalence with $T$-local target $f\colon X\rightarrow X_{T}$.
+Remark 7.2.
 Every morphism in $T$ and every weak equivalence in $\mathcal{M}$ is a $T$-local equivalence.
-Theorem 7.3 ([Bar10, Theorem 4.7]).
+Conversely, every $T$-local equivalence of $T$-local objects is a weak equivalence in $\mathcal{M}$ [26, Theorem 3.2.13(1)].
+In light of our remark, the left Bousfield localization of $\mathcal{M}$ at $T$ enlarges the class of weak equivalences to incorporate the maps in $T$ while preserving the class of cofibrations.
+Consequently, the class of fibrations, and thus the class of fibrant objects, is forced to shrink.
+Though the fibrations in a left Bousfield localization can be quite difficult to describe, at least the fibrant objects are always known to us.
+Theorem 7.3 ([2, Theorem 4.7]).
 Let $\mathcal{M}$ be a combinatorial, left proper, and simplicial model category.
 Let $T$ be a set of morphisms in $\mathcal{M}$.
-There is a model structure $\mathcal{L}_{T}\mathcal{M}$ on the underlying category of $\mathcal{M}$, called the left Bousfield localization of $\mathcal{M}$ at $T$, in which a morphism $f\colon X\rightarrow Y$ is
+There is a model structure $\mathcal{L}_{T}\mathcal{M}$ on the underlying category of $\mathcal{M}$, called the left Bousfield localization of $\mathcal{M}$ at $T$, in which a morphism $f$ is
   1. (1)
@@
 Note that a $T$-localization in $\mathcal{M}$ is a fibrant replacement in $\mathcal{L}_{T}\mathcal{M}$.
-The left Bousfield localization of $\mathcal{M}$ at $T$ enlarges, in a systematic fashion, the class of weak equivalences of $\mathcal{M}$ to encompass the morphisms in the set $T$.
-Observe that $\mathcal{L}_{T}\mathcal{M}$ inherits the generating cofibrations of $\mathcal{M}$, but it generally does not also inherit the generating acyclic cofibrations of $\mathcal{M}$ as, in general, the class of acyclic cofibrations in $\mathcal{M}$ is properly contained in the class of acyclic cofibrations in $\mathcal{L}_{T}\mathcal{M}$.
+As the class of cofibrations is preserved, $\mathcal{L}_{T}\mathcal{M}$ inherits the generating cofibrations of the combinatorial model category $\mathcal{M}$.
+However, the same is not true for the generating acyclic cofibrations because the class of weak equivalences is expanded and, in general, it is difficult to describe the smaller class of fibrations in a left Bousfield localization.
+Knowing the fibrant objects in a left Bousfield localization makes it feasible to write down a set of maps that detects them by tweaking any set that detects fibrant objects in $\mathcal{M}$; however, this set can get cumbersome to write out, which defeats its practical purpose.
+In our left Bousfield localizations in this paper, we have no nice description of such a set that detects fibrant objects, so we avoid writing one out altogether.
 7.2. The set for the left Bousfield localization
 For a multiplicative subset $M$ of $\mathbb{Z}$, our characterization of $M$-local spaces from Proposition 6.9 suggests that we produce a $1$-reduced simplicial analog of the set of maps
-$\left\{S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}\mid n\geq 2\right\}$
+$\{S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}\mid n\geq 2\}$
 in order to left Bousfield localize the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ at that set.
-A sensible first suggestion is to start with the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ and use it to emulate the construction of $S^{n}_{M^{-1}\mathbb{Z}}$ in the $1$-reduced setting.
-This approach runs into a problem: the construction of $S^{n}_{M^{-1}\mathbb{Z}}$ involves self-maps of $S^{n}$ of arbitrary degree, whereas the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ only has the trivial self-maps.
+A first suggestion is to start with the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ and use it to emulate the construction of $S^{n}_{M^{-1}\mathbb{Z}}$ in the $1$-reduced setting.
+This approach runs into a problem: the construction of $S^{n}_{M^{-1}\mathbb{Z}}$ involves self-maps of $S^{n}$ of arbitrary degree, whereas, by the following lemma, the simplicial $n$-sphere $S_{\mathrm{simp}}^{n}$ only has the trivial self-maps.
 Lemma 7.5.
@@
 Proof.
-Let $f\colon S_{\mathrm{simp}}^{n}\rightarrow S_{\mathrm{simp}}^{n}$ be a self-map of $S_{\mathrm{simp}}^{n}$.
-Let $x$ denote the unique 00-simplex of $S_{\mathrm{simp}}^{n}$ and let $\epsilon$ denote the non-degenerate $n$-simplex of $S_{\mathrm{simp}}^{n}$.
-There are two possibilities for the $n$-simplex $f_{n}(\epsilon)$ of $S_{\mathrm{simp}}^{n}$.
-  1. (1)
+Let $f$ be a self-map of $S_{\mathrm{simp}}^{n}$.
+Let $x$ denote the 0-simplex of $S_{\mathrm{simp}}^{n}$, and let $\epsilon$ denote the non-degenerate $n$-simplex of $S_{\mathrm{simp}}^{n}$.
 If $f_{n}(\epsilon)=\epsilon$, then $f$ is the identity map of $S_{\mathrm{simp}}^{n}$.
-  2. (2)
 If $f_{n}(\epsilon)$ is the iterated degeneracy of $x$, then $f$ is the constant self-map of $S_{\mathrm{simp}}^{n}$.
-Thus, we have to settle for an indirect route. We start with the CW inclusion
-$S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}.$
-We apply the singular complex functor to obtain the inclusion of simplicial sets
-Sing($\operatorname{Sing}(S^{n})\rightarrow\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}).$
-Then, we apply the $1$st Eilenberg subcomplex functor to obtain the inclusion of $1$-reduced simplicial sets
+∎
+Thus, we have to settle for an indirect route.
+We start with the CW inclusion $S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$, and we apply the singular complex functor to obtain the inclusion of simplicial sets $\operatorname{Sing}(S^{n})\rightarrow\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})$.
+Then, we take 11st Eilenberg subcomplexes to get the inclusion of $1$-reduced simplicial sets
 $E_{1}\operatorname{Sing}(S^{n})\rightarrow E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}).$
-To confirm that this map is the analog we are looking for, we study its geometric realization
-$|E_{1}\operatorname{Sing}(S^{n})|\rightarrow|E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|.$
-In particular, we study the commutative diagram of CW complexes
-${{|E_{1}\operatorname{Sing}(S^{n})|}}$${{|\operatorname{Sing}(S^{n})|}}$${{S^{n}}}$${{|E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|}}$|Sing(${{|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|}}$${{S^{n}_{M^{-1}\mathbb{Z}},}}$
-where the horizontal maps in the left-hand square are CW subcomplex inclusions and the horizontal maps in the right-hand square are natural maps from the geometric realization-singular complex adjunction.
-We show that all the horizontal maps are weak homotopy equivalences.
+To confirm that this map is our desired analog, we study the commutative diagram of CW complexes
+$|E_{1}\operatorname{Sing}(S^{n})|$$|\operatorname{Sing}(S^{n})|$$S^{n}$$|E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|$$|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|$$S^{n}_{M^{-1}\mathbb{Z}},$}}
+where the right-hand horizontal maps arise from the geometric realization-singular complex adjunction.
 Proposition 7.6.
 Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $n\geq 2$. In the commutative diagram
-${{|E_{1}\operatorname{Sing}(S^{n})|}}$${{|\operatorname{Sing}(S^{n})|}}$${{S^{n}}}$${{|E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|}}$|Sing(${{|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|}}$${{S^{n}_{M^{-1}\mathbb{Z}},}}$
+$|E_{1}\operatorname{Sing}(S^{n})|$$|\operatorname{Sing}(S^{n})|$$S^{n}$$|E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|$$|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|$$S^{n}_{M^{-1}\mathbb{Z}},$}}
 all the horizontal maps are weak homotopy equivalences.
@@
 By Corollary 4.4, the horizontal maps in the left-hand square are weak homotopy equivalences. By Theorem 4.1(2), the horizontal maps in the right-hand square are weak homotopy equivalences, as well. ∎
-Proposition 7.6 translates to the success of our approach. Thus, for a multiplicative subset $M$ of $\mathbb{Z}$, we left Bousfield localize ${\mathcal{SS}\mathrm{ets}}_{1}$ at the set of maps
-$T_{M^{-1}\mathbb{Z}}=\left\{E_{1}\operatorname{Sing}(S^{n})\rightarrow E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\right\}.$
-7.3. The family of model structures on $1$-reduced simplicial sets
+7.3. Description of the left Bousfield localization
 For a multiplicative subset $M$ of $\mathbb{Z}$, we describe the left Bousfield localization of the model category ${\mathcal{SS}\mathrm{ets}}_{1}$ at the set of maps
-$T_{M^{-1}\mathbb{Z}}=\left\{E_{1}\operatorname{Sing}(S^{n})\rightarrow E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\right\}.$
-In particular, we describe the $T_{M^{-1}\mathbb{Z}}$-local objects and the $T_{M^{-1}\mathbb{Z}}$-local equivalences in ${\mathcal{SS}\mathrm{ets}}_{1}$.
-To state our characterizations, we translate the definitions from Section 6.1 from the language of $1$-connected topological spaces to that of $1$-reduced simplicial sets using geometric realization.
+$B_{M^{-1}\mathbb{Z}}=\{E_{1}\operatorname{Sing}(S^{n})\rightarrow E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\}.$
+In particular, we characterize the $B_{M^{-1}\mathbb{Z}}$-local objects and the $B_{M^{-1}\mathbb{Z}}$-local equivalences in ${\mathcal{SS}\mathrm{ets}}_{1}$.
+To state our characterizations, we produce analogs of the notions from Section 6.1 for $1$-reduced simplicial sets.
 Definition 7.7.
@@
   2. (2)
-    A $1$-reduced simplicial set $K$ is $M$-local if $K$ is a Kan complex and $|K|$ is an $M$-local space.
+    A $1$-reduced simplicial set $K$ is $M$-local if $|K|$ is an $M$-local space.
   3. (3)
-    An $M$-localization of a $1$-reduced simplicial set $X$ is an $M$-local homotopy equivalence $f\colon X\rightarrow X_{M^{-1}\mathbb{Z}}$ with $M$-local target.
+    An $M$-localization of a $1$-reduced simplicial set $X$ is an $M$-local homotopy equivalence with $M$-local target $f\colon X\rightarrow X_{M^{-1}\mathbb{Z}}$.
 When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we speak of a rational homotopy equivalence, a rational simplicial set, and a rationalization.
-Example 7.8.
-By Proposition 7.6, every map in $T_{M^{-1}\mathbb{Z}}$ is an $M$-localization. Note that Corollary 3.19 says that the target of every map in $T_{M^{-1}\mathbb{Z}}$ is a Kan complex.
-The following proposition helps us transfer the rational homotopy theory of $1$-connected spaces from Section 6 to $1$-reduced simplicial sets.
-Proposition 7.9 ([Hir03, Proposition 1.1.11]).
-If $A$ and $X$ are pointed simplicial sets and $X$ is a Kan complex, then there is a weak homotopy equivalence
-$\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(A,X\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|A|,|X|\right)$
-that is natural in $A$ and in $X$.
-By a series of characterizations, we show that a $1$-reduced simplicial set is $T_{M^{-1}\mathbb{Z}}$-local if and only if it is $M$-local.
+The following proposition helps us transition from $1$-reduced simplicial sets to $1$-connected spaces.
+Proposition 7.8 ([26, Proposition 1.1.11]).
+If $A$ and $X$ are pointed simplicial sets and $X$ is a Kan complex, then there is a natural weak equivalence $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(A,X)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|A|,|X|)$.
+We show that a $1$-reduced simplicial set is $B_{M^{-1}\mathbb{Z}}$-local if and only if it is an $M$-local Kan complex.
 Proposition 7.9.
-Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $X$ be a $1$-reduced simplicial set. The following statements are equivalent.
+Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $X$ be a $1$-reduced Kan complex. Let $n\geq 2$. The following statements are equivalent.
   1. (1)
-    The $1$-reduced simplicial set $X$ is $T_{M^{-1}\mathbb{Z}}$-local.
+    The map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(E_{1}\operatorname{Sing}(S^{n}),X)$ is a weak equivalence.
   2. (2)
-    The $1$-reduced simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
-    $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(E_{1}\operatorname{Sing}(S^{n}),X\right)$
-    is a weak homotopy equivalence.
+    The map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(\operatorname{Sing}(S^{n}),X)$ is a weak equivalence.
   3. (3)
-    The $1$-reduced simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
-    $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}),X\right)$
-    is a weak homotopy equivalence.
+    The map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|\operatorname{Sing}(S^{n})|,|X|)$ is a weak equivalence.
   4. (4)
-    The $1$-reduced simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
-    $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n})|,|X|\right)$
-    is a weak homotopy equivalence.
+    The map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n}_{M^{-1}\mathbb{Z}},|X|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n},|X|)$ is a weak equivalence.
   5. (5)
-    The $1$-reduced simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
-    $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n}_{M^{-1}\mathbb{Z}},|X|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n},|X|)$
-    is a weak homotopy equivalence.
-  6. (6)
-    The $1$-reduced simplicial set $X$ is a Kan complex such that $|X|$ is an $M$-local space.
-  7. (7)
-    The $1$-reduced simplicial set $X$ is $M$-local.
-Hence, $X$ is $T_{M^{-1}\mathbb{Z}}$-local if and only if $X$ is $M$-local.
+    The space $|X|$ is $M$-local.
 Proof.
-Firstly, (1) and (2) are equivalent by the definition of a $T_{M^{-1}\mathbb{Z}}$-local object in ${\mathcal{SS}\mathrm{ets}}_{1}$ combined with the characterization of fibrant objects in ${\mathcal{SS}\mathrm{ets}}_{1}$ from Theorem 3.17.
-Then, (2) and (3) are equivalent by the commutative diagram
-${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}),X\right)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(E_{1}\operatorname{Sing}(S^{n}),X\right),}}$
-whose horizontal maps are weak homotopy equivalences by Proposition 7.6. Moreover, (3) and (4) are equivalent by the commutative diagram
-${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}\left(\operatorname{Sing}(S^{n}),X\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n})|,|X|\right),}}$
-whose horizontal maps are weak homotopy equivalences by Proposition 7.9. Furthermore, (4) and (5) are equivalent by the commutative diagram
-${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n}_{M^{-1}\mathbb{Z}},|X|)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n},|X|)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|\operatorname{Sing}(S^{n})|,|X|\right),}}$
-whose horizontal maps are weak homotopy equivalences by Proposition 7.6.
-Also, (5) and (6) are equivalent by Proposition 6.9.
-Lastly, (6) and (7) are equivalent by the definition of an $M$-local simplicial set.
+First, (1) and (2) are equivalent by Proposition 7.6.
+Then, (2) and (3) are equivalent by Proposition 7.8.
+Now, (3) and (4) are equivalent by Proposition 7.6.
+Lastly, (4) and (5) are equivalent by Proposition 6.9.
 ∎
-By a similar series of characterizations, we establish that a map of $1$-reduced simplicial sets is a $T_{M^{-1}\mathbb{Z}}$-local equivalence if and only if it is an $M$-local homotopy equivalence.
-To that end, we need two auxiliary lemmas.
-Lemma 7.11.
-Let $M$ be a multiplicative subset of $\mathbb{Z}$. If $Z$ is an $M$-local pointed space, then $E_{1}\operatorname{Sing}(Z)$ is an $M$-local simplicial set.
-Proof.
-By Corollary 3.19, $E_{1}\operatorname{Sing}(Z)$ is a Kan complex. Then, Corollary 4.5 and Theorem 4.1(2) provide the weak homotopy equivalences
-$|E_{1}\operatorname{Sing}(Z)|\rightarrow|\operatorname{Sing}(Z)|\rightarrow Z,$
-so $|E_{1}\operatorname{Sing}(Z)|$ is $M$-local. Overall, $E_{1}\operatorname{Sing}(Z)$ is $M$-local, as required. ∎
+By a similar series of characterizations, we establish that a map of $1$-reduced simplicial sets is a $B_{M^{-1}\mathbb{Z}}$-local equivalence if and only if it is an $M$-local homotopy equivalence.
+To that end, we need an auxiliary lemma.
 Lemma 7.10.
@@
   1. (1)
-    For every $M$-local simplicial set $K$, the map
-    $|f|^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,|K|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,|K|)$
-    is a weak homotopy equivalence.
+    For every $M$-local Kan complex $K$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|f|,|K|)$ is a weak equivalence.
   2. (2)
-    For every $M$-local pointed space $Z$, the map
-    $|f|^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,Z)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,Z)$
-    is a weak homotopy equivalence.
+    For every $M$-local pointed space $Z$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|f|,Z)$ is a weak equivalence.
 Proof.
-Suppose (1) holds. Let $Z$ be an $M$-local pointed space. Then, by Lemma 7.11, $E_{1}\operatorname{Sing}(Z)$ is an $M$-local simplicial set. There is a commutative diagram
-${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|Y|,|E_{1}\operatorname{Sing}(Z)|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|Y|,|\operatorname{Sing}(Z)|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,Z)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|X|,|E_{1}\operatorname{Sing}(Z)|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|X|,|\operatorname{Sing}(Z)|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,Z)}}$
-whose horizontal maps are weak homotopy equivalences by Corollary 4.5 and Theorem 4.1(2) and whose left-hand vertical map is a weak homotopy equivalence by (1).
-Thus, the right-hand vertical map is a weak homotopy equivalence, which proves (2).
-Conversely, suppose (2) holds. Let $K$ be an $M$-local simplicial set. Then, $|K|$ is an $M$-local pointed space, so
-$|f|^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|Y|,|K|\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(|X|,|K|\right)$
-is a weak homotopy equivalence, which proves (1). ∎
-We are now ready for the series of characterizations that describes the $T_{M^{-1}\mathbb{Z}}$-local equivalences.
+Suppose (1) holds.
+Let $Z$ be an $M$-local pointed space.
+By Corollary 3.18, $E_{1}\operatorname{Sing}(Z)$ is a Kan complex.
+Corollary 4.4 and Theorem 4.1(2) provide the weak homotopy equivalences $|E_{1}\operatorname{Sing}(Z)|\rightarrow|\operatorname{Sing}(Z)|\rightarrow Z$, so $E_{1}\operatorname{Sing}(Z)$ is an $M$-local Kan complex.
+Now, there is a commutative diagram
+$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,|E_{1}\operatorname{Sing}(Z)|)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,|\operatorname{Sing}(Z)|)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,Z)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,|E_{1}\operatorname{Sing}(Z)|)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,|\operatorname{Sing}(Z)|)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,Z)$}}
+whose horizontal maps are weak equivalences by Corollary 4.4 and Theorem 4.1(2) and whose left-hand vertical map is a weak equivalence by (1); thus, so is the right-hand vertical map, proving (2).
+Conversely, suppose (2) holds.
+Let $K$ be an $M$-local simplicial set.
+Then, $|K|$ is an $M$-local pointed space, which implies (1).
+∎
+Using our lemma, we show that a $B_{M^{-1}\mathbb{Z}}$-local equivalence is precisely an $M$-local homotopy equivalence.
 Proposition 7.11.
-Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets. The following statements are equivalent.
+Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $f$ be a map of $1$-reduced simplicial sets. The following statements are equivalent.
   1. (1)
-    The map $f$ is a $T_{M^{-1}\mathbb{Z}}$-local equivalence.
+    For every $B_{M^{-1}\mathbb{Z}}$-local simplicial set $K$, the map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(f,K)$ is a weak equivalence.
   2. (2)
-    For every $T_{M^{-1}\mathbb{Z}}$-local simplicial set $K$, the map
-    $f^{\ast}\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(X,K)$
-    is a weak homotopy equivalence.
+    For every $M$-local Kan complex $K$, the map $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(f,K)$ is a weak equivalence.
   3. (3)
-    For every $M$-local simplicial set $K$, the map
-    $f^{\ast}\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(X,K)$
-    is a weak homotopy equivalence.
+    For every $M$-local Kan complex $K$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|f|,|K|)$ is a weak equivalence.
   4. (4)
-    For every $M$-local simplicial set $K$, the map
-    $|f|^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,|K|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,|K|)$
-    is a weak homotopy equivalence.
-  5. (5)
-    For every $M$-local pointed space $Z$, the map
-    $|f|^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,Z)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,Z)$
-    is a weak homotopy equivalence.
+    For every $M$-local pointed space $Z$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|f|,Z)$ is a weak equivalence.
   5. (5)
     The map $|f|$ is an $M$-local homotopy equivalence.
-  7. (7)
-    The map $f$ is an $M$-local homotopy equivalence.
-Hence, $f$ is a $T_{M^{-1}\mathbb{Z}}$-local equivalence if and only $f$ is an $M$-local homotopy equivalence.
 Proof.
-Firstly, (1) and (2) are equivalent by the definition of a $T_{M^{-1}\mathbb{Z}}$-local equivalence.
+First, (1) and (2) are equivalent by Proposition 7.9.
 Then, (2) and (3) are equivalent by Proposition 7.8.
-Now, (3) and (4) are equivalent by the commutative diagram
-${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(Y,K)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,|K|)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(X,K)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,|K|),}}$$\scriptstyle{f^{\ast}}$$\scriptstyle{|f|^{\ast}}$
-whose horizontal maps are weak homotopy equivalences by Proposition 7.9.
-Moreover, (4) and (5) are equivalent by Lemma 7.12.
-Furthermore, (5) and (6) are equivalent by Proposition 6.12.
-Lastly, (6) and (7) are equivalent by the definition of an $M$-local homotopy equivalence.
+Now, (3) and (4) are equivalent by Lemma 7.10.
+Lastly, (4) and (5) are equivalent by Proposition 6.12.
 ∎
-Propositions 7.10 and 7.13 together describe the left Bousfield localization of ${\mathcal{SS}\mathrm{ets}}_{1}$ at the set $T_{M^{-1}\mathbb{Z}}$.
+Propositions 7.9 and 7.11 together describe the left Bousfield localization of ${\mathcal{SS}\mathrm{ets}}_{1}$ at the set $B_{M^{-1}\mathbb{Z}}$.
 Theorem 7.12.
@@
     a cofibration if and only if $f$ is a monomorphism.
-A $1$-reduced simplicial set $X$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $X$ is $M$-local.
+A $1$-reduced simplicial set $X$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $X$ is an $M$-local Kan complex.
 The model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ is combinatorial, left proper, and simplicial with the structure from Theorem 5.5.
-Theorem 7.14 recovers the family of model structures on $1$-reduced simplicial sets in [Qui69, Part II, Theorem 2.2], as well as the characterization of fibrant objects in [Qui69, Part II, Corollary 2.6].
-The most prominent model structure in this family is that corresponding to the case $M^{-1}\mathbb{Z}=\mathbb{Q}$, so we present it separately.
-Theorem 7.15.
-There is a model structure $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ on the category of $1$-reduced simplicial sets in which a map $f$ is
+Theorem 7.12 recovers the family of model structures on $1$-reduced simplicial sets in [33, Part II, Theorem 2.2], as well as the characterization of fibrant objects in [33, Part II, Corollary 2.6].
+The most prominent model structure in this family is that corresponding to the case $M^{-1}\mathbb{Z}=\mathbb{Q}$, which we present separately.
+Corollary 7.13.
+There is a model structure $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ on the category of $1$-reduced simplicial sets in which
   1. (1)
-    a weak equivalence if and only if $f$ is a rational homotopy equivalence; and
+    a map is a weak equivalence if and only if it is a rational homotopy equivalence; and
   2. (2)
-    a cofibration if and only if $f$ is a monomorphism.
-A $1$-reduced simplicial set $X$ is fibrant in $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $X$ is rational.
-The model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ is combinatorial, left proper, and simplicial with the structure from Theorem 5.15.
+    a $1$-reduced simplicial set is fibrant if and only if it is a rational Kan complex.
+Lastly, note that our approach establishes some additional model-categorical features of the model categories in Quillen’s family; namely, we know their simplicial structure, as well as a generating set of cofibrations.
 8. An approach using localization with respect to homology
 We give an alternative description of the family of model structures from Theorem 7.12.
-This description uses Bousfield localization with respect to homology, as developed in [Bou75], together with Theorem 3.15 on left-induced model structures.
-The idea is that, instead of first left-transferring the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ and then localizing, we can first localize the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ and then left-transfer it.
-As a pointed simplicial set need not be $1$-connected, we do not localize ${\mathcal{SS}\mathrm{ets}}_{\ast}$ with respect to homotopy, but rather with respect to homology.
-Because the two notions are in agreement for $1$-connected spaces, the left transfer to ${\mathcal{SS}\mathrm{ets}}_{1}$ yields the model structures from Theorem 7.14, and we extract our alternative description.
+This description uses Bousfield localization with respect to homology, as developed by Bousfield in [7], together with Theorem 3.15 on left-induced model structures.
+The idea is that, instead of first left-transferring the model structure on $\mathcal{SS}\mathrm{ets}$$_{\ast}$ and then localizing, we can first localize the model structure on $\mathcal{SS}\mathrm{ets}$$_{\ast}$ and then do a left transfer.
+We localize $\mathcal{SS}\mathrm{ets}$$_{\ast}$ with respect to homology.
+Because homotopy and homology are in agreement for $1$-connected spaces, the left transfer to ${\mathcal{SS}\mathrm{ets}}_{1}$ yields the family of model structures from Theorem 7.12.
 8.1. Bousfield localization with respect to homology
@@
 Let $M$ be a multiplicative subset of $\mathbb{Z}$.
-  1. (1)
-    A map of topological spaces $f\colon X\rightarrow Y$ is an $M$-local homology isomorphism if, for every $n\geq 0$, the induced map
-    $f_{\ast}\colon H_{n}(X;M^{-1}\mathbb{Z})\rightarrow H_{n}(Y;M^{-1}\mathbb{Z})$
-    is an isomorphism.
-  2. (2)
+A map of topological spaces $f\colon X\rightarrow Y$ is an $M$-local homology isomorphism if, for every $n\geq 0$, the map $H_{n}(f;M^{-1}\mathbb{Z})\colon H_{n}(X;M^{-1}\mathbb{Z})\rightarrow H_{n}(Y;M^{-1}\mathbb{Z})$ is an isomorphism.
 A simplicial map $g$ is an $M$-local homology isomorphism if $|g|$ is an $M$-local homology isomorphism of topological spaces.
@@
 Remark 8.2.
-As the abelian group $M^{-1}\mathbb{Z}$ is flat [Rot09, Theorem 4.80], the universal coefficient theorem for homology yields the natural isomorphism
-$H_{n}(X;M^{-1}\mathbb{Z})\cong H_{n}(X;\mathbb{Z})\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}.$
-Thus, $M$-local homology discards torsion from integral homology in the same fashion $M$-local homotopy discards torsion from higher homotopy groups.
-When $M^{-1}\mathbb{Z}=\mathbb{Q}$, all torsion is discarded in both cases.
+As the abelian group $M^{-1}\mathbb{Z}$ is flat [35, Theorem 4.80], the universal coefficient theorem for homology yields the natural isomorphism $H_{n}(X;M^{-1}\mathbb{Z})\cong H_{n}(X;\mathbb{Z})\otimes M^{-1}\mathbb{Z}$.
+Thus, $M$-local homology discards torsion from integral homology just as $M$-local homotopy discards torsion from the higher homotopy groups.
+The following example is invoked later in this paper.
 Example 8.3.
@@
 Also for later use, we need to recall the following theorem that relates integral homology and homotopy.
-Theorem 8.5 ([DK01, Corollary 6.69(1)]).
-Every weak homotopy equivalence is an integral homology isomorphism.
-We also need to recall another theorem of a similar nature for future reference.
-This theorem informs us that, under the hypothesis of $1$-connectedness, $M$-local homology agrees with $M$-local homotopy.
-Theorem 8.6 ([FHT01, Theorem 8.6]).
+Theorem 8.5 ([16, Corollary 6.69(1)]).
+A weak homotopy equivalence is an integral homology isomorphism.
+We also need to recall another theorem of a similar nature for future reference; this theorem informs us that, under the hypothesis of $1$-connectedness, local homology agrees with local homotopy.
+Theorem 8.6 ([22, Theorem 8.6]).
 Let $M$ be a multiplicative subset of $\mathbb{Z}$. A map of $1$-connected spaces $f$ is an $M$-local homology isomorphism if and only if $f$ is an $M$-local homotopy equivalence.
 For a multiplicative subset $M$ of $\mathbb{Z}$, Bousfield localization with respect to homology endows the category $\mathcal{SS}\mathrm{ets}$ of simplicial sets with a model structure whose weak equivalences are the $M$-local homology isomorphisms.
-Theorem 8.7 ([Bou75, Theorem 10.2]).
-Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}$ on the category of simplicial sets in which a map $f$ is
+We use the pointed version of this result, which holds by the same proof.
+Theorem 8.7 ([7, Theorem 10.2]).
+Let $M$ be a multiplicative subset of $\mathbb{Z}$.
+There exists a model structure $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ on the category of pointed simplicial sets in which a map $f$ is
   1. (1)
@@
 Remark 8.8.
-By [Bou97, §2.5], $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}$ is the left Bousfield localization of $\mathcal{SS}\mathrm{ets}$ at a single map, which allows us to infer that $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}$ inherits the properties of being combinatorial, left proper, and simplicial.
-When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we get a model structure whose weak equivalences are the rational homology isomorphisms.
-Theorem 8.9 ([Bou75, Theorem 10.2]).
-There is a model structure $\mathcal{L}_{H\mathbb{Q}}\mathcal{SS}\mathrm{ets}$ on the category of simplicial sets in which a map $f$ is
-  1. (1)
-    a weak equivalence if and only if $f$ is a rational homology isomorphism; and
-  2. (2)
-    a cofibration if and only if $f$ is a monomorphism.
-This model category is combinatorial, left proper, and simplicial with the structure from Example 5.2.
-We use the pointed version of Theorem 8.7, which holds by the same proof.
-Theorem 8.10.
-Let $M$ be a multiplicative subset of $\mathbb{Z}$.
-There is a model structure $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ on the category of pointed simplicial sets in which a map $f$ is
-  1. (1)
-    a weak equivalence if and only if $f$ is an $M$-local homology isomorphism; and
-  2. (2)
-    a cofibration if and only if $f$ is a monomorphism.
-This model category is combinatorial, left proper, and simplicial with the structure from Example 5.3.
-8.2. An alternative description of the family of model structures
+By [9, §2.5], $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ is the left Bousfield localization of $\mathcal{SS}\mathrm{ets}_{\ast}$ at a single map, which allows us to infer that $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ inherits the properties of being combinatorial, left proper, and simplicial.
+In particular, we have a model structure whose weak equivalences are the rational homology isomorphisms.
+8.2. An alternative description of the family of model structures on $1$-reduced simplicial sets
 For a multiplicative subset $M$ of $\mathbb{Z}$, we show that the model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ from Theorem 7.12 is the left transfer of the model category $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ from Theorem 8.7 along the adjunction
 $F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows\mathcal{SS}\mathrm{ets}_{\ast}\colon E_{1}$
-from Lemma 2.23(2). For this left transfer, we rely on Theorem 3.15, so we verify the condition in its hypothesis.
+from Proposition 2.19(2). Again, we use Theorem 3.15, so we verify the homotopical condition in its hypothesis.
 Lemma 8.9.
 Let $M$ be a multiplicative subset of $\mathbb{Z}$.
-Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets.
+Let $f$ be a map of $1$-reduced simplicial sets.
 If $f$ has the right lifting property with respect to every monomorphism of $1$-reduced simplicial sets, then $f$ is an $M$-local homology isomorphism.
@@
 Suppose $f$ has the right lifting property with respect to every monomorphism of $1$-reduced simplicial sets.
-By Lemma 3.16, we know that $f$ is a weak homotopy equivalence.
-Then, Theorem 8.5 informs us that $f$ is an integral homology isomorphism.
-Therefore, by Lemma 8.4, $f$ is an $M$-local homology isomorphism.
+By Lemma 3.16, we know that $|f|$ is a weak homotopy equivalence.
+Then, Theorem 8.5 informs us that $|f|$ is an integral homology isomorphism.
+Therefore, by Lemma 8.4, $|f|$ is an $M$-local homology isomorphism.
 ∎
-We invoke Theorem 3.15 to give an alternative approach to the family of model structures from Theorem 7.14.
-Proposition 8.12.
+We use Theorem 3.15 to give an alternative approach to the family of model structures from Theorem 7.12.
+Theorem 8.10.
 Let $M$ be a multiplicative subset of $\mathbb{Z}$.
-The model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ is the left transfer of the model category $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ along the adjunction
-$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}.$
+The model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ is the left transfer of the model category $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ along the adjunction $F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}$.
 Proof.
-Theorem 3.15 says that we can left transfer the model structure $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ along the adjunction
-$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}$
-to a model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$ in which a map $f$ is
-  1. (1)
-    a weak equivalence if and only if $f$ is an $M$-local homology isomorphism; and
-  2. (2)
-    a cofibration if and only if $f$ is a monomorphism.
-By Theorem 8.6, a map of $1$-reduced simplicial sets $f$ is an $M$-local homology isomorphism if and only if $f$ is an $M$-local homotopy equivalence.
-Thus, the left transfer gives us the model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$.
+Theorem 3.15 says that we can left transfer the model structure $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ along the adjunction to a model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$ whose weak equivalences are the $M$-local homology isomorphisms and whose cofibrations are the monomorphisms.
+By Theorem 8.6, a map of $1$-reduced simplicial sets is an $M$-local homology isomorphism if and only if it is an $M$-local homotopy equivalence, so the model structure is $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$.
 ∎
-9. Rational homotopy theory beyond $1$-connected spaces
+In particular, we obtain another modern description of the model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ from Corollary 7.13.
+9. Rational homotopy theory of non-simply connected spaces
 Let $M$ be a multiplicative subset of $\mathbb{Z}$.
-We study some elements of rational homotopy theory beyond $1$-connected spaces in order to extend Quillen’s family of model structures from $1$-reduced simplicial sets to all simplicial sets.
-In particular, we characterize $M$-local spaces in terms of function complexes, which prepares us for the left Bousfield localization in Section 10.
-We expect this characterization to be familiar to the rational homotopy theory community, but we could not find it in the literature, so we provide a self-contained proof.
-9.1. Localization of spaces
-We adopt the generalization of rational homotopy theory to all spaces from [GTHT00, §6]; in particular, we generalize Definition 6.1 to all spaces as follows.
-Definition 9.1.
+We study some elements of the rational homotopy theory of non-simply connected spaces in order to generalize Quillen’s family of model structures beyond $1$-reduced simplicial sets.
+In particular, we retain our description of $M$-local spaces in terms of function complexes beyond the $1$-connected setting, which allows us to describe our left Bousfield localization of all spaces in Section 10.
+9.1. Rationalization of spaces
+Firstly, Definition 6.1 generalizes to all spaces as follows.
+Definition 9.1 ([24, §6]).
 Let $M$ be a multiplicative subset of $\mathbb{Z}$.
   1. (1)
-    A continuous map $f\colon X\rightarrow Y$ is an $M$-local homotopy equivalence if, for every $n\leq 1$, the induced map
-    $f_{\ast}\colon\pi_{n}(X)\rightarrow\pi_{n}(Y)$
-    is an isomorphism and, for every $n\geq 2$, the induced map
-    $f_{\ast}\underset{\mathbb{Z}}{\otimes}\operatorname{id}_{M^{-1}\mathbb{Z}}\colon\pi_{n}(X)\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}\rightarrow\pi_{n}(Y)\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}$
-    is an isomorphism.
+    A map $f\colon X\rightarrow Y$ is an $M$-local homotopy equivalence if $f_{\ast}\colon\pi_{n}(X)\rightarrow\pi_{n}(Y)$ is an isomorphism for every $n\leq 1$ and $f_{\ast}\otimes\operatorname{id}_{M^{-1}\mathbb{Z}}\colon\pi_{n}(X)\otimes M^{-1}\mathbb{Z}\rightarrow\pi_{n}(Y)\otimes M^{-1}\mathbb{Z}$ is an isomorphism for every $n\geq 2$.
   2. (2)
@@
   3. (3)
-    An $M$-localization of a space $Y$ is an $M$-local homotopy equivalence $f\colon Y\rightarrow Y_{M^{-1}\mathbb{Z}}$ with $M$-local target.
+    An $M$-localization of a space $Y$ is an $M$-local homotopy equivalence with $M$-local target $f\colon Y\rightarrow Y_{M^{-1}\mathbb{Z}}$.
 When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we speak of a rational homotopy equivalence, a rational space, and a rationalization.
-Our goal of extending Quillen’s family of model structures to all simplicial sets precludes us from imposing any connectedness or nilpotency hypothesis on spaces.
-The absence of any connectedness hypothesis forces us to keep track of the basepoint across Definition 9.1 and approach disconnected spaces component-by-component.
-The absence of any nilpotency hypothesis deprives us of any sensible notion of tensoring the fundamental group with $\mathbb{Q}$, so we have to ask that a rational homotopy equivalence induce an isomorphism of fundamental groups.
+The absence of any nilpotency hypothesis deprives us of any sensible notion of tensoring the fundamental group with $\mathbb{Q}$, so we ask that a rational homotopy equivalence induce an isomorphism of fundamental groups.
 Example 9.2.
 Every weak homotopy equivalence is an $M$-local homotopy equivalence. Also, every $M$-local homotopy equivalence of $M$-local spaces is a weak homotopy equivalence.
+The following example ought to be contrasted with Example 8.3.
+Example 9.3.
+The unique map from $\mathbb{R}P^{2}$ to the point is not a rational homotopy equivalence; the nontriviality of $\pi_{1}(\mathbb{R}P^{2})\cong\mathbb{Z}/2\mathbb{Z}$ precludes it from inducing an isomorphism of fundamental groups.
 Example 9.4.
@@
 Example 9.4 begs the question of why our definition of an $M$-local space imposes no condition on the fundamental group.
-In particular, we can ask that, for every element $m$ of $M$, the $m$-th power map $g\mapsto g^{m}$ be a self-bijection of the fundamental group, even though it is generally not a homomorphism for non-abelian groups.
-The problem with imposing this additional condition is that, with this stricter definition, there are spaces that do not admit any $M$-localization.
-This striking result is proven by Casacuberta in [Cas93, Proposition 2.1].
-On the contrary, in our generalization of rational homotopy theory, every space can be localized, albeit not with the universal property that the localization of $1$-connected spaces enjoys.
-Theorem 9.4 ([GTHT00, §6]).
+In particular, we can ask that, for every element $m$ of $M$, the $m$-th power map $g\mapsto g^{m}$ be a self-bijection of the fundamental group, even though it need not be a homomorphism for non-abelian groups.
+The problem is that, with this stricter definition, there are spaces that do not admit any $M$-localization; this result is proven by Casacuberta in [13, Proposition 2.1].
+In our generalization of rational homotopy theory, spaces can be rationalized, albeit not with the universal property of the rationalization of $1$-connected spaces.
+Theorem 9.5 ([24, §6]).
 Let $M$ be a multiplicative subset of $\mathbb{Z}$. Every space admits an $M$-localization.
-The proof of the theorem is concise and it sheds light on what goes wrong with the universal property of localization, so we present it to provide some commentary afterwards.
+The proof of the theorem sheds light on what goes wrong with the universal property of rationalization.
 Proof.
-If $X$ is a connected space, then its $1$st Postnikov approximation yields the fibration
-$\widetilde{X}\rightarrow X\rightarrow K(\pi_{1}(X),1),$
-where $\widetilde{X}$ is the universal cover of $X$ and $K(\pi_{1}(X),1)$ is an Eilenberg-Mac Lane space.
-Since $\widetilde{X}$ is a $1$-connected space, we invoke fiberwise $M$-localization, in the sense of [Hir03, Theorem 6.1.3], to obtain a map of fibrations
+If $X$ is a connected space, its 11st Postnikov approximation yields the fibration $\widetilde{X}\rightarrow X\rightarrow K(\pi_{1}(X),1)$, where $\widetilde{X}$ is the universal cover of $X$ and $K(\pi_{1}(X),1)$ is an Eilenberg-Mac Lane space.
+Since $\widetilde{X}$ is a $1$-connected space, we invoke fiberwise $M$-localization, in the sense of [26, Theorem 6.1.3], to obtain a map of fibrations
 […]
-where the map on homotopy fibers $\widetilde{X}\rightarrow\widetilde{X}_{M^{-1}\mathbb{Z}}$ is an $M$-localization of the $1$-connected space $\widetilde{X}$ and, by the induced map of long exact sequences of homotopy groups, the map on total spaces $X\rightarrow X_{M^{-1}\mathbb{Z}}$ is the required $M$-localization of $X$.
-Applying this treatment to each component yields the claim for disconnected spaces.
+where the map on fibers $\widetilde{X}\rightarrow\widetilde{X}_{M^{-1}\mathbb{Z}}$ is an $M$-localization of the $1$-connected space $\widetilde{X}$ and, by the induced map of long exact sequences of homotopy groups, the map on total spaces $X\rightarrow X_{M^{-1}\mathbb{Z}}$ is the required $M$-localization of $X$.
+Applying this treatment to each path component yields the claim for disconnected spaces, as well.
 ∎
 The downside with fiberwise localization in our setting is that, because it applies to fibrations over a fixed base space $B$, its universal property states that it is initial among fibrations over $B$ with local fibers.
-Thus, our application of fiberwise localization in the proof of Theorem 9.4 only yields a universal property of $M$-localization over the Eilenberg-Mac Lane space $K(\pi_{1}(X),1)$, and we are neither aware of how to get rid of this restriction nor convinced that we should be able to do so in the first place.
-The absence of an analog of the universal property of localization for $1$-connected spaces has the following domino effect.
-The universal property of localization was instrumental for our characterization of local homotopy equivalences in terms of function complexes, which in turn gave us the description of the weak equivalences in the left Bousfield localization in Section 7.
-Thus, the left Bousfield localization in Section 10 is only a partial extension of that in Section 7 to all simplicial sets: we produce model structures whose fibrant objects are the local simplicial sets, but we do not know whether the weak equivalences are the local homotopy equivalences.
+Thus, our application of fiberwise localization only yields a universal property of $M$-localization over $K(\pi_{1}(X),1)$, and we are neither aware of how to get rid of this restriction nor convinced that we should even be able to do so.
+The universal property of rationalization was pivotal for our characterization of rational homotopy equivalences in terms of function complexes, which gave us the description of the weak equivalences in the left Bousfield localization in Section 7.
+Nevertheless, fiberwise localization still allows us to infer that, in our localization of all spaces in Section 10, the weak equivalences are the $M$-local homotopy equivalences in our generalized sense.
 9.2. The construction of local spheres
 Let $M$ be a multiplicative subset of $\mathbb{Z}$ and let $n\geq 2$.
-To salvage our function complex characterization of $M$-local spaces in the absence of a universal property for $M$-localizatrion, we need the details of the CW complex construction of the $M$-local $n$-sphere $S^{n}_{M^{-1}\mathbb{Z}}$ from [FHT01, Chapter 9(a)].
-Let $\mathbb{N}$ denote the set of positive integers. We enumerate
-$M\cap\mathbb{N}=\left\{k_{1},k_{2},k_{3},\dots\right\}.$
-We denote the $(n+1)$-disk by $D^{n+1}$. The $M$-local $n$-sphere $S^{n}_{M^{-1}\mathbb{Z}}$ is constructed as the pushout
+To salvage our function complex characterization of $M$-local spaces in the absence of a universal property for $M$-localizatrion, we need the details of the CW complex construction of the $M$-local $n$-sphere $S^{n}_{M^{-1}\mathbb{Z}}$ from [22, Chapter 9(a)].
+We enumerate $M\cap\mathbb{N}=\{m_{1},m_{2},m_{3},\dots\}.$ We denote the $(n+1)$-disk by $D^{n+1}$. Then, $S^{n}_{M^{-1}\mathbb{Z}}$ is the pushout
+$\coprod\limits_{j=1}^{\infty}(S^{n})_{j}$$\bigvee\limits_{i=0}^{\infty}(S^{n})_{i}$$\coprod\limits_{j=1}^{\infty}(D^{n+1})_{j}$$S^{n}_{M^{-1}\mathbb{Z}},$$\scriptstyle{h}$
+where, for $j\geq 1$, the $j$-th $(n+1)$-cell $(D^{n+1})_{j}$ is attached to $\bigvee\limits_{i=0}^{\infty}(S^{n})_{i}$ by an attaching map $h$ given by
+$S^{n}\xrightarrow{\textrm{pinch}}S^{n}\vee S^{n}\xrightarrow{\operatorname{id}_{S^{n}}\vee(-m_{j})}(S^{n})_{j-1}\vee(S^{n})_{j}\xrightarrow{\textrm{inclusion}}\bigvee\limits_{i=0}^{\infty}(S^{n})_{i}.$
+The map $(-m_{j})$ is a self-map of $S^{n}$ that has degree $(-m_{j})$. In other words, the $1$-connected CW complex
 […]
-where, for $j\geq 1$, the $j$-th $(n+1)$-cell $(D^{n+1})_{j}$ is attached to $\bigvee\limits_{i=0}^{\infty}(S^{n})_{i}$ by an attaching map $h$ given by
-[…]
-The map
-$(-k_{j})\colon S^{n}\rightarrow(S^{n})_{j}$
-is a self-map of the $n$-sphere $S^{n}$ that has degree $(-k_{j})$. In other words, the $1$-connected CW complex
-[…]
-is the mapping telescope of the sequence of maps
-$S^{n}\xrightarrow{k_{1}}S^{n}\xrightarrow{k_{2}}S^{n}\xrightarrow{k_{3}}\cdots.$
-9.3. A characterization of local spaces
-Let $M$ be a multiplicative subset of $\mathbb{Z}$. We use the construction of $S^{n}_{M^{-1}\mathbb{Z}}$ from Section 9.2, including our notation, to prove the following key lemma.
-Lemma 9.5.
+is the mapping telescope of the sequence of maps $S^{n}\xrightarrow{m_{1}}S^{n}\xrightarrow{m_{2}}S^{n}\xrightarrow{m_{3}}\cdots$.
+9.3. A characterization of rational spaces
+We use the construction of $S^{n}_{M^{-1}\mathbb{Z}}$ to prove a key result.
+Proposition 9.6.
 Let $M$ be a multiplicative subset of $\mathbb{Z}$ and let $n\geq 2$.
@@
 Proof.
-Given a map $f\colon S^{n}\rightarrow Y$, we know that $[f]$ is an element of the $M^{-1}\mathbb{Z}$-module $\pi_{n}(Y)$. Thus, there exists a unique element $[u]$ of $\pi_{n}(Y)$ such that
-$[f]=k_{1}\cdot[u],$
-and the representative $u\colon S^{n}\rightarrow Y$ of $[u]$ is unique up to homotopy. Using the notation in our construction of
-[…]
-we extend $f$ from $S^{n}=(S^{n})_{0}$ to $(S^{n})_{0}\vee(S^{n})_{1}$ uniquely up to homotopy by setting
-$g|_{(S^{n})_{1}}=u.$
-As
-$[f]-\left(k_{1}\cdot[u]\right)=0\in\pi_{n}(Y),$
-we can further extend $f$ to
+Given a map $f\colon S^{n}\rightarrow Y$, we know that $[f]$ is an element of the $M^{-1}\mathbb{Z}$-module $\pi_{n}(Y)$.
+Thus, there exists a unique element $[u]$ of $\pi_{n}(Y)$ such that $f]=m_{1}\cdot[u]$, and the representative $u\colon S^{n}\rightarrow Y$ of $[u]$ is unique up to homotopy.
+Using the notation in our construction of $S^{n}_{M^{-1}\mathbb{Z}}$, we extend $f$ from $S^{n}=(S^{n})_{0}$ to $(S^{n})_{0}\vee(S^{n})_{1}$ uniquely up to homotopy by setting $g|_{(S^{n})_{1}}=u$.
+As $f]-(m_{1}\cdot[u])=0$, we can further extend $f$ to
 […]
 uniquely up to homotopy. Continuing further down the telescope in this fashion yields the claim. ∎
-In particular, we need the following consequence of Lemma 9.5 to give a characterization of $M$-local spaces.
+Proposition 9.6 salvages our characterization of $M$-local spaces in terms of function complexes.
 Corollary 9.7.
-Let $M$ be a multiplicative subset of $\mathbb{Z}$ and let $n\geq 2$.
-Let $Y$ be an $M$-local space.
-Then, every map $f\colon S^{n}\rightarrow Y$ extends uniquely up to homotopy to a map $g\colon S^{n}_{M^{-1}\mathbb{Z}}\rightarrow Y$ such that the diagram
-[…]
-commutes.
-We salvage our characterization of $M$-local spaces for the left Bousfield localization in the next section.
-Proposition 9.7.
-Let $M$ be a multiplicative subset of $\mathbb{Z}$. A space $Y$ is $M$-local if and only if, for every $n\geq 2$, the map
-$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n},Y\right)$
-induced by the inclusion $S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$ is a weak homotopy equivalence.
+Let $M$ be a multiplicative subset of $\mathbb{Z}$.
+A space $Y$ is $M$-local if and only if, for every $n\geq 2$, the restriction map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},Y)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},Y)$ is a weak equivalence.
 Proof.
-If $Y$ is $M$-local, then Corollary 9.6 implies that, for every $n\geq 2$, the map
-$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n},Y\right)$
-is a weak homotopy equivalence. The converse implication is established as in the proof of Proposition 6.8. ∎
-10. The family of model structures on simplicial sets
+If $Y$ is $M$-local, then Proposition 9.6 implies that the restriction map is a weak equivalence. The converse implication is established as in the proof of Proposition 6.8. ∎
+10. The family of model structures on spaces
 Let $M$ be a multiplicative subset of $\mathbb{Z}$.
-We left Bousfield localize the model category ${\mathcal{SS}\mathrm{ets}}$ of simplicial sets to obtain a model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ whose fibrant objects are the $M$-local ones.
-In particular, we produce a model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ whose fibrant objects are the rational simplicial sets.
-Thus, we partially extend Quillen’s family of model structures from $1$-reduced simplicial sets to all simplicial sets.
-Our family of model structures is meant to axiomatize the generalization of rational homotopy theory that we studied in Section 9.
-10.1. The set for the left Bousfield localization
-Let $M$ be a multiplicative subset of $\mathbb{Z}$. We left Bousfield localize ${\mathcal{SS}\mathrm{ets}}$ at the set of maps
-$T_{M^{-1}\mathbb{Z}}=\left\{E_{1}\operatorname{Sing}(S^{n})\rightarrow E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\right\}$
-from Section 7.2.
-We observe that, by Proposition 7.6, the application of the $1$st Eilenberg subcomplex functor $E_{1}$ is superfluous for this localization, for we are no longer working within the confines of $1$-reduced simplicial sets.
-In other words, left Bousfield localizing ${\mathcal{SS}\mathrm{ets}}$ at the set of maps
-$B_{M^{-1}\mathbb{Z}}=\left\{\operatorname{Sing}(S^{n})\rightarrow\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\right\}$
-yields the same model category, so we equivalently localize ${\mathcal{SS}\mathrm{ets}}$ at the set of maps $B_{M^{-1}\mathbb{Z}}$ instead.
+We left Bousfield localize the model category $\mathcal{SS}\mathrm{ets}$} of simplicial sets to obtain a model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ in which the weak equivalences are the $M$-local homotopy equivalences in our extended sense and the fibrant objects are the $M$-local Kan complexes.
+In particular, we produce a model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ whose weak equivalences are the rational homotopy equivalences and whose fibrant objects are the rational Kan complexes.
+Thus, we generalize Quillen’s family of model structures to all spaces.
+Lastly, we show that our localization of spaces at rational homotopy does not coincide with Bousfield’s localization at rational homology as rational homotopy and rational homology do not agree beyond the $1$-connected setting.
 10.1. The family of model structures on simplicial sets
-For a multiplicative subset $M$ of $\mathbb{Z}$, we describe the left Bousfield localization of the model category ${\mathcal{SS}\mathrm{ets}}$ at the set of maps
-$B_{M^{-1}\mathbb{Z}}=\left\{\operatorname{Sing}(S^{n})\rightarrow\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\right\}.$
-In particular, we characterize the $B_{M^{-1}\mathbb{Z}}$-local objects in ${\mathcal{SS}\mathrm{ets}}$ in terms of the following definition.
+Proposition 9.6 allows us to describe the left Bousfield localization of $\mathcal{SS}\mathrm{ets}$} at the set of maps
+$T_{M^{-1}\mathbb{Z}}=\{\operatorname{Sing}(S^{n})\rightarrow\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\}.$
+To do so, we first formulate the analog of Definition 9.1 for simplicial sets.
 Definition 10.1.
 Let $M$ be a multiplicative subset of $\mathbb{Z}$.
-A simplicial set $K$ is $M$-local if $K$ is a Kan complex and $|K|$ is an $M$-local space.
-When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we speak of a rational simplicial set.
-Example 10.2.
-By Proposition 7.6, the target of every map in $B_{M^{-1}\mathbb{Z}}$ is an $M$-local simplicial set.
-Note that Example 3.2 says that the target of every map in $B_{M^{-1}\mathbb{Z}}$ is a Kan complex.
-The following proposition helps us transfer the rational homotopy theory from Section 9 to simplicial sets.
-Proposition 10.3 ([Hir03, Proposition 1.1.11]).
-If $A$ and $X$ are simplicial sets and $X$ is a Kan complex, then there is a weak homotopy equivalence
-$\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}\left(A,X\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|A|,|X|\right)$
-that is natural in $A$ and in $X$.
-By a series of characterizations, we show that a simplicial set is $B_{M^{-1}\mathbb{Z}}$-local if and only if it is $M$-local.
+  1. (1)
+    A simplicial map $f$ is an $M$-local homotopy equivalence if $|f|$ is an $M$-local homotopy equivalence.
+  2. (2)
+    A simplicial set $K$ is $M$-local if $|K|$ is an $M$-local space.
+  3. (3)
+    An $M$-localization of $X$ is an $M$-local homotopy equivalence with $M$-local target $X\rightarrow X_{M^{-1}\mathbb{Z}}$.
+When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we speak of a rational homotopy equivalence, a rational simplicial set, and a rationalization.
+Having salvaged our characterization of $M$-local spaces in terms of function complexes, we can replicate our proof of Proposition 7.9 to show that a simplicial set is $T_{M^{-1}\mathbb{Z}}$-local if and only if it is an $M$-local Kan complex.
+It remains to show that a $T_{M^{-1}\mathbb{Z}}$-local equivalence is precisely an $M$-local homotopy equivalence.
+To that end, we invoke a general result on fiberwise localizations for the fiberwise $M$-localization from Proposition 9.5.
+Lemma 10.2.
+[26, Theorem 6.1.3(3)] Let $M$ be a multiplicative subset of $\mathbb{Z}$. Every fiberwise $M$-localization map is a $T_{M^{-1}\mathbb{Z}}$-local equivalence.
+Lemma 10.2 allows us to characterize the $T_{M^{-1}\mathbb{Z}}$-local equivalences in our desired fashion.
 Proposition 10.3.
-Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $X$ be a simplicial set. The following statements are equivalent.
+Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $f\colon X\rightarrow Y$ be a simplicial map. The following statements are equivalent.
   1. (1)
-    The simplicial set $X$ is $B_{M^{-1}\mathbb{Z}}$-local.
+    The map $f$ is a $T_{M^{-1}\mathbb{Z}}$-local equivalence.
   2. (2)
-    The simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
-    $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}\left(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}\left(\operatorname{Sing}(S^{n}),X\right)$
-    is a weak homotopy equivalence.
+    The induced map on fiberwise $M$-localizations $f_{M^{-1}\mathbb{Z}}$ is a $T_{M^{-1}\mathbb{Z}}$-local equivalence.
   3. (3)
-    The simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
-    $\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n})|,|X|\right)$
-    is a weak homotopy equivalence.
+    The induced map on fiberwise $M$-localizations $f_{M^{-1}\mathbb{Z}}$ is a weak equivalence.
   4. (4)
-    The simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
-    $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},|X|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},|X|)$
-    is a weak homotopy equivalence.
+    The induced map on fiberwise $M$-localizations $f_{M^{-1}\mathbb{Z}}$ is an $M$-local homotopy equivalence.
   5. (5)
-    The simplicial set $X$ is a Kan complex such that $|X|$ is an $M$-local space.
+    The map $f$ is an $M$-local homotopy equivalence.
+Proof.
+Fiberwise $M$-localization induces the commutative square
+[…]
+Firstly, (1) and (2) are equivalent because the vertical maps are $T_{M^{-1}\mathbb{Z}}$-local equivalences by Lemma 10.2.
+Then, (2) and (3) are equivalent by Remark 7.2.
+Now, (3) and (4) are equivalent by Example 9.2.
+Lastly, (4) and (5) are equivalent because the vertical maps are $M$-local homotopy equivalences by Proposition 9.5.
+We arrive at our promised generalization of Quillen’s family of model structures to all spaces.
+Theorem 10.4.
+Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ on the category of simplicial sets in which a map $f$ is
   1. (1)
-    The simplicial set $X$ is $M$-local.
-Hence, $X$ is $B_{M^{-1}\mathbb{Z}}$-local if and only if $X$ is $M$-local.
-Proof.
-Firstly, (1) and (2) are equivalent by the definition of a $B_{M^{-1}\mathbb{Z}}$-local object in ${\mathcal{SS}\mathrm{ets}}$. Then, (2) and (3) are equivalent by the commutative diagram
-${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}\left(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}\left(\operatorname{Sing}(S^{n}),X\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n})|,|X|\right),}}$
-whose horizontal maps are weak homotopy equivalences by Proposition 10.3. Now, (3) and (4) are equivalent by the commutative diagram
-${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},|X|)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},|X|)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n})|,|X|\right),}}$
-whose horizontal maps are weak homotopy equivalences by Proposition 7.6.
-Also, (4) and (5) are equivalent by Proposition 9.7.
-Lastly, (5) and (6) are equivalent by the definition of an $M$-local simplicial set.
-∎
-Proposition 10.4 helps us describe the left Bousfield localization of ${\mathcal{SS}\mathrm{ets}}$ at the set of maps $B_{M^{-1}\mathbb{Z}}$.
-Theorem 10.5.
-Let $M$ be a multiplicative subset of $\mathbb{Z}$.
-There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ on the category ${\mathcal{SS}\mathrm{ets}}$ of simplicial sets in which a map $f\colon X\rightarrow Y$ is
-  1. (1)
-    a weak equivalence if and only if, for every $M$-local simplicial set $K$, the map
-    $f^{\ast}\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(X,K)$
-    is a weak homotopy equivalence; and
+    a weak equivalence if and only if $f$ is an $M$-local homotopy equivalence; and
   2. (2)
     a cofibration if and only if $f$ is a monomorphism.
-A simplicial set $X$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ if and only if $X$ is $M$-local.
+A simplicial set $X$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ if and only if $X$ is an $M$-local Kan complex.
 The model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ is combinatorial, left proper, and simplicial with the structure from Example 5.2.
-Theorem 10.5 partially extends Quillen’s family of model structures from $1$-reduced simplicial sets to all simplicial sets.
-A distinguished model structure in this family is that corresponding to the case $M^{-1}\mathbb{Z}=\mathbb{Q}$.
-Theorem 10.6.
-There is a model structure $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ on the category ${\mathcal{SS}\mathrm{ets}}$ of simplicial sets in which a map $f\colon X\rightarrow Y$ is
+Again, a distinguished model structure in this family is that corresponding to the rational case $M^{-1}\mathbb{Z}=\mathbb{Q}$.
+Corollary 10.5.
+There is a model structure $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ on the category of simplicial sets in which
   1. (1)
-    a weak equivalence if and only if, for every rational simplicial set $K$, the map
-    $f^{\ast}\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(X,K)$
-    is a weak homotopy equivalence; and
+    a map is a weak equivalence if and only if it is a rational homotopy equivalence; and
   2. (2)
-    a cofibration if and only if $f$ is a monomorphism.
-A simplicial set $X$ is fibrant in $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ if and only if $X$ is rational.
-The model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ is combinatorial, left proper, and simplicial with the structure from Example 5.2.
-11. The family of model structures on topological spaces
-Quillen’s family of model structures on the category of $1$-reduced simplicial sets finds no counterpart for the category of $1$-connected spaces, for the latter category cannot be endowed with any model structure because it does not have all finite limits and colimits; this failure necessitates the use of $1$-reduced simplicial sets in Quillen’s model-categorical approach of rational homotopy theory.
-To see that the category of $1$-connected spaces does not have all finite limits, consider the pullback square
-[…]
-where $f\colon\mathbb{R}^{2}\rightarrow\mathbb{R}$ is given by the formula $f(x,y)=x^{2}+y^{2}$, so $f^{-1}(1)=S^{1}$ is the unit circle in the plane.
-In this pullback square, we start with the contractible spaces $|\Delta[0]|$, $\mathbb{R}^{2}$, and $\mathbb{R}$, and we obtain $S^{1}$, which is not a $1$-connected space because $\pi_{1}(S^{1})\cong\mathbb{Z}$.
-Thus, we see that connectedness conditions on spaces do not interact well with limits and colimits, in contrast to skeletal conditions on simplicial sets.
-However, now that we have a family of model structures on the category of simplicial sets, we continue our work by presenting its analog for the category $\mathcal{T}\mathrm{op}$ of topological spaces.
-This analog is produced by left Bousfield localizing the model structure on $\mathcal{T}\mathrm{op}$, which we recall in this section.
-Moreover, because we first characterized local topological spaces and then transferred this characterization to simplicial sets in this section, there is no additional work that we need to do in order to obtain a family of model structures on topological spaces.
+    a simplicial set is fibrant if and only if it is a rational Kan complex.
+Thus, our modern approach to Quillen’s model structures on $1$-reduced simplicial sets generalizes to all spaces and yields a model category that encodes the rational homotopy theory of non-simply connected spaces.
 10.2. The model category of topological spaces
-We recall the model structure on $\mathcal{T}\mathrm{op}$ from [DS95, §8]. To describe it, we first need to introduce the following notions.
+Quillen’s family of model structures on the category of $1$-reduced simplicial sets finds no counterpart for the category of $1$-connected spaces, for the latter category does not have all finite limits and colimits.
+For instance, consider the pullback
+$S^{1}$$\mathbb{R}^{2}$$|\Delta[0]|$$\mathbb{R}$$\scriptstyle{f}$$\scriptstyle{1}$
+where $f$ is given by the formula $f(x,y)=x^{2}+y^{2}$, so $f^{-1}(1)=S^{1}$ is the unit circle in the plane.
+In this pullback, we start with the contractible spaces $|\Delta[0]|$, $\mathbb{R}^{2}$, and $\mathbb{R}$ and get $S^{1}$, which has fundamental group $\pi_{1}(S^{1})\cong\mathbb{Z}$.
+However, now that we have a family of model structures on simplicial sets, we can produce its analog for the category $\mathcal{T}\mathrm{op}$ of topological spaces by left Bousfield localizing the model structure on $\mathcal{T}\mathrm{op}$ at our usual set.
+We first recall the model structure on $\mathcal{T}\mathrm{op}$ from [18, §8]; the description is in terms of the following definition.
 Definition 10.6.
   1. (1)
-    Given a sequence of subspace inclusions
-    $X_{0}\rightarrow X_{1}\rightarrow\cdots$
-    such that, for every $n\geq 0$, the pair $(X_{n+1},X_{n})$ is a relative CW complex, we say that the map
-    $X_{0}\rightarrow\operatorname{colim}(X_{n})$
-    is a generalized relative CW inclusion.
+    Given a sequence of inclusions $X_{0}\rightarrow X_{1}\rightarrow\cdots$ such that the pairs $(X_{n+1},X_{n})$ are relative CW complexes, we say that $X_{0}\rightarrow\operatorname{colim}(X_{n})$ is a generalized relative CW inclusion.
   2. (2)
-    A Serre fibration is a continuous map $p\colon X\rightarrow Y$ that has the right lifting property with respect to the class of inclusions
-    $\left\{A\times\{0\}\rightarrow A\times[0,1]\mid A\textrm{ is a CW complex}\right\}.$
-    In other words, a Serre fibration is a continuous map that has the homotopy lifting property with respect to every CW complex.
-Remark 11.2.
-A continuous map is often said to be a fibration if it has the homotopy lifting property with respect to every space.
-This notion is stronger than that of a Serre fibration, and it is sometimes called a Hurewicz fibration to emphasize this distinction.
-Every Hurewicz fibration is a Serre fibration; the converse implication finds a pathological counterexample in [Bro66, §1], but it holds for maps of CW complexes [Cau92].
-Example 11.3.
-Every space $X$ is Serre fibrant: the unique map from $X$ to the point is a Serre fibration.
-We are now ready to describe the model category of topological spaces.
-Theorem 11.4 ([DS95, Propositions 8.3 and 8.9]).
+    A Serre fibration is a map that has the right lifting property with respect to the class of inclusions $\{A\times\{0\}\rightarrow A\times[0,1]\mid A\textrm{ is a CW complex}\}$.
+In other words, a Serre fibration is a map with the homotopy lifting property against all CW complexes.
+Theorem 10.7 ([18, Propositions 8.3 and 8.9]).
 The category $\mathcal{T}\mathrm{op}$} of topological spaces has a model structure in which a map $f$ is
@@
     a fibration if and only if $f$ is a Serre fibration.
-The model category of topological spaces is simplicial with the following structure.
-Proposition 11.5 ([Hir03, Example 9.1.15]).
+The model category of spaces is also cofibrantly generated and simplicial with the following structure.
+Proposition 10.8 ([26, Example 11.1.8]).
+The model category $\mathcal{T}\mathrm{op}$} is cofibrantly generated.
+  1. (1)
+    A generating set of cofibrations is $|I|=\{|i_{n}|\colon|\partial\Delta[n]|\rightarrow|\Delta[n]|\mid n\geq 0\}$.
+  2. (2)
+    A generating set of acyclic cofibrations is $|J|=\{|j_{n,k}|\colon|\Lambda[n,k]|\rightarrow|\Delta[n]|\mid n\geq 1\textrm{ and }0\leq k\leq n\}$.
+Proposition 10.9 ([26, Example 9.1.15]).
 Given two spaces $X$ and $Y$ and a simplicial set $Z$, we define
   1. (1)
-    the function complex $\operatorname{Map}_{{\mathcal{T}\mathrm{op}}}\left(X,Y\right)$ to be the simplicial set with $n$-simplices $\operatorname{Hom}_{{\mathcal{T}\mathrm{op}}}\left(X\times|\Delta[n]|,Y\right)$;
+    the function complex $\operatorname{Map}_{{\mathcal{T}\mathrm{op}}}(X,Y)$ to be the simplicial set with $n$-simplices $\operatorname{Hom}_{{\mathcal{T}\mathrm{op}}}(X\times|\Delta[n]|,Y)$;
   2. (2)
@@
 With these definitions, $\mathcal{T}\mathrm{op}$} has the structure of a simplicial model category.
-The model category of topological spaces is also cofibrantly generated; its generating cofibrations arise from geometrically realizing those for the model category of simplicial sets.
-Proposition 11.6 ([Hir03, Example 11.1.8]).
-The model category ${\mathcal{T}\mathrm{op}}$ is cofibrantly generated.
+The model category of topological spaces differs from that of simplicial sets in three noteworthy respects.
   1. (1)
-    A generating set of cofibrations for ${\mathcal{T}\mathrm{op}}$ is
-    $|I|=\left\{|i_{n}|\colon|\partial\Delta[n]|\rightarrow|\Delta[n]|\mid n\geq 0\right\}.$
+    It is not combinatorial because the category $\mathcal{T}\mathrm{op}$} is not locally presentable [1, Example 1.18(5)].
   2. (2)
-    A generating set of acyclic cofibrations for ${\mathcal{T}\mathrm{op}}$ is
-    $|J|=\left\{|j_{n,k}|\colon|\Lambda[n,k]|\rightarrow|\Delta[n]|\mid n\geq 1\textrm{ and }0\leq k\leq n\right\}.$
-The model category of topological spaces differs from that of simplicial sets in two noteworthy respects.
+    It is left proper [18, Theorem 13.1.10], but not by Proposition 3.12, for not every space is cofibrant.
   3. (3)
-    It is left proper [DS95, Theorem 13.1.10], but not by Proposition 3.12, for not every space is cofibrant.
+    Every topological space is Serre fibrant, whereas not every simplicial set is a Kan complex.
+Nevertheless, the geometric realization-singular complex adjunction is a Quillen equivalence [32, Chapter I.4, Example 2], which renders simplicial sets a model for spaces.
+Still, the first difference raises the question of whether $\mathcal{T}\mathrm{op}$} can be localized.
+We note that $\mathcal{T}\mathrm{op}$} shares with $\mathcal{SS}\mathrm{ets}$} the feature of being a cellular model category [26, Proposition 12.1.4].
+This property is a different way of strengthening cofibrant generation, and it lets us localize $\mathcal{T}\mathrm{op}$}, for Theorem 7.3 holds if we replace “combinatorial” with “cellular” [26, Theorem 4.1.1].
+10.3. The family of model structures on topological spaces
+Corollary 9.7 and Proposition 10.3 describe the left Bousfield localization of $\mathcal{T}\mathrm{op}$ at the set of maps $\{S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}\mid n\geq 2\}$ with no additional work needed.
+Theorem 10.10.
+Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{T}\mathrm{op}}$ on the category of topological spaces in which a map $f$ is
   1. (1)
-    It is not combinatorial because the category ${\mathcal{T}\mathrm{op}}$ is not locally presentable [AR94, Example 1.18(5)].
-The latter difference raises the question of whether ${\mathcal{T}\mathrm{op}}$ can be left Bousfield localized.
-To that end, we remark that ${\mathcal{T}\mathrm{op}}$ shares with ${\mathcal{SS}\mathrm{ets}}$ the feature of being a cellular model category [Hir03, Proposition 12.1.4].
-This property is a different way of strengthening the property of being a cofibrantly generated model category, and it allows us to left Bousfield localize ${\mathcal{T}\mathrm{op}}$, for Theorem 7.3 holds true if we replace “combinatorial” with “cellular” throughout [Hir03, Theorem 4.1.1].
-11.2. The family of model structures on topological spaces
-Our work in Section 9 yields the promised family of model structures on topological spaces.
-Theorem 11.7.
-Let $M$ be a multiplicative subset of $\mathbb{Z}$.
-There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{T}\mathrm{op}}$ on the category ${\mathcal{T}\mathrm{op}}$ of topological spaces in which a map $f\colon X\rightarrow Y$ is
-  1. (1)
-    a weak equivalence if and only if, for every $M$-local space $K$, the map
-    $f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,K)$
-    is a weak homotopy equivalence; and
+    a weak equivalence if and only if $f$ is an $M$-local homotopy equivalence; and
   2. (2)
@@
 The model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{T}\mathrm{op}}$ is cellular, left proper, and simplicial with the structure from Proposition 10.9.
-Proof.
-Proposition 9.7 informs us that $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{T}\mathrm{op}}$ is the left Bousfield localization of ${\mathcal{T}\mathrm{op}}$ at the set of maps
-$\left\{S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}\mid n\geq 2\right\}.$
+In particular, we obtain a model category whose weak equivalences are the rational homotopy equivalences and whose fibrant objects are the rational spaces.
 Remark 10.11.
 By Proposition 7.6, $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{T}\mathrm{op}}$ is also the left Bousfield localization of $\mathcal{T}\mathrm{op}$} at the set of maps
-$|B_{M^{-1}\mathbb{Z}}|=\left\{|\operatorname{Sing}(S^{n})|\rightarrow|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|\mid n\geq 2\right\}.$
-Thus, by [Hir03, Theorem 3.3.20], the Quillen equivalence
-$|-|\colon\mathcal{SS}\mathrm{ets}\rightleftarrows\mathcal{T}\mathrm{op}\colon\operatorname{Sing}$
-induces a Quillen equivalence
+$|T_{M^{-1}\mathbb{Z}}|=\{|\operatorname{Sing}(S^{n})|\rightarrow|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|\mid n\geq 2\}.$
+Thus, by [26, Theorem 3.3.20], the Quillen equivalence $|-|\colon\mathcal{SS}\mathrm{ets}\rightleftarrows\mathcal{T}\mathrm{op}\colon\operatorname{Sing}$ induces a Quillen equivalence
 $|-|\colon\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}\rightleftarrows\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{T}\mathrm{op}\colon\operatorname{Sing}.$
-This Quillen equivalence captures the fact that the model categories $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}$ and $\mathcal{L}_{M^{-1}\mathbb{Z}}\mathcal{T}\mathrm{op}$ both describe the $M$-local homotopy theory beyond $1$-connected spaces from Section 9.
-Lastly, we distinguish the model category corresponding to the case $M^{-1}\mathbb{Z}=\mathbb{Q}$.
-Theorem 11.9.
-There is a model structure $\mathcal{L}_{\mathbb{Q}}{\mathcal{T}\mathrm{op}}$ on the category ${\mathcal{T}\mathrm{op}}$ of topological spaces in which a map $f\colon X\rightarrow Y$ is
-  1. (1)
-    a weak equivalence if and only if, for every rational space $K$, the map
-    $f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,K)$
-    is a weak homotopy equivalence; and
-  2. (2)
-    a cofibration if and only if $f$ is a retract of a generalized relative CW inclusion.
-A space $X$ is fibrant in $\mathcal{L}_{\mathbb{Q}}{\mathcal{T}\mathrm{op}}$ if and only if $X$ is rational.
-The model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{T}\mathrm{op}}$ is cellular, left proper, and simplicial with the structure from Proposition 11.5.
-12. Rational homology versus rational homotopy
-We conclude this paper by contrasting Bousfield’s localization of spaces with respect to rational homology to our localization of spaces with respect to rational homotopy.
-Even though rational homology and rational homotopy agree for $1$-connected spaces, this agreement does not extend to all spaces.
-Therefore, we verify that the two localized model structures on spaces are not equal.
-12.1. Localization of topological spaces with respect to homology
-Our argument is best presented using topological spaces, so we formulate the analog of Theorem 8.7 for the category $\mathcal{T}\mathrm{op}$ of topological spaces.
-Theorem 12.1 ([Bou75, Theorem 10.2]).
+10.4. Rational homotopy versus rational homology
+We conclude this paper by contrasting our localization of all spaces at rational homotopy to Bousfield’s localization at rational homology.
+Our argument is best presented using topological spaces, so we first formulate Theorem 8.7 for topological spaces.
+Theorem 10.12 ([7, Theorem 10.2]).
 Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{T}\mathrm{op}$ on the category of topological spaces in which a map $f$ is
@@
 Remark 10.13.
-By [Bou97, §2.5], $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{T}\mathrm{op}$ is the left Bousfield localization of $\mathcal{T}\mathrm{op}$ at a single map, so it inherits the properties of being cellular, left proper, and simplicial.
-When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we get a model structure whose weak equivalences are the rational homology isomorphisms.
-Theorem 12.3 ([Bou75, Theorem 10.2]).
-There is a model structure $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ on the category of topological spaces in which a map $f$ is
-  1. (1)
-    a weak equivalence if and only if $f$ is a rational homology isomorphism; and
-  2. (2)
-    a cofibration if and only if $f$ is a retract of a generalized relative CW inclusion.
-This model category is cellular, left proper, and simplicial with the structure from Proposition 11.5.
-12.2. Rational homology versus rational homotopy
-We show that Bousfield’s localization $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ with respect to rational homology does not coincide with our localization $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$ with respect to rational homotopy from Theorem 11.9.
-This result stems from the fact that rational homology and rational homotopy cease to agree beyond the realm of $1$-connected spaces, and our proof comprises a particular instance of this phenomenon.
+By [9, §2.5], $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{T}\mathrm{op}$ is the left Bousfield localization of $\mathcal{T}\mathrm{op}$ at a single map, so it inherits the properties of being cellular, left proper, and simplicial.
+We show that our localization at rational homotopy $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$ differs from that at rational homology $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$.
 Proposition 10.14.
@@
 Proof.
-We show that the unique map
-$f\colon\mathbb{R}P^{2}\rightarrow|\Delta[0]|$
-is a weak equivalence in $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ but not a weak equivalence in $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$.
-Firstly, $f$ is a weak equivalence in $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ because, by Example 8.3, $f$ is a rational homology isomorphism.
-To show that $f$ is not a weak equivalence in $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$, we recall from Example 9.3 that $\mathbb{R}P^{\infty}$ is a rational space because its higher homotopy groups vanish.
-This recollection renders it sufficient to show that
-$f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(|\Delta[0]|,\mathbb{R}P^{\infty})\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(\mathbb{R}P^{2},\mathbb{R}P^{\infty})$
-is not a weak homotopy equivalence. In particular, it suffices to show that the induced map on path components
-$f^{\ast}\colon[|\Delta[0]|,\mathbb{R}P^{\infty}]\rightarrow[\mathbb{R}P^{2},\mathbb{R}P^{\infty}]$
-is not a bijection. To that end, we find a map from $\mathbb{R}P^{2}$ to $\mathbb{R}P^{\infty}$ that is not null. Such a map is the inclusion
-$i\colon\mathbb{R}P^{2}\rightarrow\mathbb{R}P^{\infty},$
-for its induced map on cohomology rings with $\mathbb{Z}/2\mathbb{Z}$-coefficients is the quotient map
-$i^{\ast}\colon(\mathbb{Z}/2\mathbb{Z})[x]\rightarrow(\mathbb{Z}/2\mathbb{Z})[x]/(x^{3}),$
-which is not trivial. Overall, $f$ is a weak equivalence in $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ but not a weak equivalence in $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$. ∎
-Thanks to Example 3.2 and Theorem 4.1(2), we can safely apply the singular complex functor throughout our proof of Proposition 12.4 to obtain the analogous result for simplicial sets.
-Proposition 12.5.
-The model structures $\mathcal{L}_{H\mathbb{Q}}\mathcal{SS}\mathrm{ets}$ and $\mathcal{L}_{\mathbb{Q}}\mathcal{SS}\mathrm{ets}$ on simplicial sets are not equal.
+We show that $c\colon\mathbb{R}P^{2}\rightarrow|\Delta[0]|$ is a weak equivalence in $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ but not in $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$.
+Firstly, $c$ is a weak equivalence in $\mathcal{L}_{H\mathbb{Q}}\mathcal{T}\mathrm{op}$ because, by Example 8.3, $c$ is a rational homology isomorphism.
+However, $c$ is not a weak equivalence in $\mathcal{L}_{\mathbb{Q}}\mathcal{T}\mathrm{op}$ because, by Example 9.3, $c$ is not a rational homotopy equivalence.
+∎